A Bibliography of Publications in SIAM Journal on Numerical Analysis: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

24 August 2024 Version 1.14

Title word cross-reference

(p, p − 1) [ZW24]. (ϑ) [AK20a]. 1 [KLMZ21]. 2 [DFFS22, HM22]. 3 [FGOS20, FLH23, Han23, HJHUT20, KLMZ21]. ^{k} [CCMP24]. A [LWC23]. H(curl) [Hu23]. H^{1} [FH24]. H(div) [FH24]. o (\frac{1}{k} ) [BN23]. C^{1} [CJZ22]. δ [HX20]. ℓ_{2} [GPHHA20]. ℓ_{∞} [GPHHA20]. Γ [GKMR22]. H [WX20]. H(curl) [CFV23]. H(curl) [WYZ20b]. H(DIV) [Mu20]. H^{1} [QW23]. H^{2} [AP20]. H^{m} [CHW22]. H ∈ (0, 1) [ND22]. hp [AH21, AP21a, AP21b, DMS21, FMMS23]. κ [DH24]. L^{2} [AXSZ22, CSWY23, DST21, LSXY24]. L^{∞} [CMZ21]. L^{p} [DST21]. L_{1} [Mus20, NT21a]. p [BDS23, BHL21, CDCVV21, CFV23, KR23a, KR23b, KR23c, MPV20, WL20]. ϕ [DL20b]. Q [GWY22]. r [ARSY20]. θ [WZZ22].

-Conforming [CHW22]. -Convergence [GKMR22]. -Curl [WL20]. -Dimensional [HJHUT20]. -Discontinuous [DMS21]. -Elliptic [Hu23]. -FEM [FMMS23, AP21a, AP21b, DL20b]. -Field [FGOS20]. -HDG [ZW24]. -Laplace-Type [BHL21]. -Laplacian [BDS23, CDCVV21]. -Navier [KR23a, KR23b, KR23c]. -Norm [QW23]. -Projection [DST21]. -Robust [CFV23, MPV20]. -Singularities [HX20]. -stability [AK20a, DST21]. -Stable [AP20, LWC23]. -Tensor [GWY22].

1-Wasserstein [XLZZ24]. 1D [MKR24].

2 [ARSY20].

3D [MKR24].

Absorption [GSZ20]. Abstract [SV24]. Acceleration [BC22, BM22, EPRX20]. Accuracy [ARSY20, HKYY21, LAZ22, MR21, WZ23, YHLR22]. Accurate [CS22a, EW21, EG22, FR21, LWZ20]. Accurately [JZZ24]. Acoustic [ZWC20]. Active [LO23]. Adaptation [CAGD22]. Adaptive [AB24, BLY21, BIP22, CN21, CM21, CLS23, EEST22, HAST22, HL23, MRV23, QZ20, WYZ20b, dVCN^{+}23]. Additive [BJ22, CHS21, KLS20, Par20, YDS21]. adjoint [CM21, LWZ20]. Advancement [AH21]. Advection [HZ20, MZZ22, SWZ20]. Affine [BG20, EEST22]. Agglomeration [ABBV23]. Agglomeration-Based [ABBV23]. Algebraic [BKSS24, MPV20, Xia21]. Algorithm [ARSY20, BdBEO21, BHYZ23, BC20, BN23, Col23, GW20, GANT20, QWL22, SSZ23, TWZZ23, ZRZ23]. Algorithmic [HHM^{+}20]. Algorithms [BW20, CL21, CCW21, CDD^{+}20, DH21, GL24, HPBL21, MR21, Ni21, SZ21a]. Allen [CGP20, JLQ22, lLTZ20]. Alternative [WR21]. American [DLZ20]. AMF [GPHHA20]. AMF-W-Methods [GPHHA20]. Ampère [BBHT24]. Amplitude [IG21]. Amplitude-based [IG21]. analogue [AK20a]. Analyses [LET20]. Analysis [AGS21, AB20, ADS23, ABCS24, BGNS20, BCJQ24, BK21, BFOQ24, BSWW22, BPS22, BH21, CJZ22, CH22a, CL21, CFV23, DN21, DEL21, DS20, DGL20, EH22, FGOS20, FA22, FGP21, Gal21, GHM21, GS23, GKS23, GMSZ22, GKMR22, GWY22, Guo21, GK24, Hal21, HY22, HN21, HS21, HS23, HPBL21, JLZ24, JZ21a, JK22, KR23a, KS20, KM20a, Kop20, KLMZ21, LS20a, LWS22, LZ23, LSXY24, LWWW22, LRK22, LZCZ22, LWXZ20, MSD22, MST22, MST20, MS20, Mus20, ND22, NR21, OS24, Otá22, Pap22, RS21a, Rub20, RN20, SGT20, SC23a, Sun21, WLPU21, Wan23, WYZ24, WYZ20a, WB21, WLF23, XZ22, YZZ20, ZZZ22b, ZN23, ZDMZ20, ZQ23, ZL21, ZW23, dVCN^{+}23, vWR23]. Analytic [CL24]. Anderson [BC22, EPRX20]. Anisotropic [AK20b, BJL23, CAGD22, Hal22]. Anisotropically [ST22]. Any [AK20b, WHL21]. AP [PCQL21]. Application [ACM24, BCT21, BGG23, CFV23, CMZ21, DH21, HLM22, JLY21, KN23, LZ23, PR22]. Applications [CS23, CL24, GP22, JLM22, Kar21, LJL^{+}21, LYL24]. Applied [BJ22, CMZ21, GANT20]. Applying [VL23]. Approach [AHW23, BHL24, CS22b, GKN21, Mai21]. Approaches [GFSZ20]. Approximability [RW24]. Approximants [NT21a]. Approximate [Lab20, WLPU21, Zha23b]. Approximated [DXY22]. Approximating [CGR20, DMM24, FL21, Ni21]. Approximation [AN21, AJH23, BPW23, BO20, BHL21, BG20, BHL24, CC23, CS23, CRR24, DLZ20, DH21, DSY21, DDO20, EHK24, EG23, FLO20, Gaw20, GSW21, GPR20, HKYY21, KR23a, KR23b, KR23c, KLMZ21, LS20b, MSS20, Mus20, NT21b, OS24, PS20, PS21, PSS22, RSS22, Spr24, WYZ20b, Xia21, ZSH20, ZZZ22b, ZW20a, ZW20b, ZDMZ20, dDFH23]. Approximations [AP21a, AP21b, ADM17, ADM21, BR20b, BLV20, BDG23, CZZ22, DGL20, EW21, GS23, GMSZ22, GKMR22, GHM20, Hel24, HL23, HK20, JZ21a, MSD22, MNO21, Wan23]. Arbitrary [CHW22, CH22b, Han23, HX20, Kam21, LMN23]. Area [JSZ23, KYB23]. Area-Preserving [JSZ23]. Arising [dDFH23]. Arnoldi [CCW21]. Artificial [DIL^{+}20, JLY21, SWZ20]. Assimilation [GANT20]. Assumptions [Kam21]. Asymptotic [DHS24, EHK24, GV20, GGHS22, GPR20, HD24]. Asymptotic-Preserving [EHK24, GV20]. Asymptotically [CHJS21, LTTF21]. Asymptotically-Preserving [CHJS21]. Attractors [DK23, LYC23]. Average [CDW24, YDS21]. Averaged [WX20]. Averages [BE24]. Axisymmetric [DN21].

B [HHM^{+}20]. B-Splines [HHM^{+}20]. Backpropagation [ABCS24]. Backward [HHL24, KV22, WZZ22]. Balance [LWWW22]. Banach [MvdZ20]. Barrier [KM20a]. Based [AHW23, ABBV23, AP24, BGNS20, BHLZ24, BdBEO21, BHYZ23, BFS23, CFV23, CCL23, CLR23, CCMP24, CDD^{+}20, DMS21, GW20, LL22a, MSD22, MRV23, PS21, Rub20, AF21, IG21, KS20, PCQL21]. Basis [AJ24, HO24]. Batch [JLL21]. Bates [BCT21]. Bathymetry [NPT22]. Bayesian [ACM24]. BDF [AK20a, ACYZ21, HS24, LWZ20, lLTZ20, LL22b, SC23b]. BDF2 [LlL22]. BDF3 [LlL22]. Beam [QS21]. Beavers [SSZ23]. Behavior [CCHHK20, MPV20]. Bellman [CFF20, FL21]. Bernoulli [DSS24]. Besov [WSH20]. Best [DH21, NT21a]. Between [SZ21a, GB20]. Beyond [GK24]. Bi [CP23]. Bi-Laplacian [CP23]. Bianchi [HLM22]. Bifurcation [SV24]. Biharmonic [CKP23, DMS21, DE22, MSS20, YZ20]. Bilinear [BFOQ24]. Biomaterials [GKMR22]. Biot [GVMRBV23]. Biperiodic [Zha23a]. Bitemperature [ADBP22]. Block [LWC23, LWX24]. Blow [CRR24, KM20b]. Blow-Up [KM20b]. Blow-Ups [CRR24]. Boltzmann [DHS24, DEL21, HQY21, PR22]. Borrvall [Pap22]. Bound [CDW24, MQ23, SX20b, CS22b]. Bound-Preserving [CDW24]. Boundary [AHW23, BFS20, CGP24, CMZ21, DFFS22, GZ21, GMSZ22, HJHUT20, HS21, JK24, JK22, KN20, Met21, NPT22, NH20, SBL22, SWZ20, WM23, YHLR22]. Bounded [HW22, LM22b, QS21, ZW22]. Bounds [BCY21, BCF22, BW24b, CZZ20, CP23, DH24, FV23, GO23, HM22, Kam21, KRS^{+}21, KM20b, WR21]. Branching [SX20a]. Bregman [BCJQ24]. Brezzi [AK20b]. Brinkman [Mu20]. Brownian [FLO20]. Bubble [BV22]. Bubble-Enriched [BV22].

C [ALOS23]. Cahn

[CSWY23, CHSZ23, CGP20, CHS21, JLQ22, lLTZ20, LMR23, Met21, QW20]. Can [CZZ20]. Cardinality [BC20]. Carleman [BdBEO21]. Carleman-Based [BdBEO21]. Carlo [CG24, FA22, GKS23, GSM24, HAST22, HZW23, KVM22, LWY23, SSX23, TS23]. Case [CL21, GHM20]. Cases [FH24]. Cauchy [BGJ22]. Cell [CDW24, GKMR22]. Cell-Induced [GKMR22]. Central [WJS23]. Certain [RW24]. Chains [BCT21]. Changing [CLR23]. Chaotic [LWXZ22]. Characteristic [MZZ22]. Characteristics [Sun21]. Characteristics-Mixed [Sun21]. Charged [HLS23, WZ21]. Charged-Particle [HLS23, WZ21]. Chebyshev [CHS20, Huy22, SST^{+}20, Wan23]. Chemotaxis [FHW21, GGC23]. Chemotaxis-Consumption [GGC23]. Chemotaxis-Stokes [FHW21]. Choice [KR20]. CIP [ZW23]. CIP-FEM [ZW23]. Class [ACM24, BCJQ24, CT21, CL24, HS21, HD24, HS24, Lab20, LWZ20, NPT22, SX20b]. Classes [AN21, RW24]. Classical [CS22a, FV23]. Close [ARSY20]. Closed [BL23, BJL23, Li21]. Clustered [LWX24]. Coefficient [BJ22, JZ21a]. Coefficients [CLR23, Hu21, Hu23, Mai21, OP24, Spr24]. Coercive [WL23]. Collocation [AB20, CL20, CCL23, EEST22, FLM21, LL22a, LTTF21]. Commutativity [FdRS24]. Compact [AKM23, CGR20, TS23]. Compact-Equivalent [AKM23]. Comparison [GB20]. Compatibility [HD24]. Compatible [BD23, LTTF21, Geo21]. Complex [AKM23, ACW21, DH21, Hal21, Hal22]. Complexes [HLM22, HZZ22]. Complexity [JK22]. Complicated [Zha23b]. Compressibility [JLY21]. Compressible [LS22]. Compression [DIL^{+}20, FP20]. Computation [ACW21, DH20, Geo21, HHM^{+}20, Huy22, LWXZ22, Sto20]. Computational [GO23, HP20, LJL^{+}21]. Computations [Col23]. Compute [CZZ20]. Computing [AB20, Col22, JLZ24, LWXZ20]. concept [AK20a]. Condition [FdRS24]. Conditions [BKSS24, BFS20, CGP24, DNT22, GSZ20, HS21, JK24, JK22, KN20, Met21, NPT22, NH20, Otá22, SBL22, SSZ23, SWZ20]. Conductor [RS21a]. Conductor-like [RS21a]. Conforming [BDL20, CJZ22, CHW22, CH22b, FK22, HLM22, KK20]. Congested [PS22]. Connected [KOO23]. Connections [SZ21a]. Conservation [AXSZ22, BR20a, CDW24, DMM24, FR21, FMR22, LZM^{+}20, LLS21, PS20]. Conservations [WZ23]. Conservative [CDP22, DWW22, TC24]. Conserving [AP21a, AP21b, AP22, GLS20, MCL20, FLM21]. Consistency [MT20, YZZ20]. Consistent [HS23]. Constant [Gal21, Sto20]. Constituted [FGOS20]. Constrained [LWY23, GZ21, HW21]. Constraints [BPW23, YZDCNC23]. Constructing [CS22b]. Consumption [GGC23]. Continuation [BDE21, BDE23]. Continuous [AHV22, BSWW22, LET20, XLZZ24]. Continuum [RS21a]. Contraction [LPW24, dVCN^{+}23]. Contractivity [SZ20]. Control [AP24, BHYZ23, BFOQ24, CC23, CL21, CGN24, CMZ21, CLR23, Col22, DH21, GL24, GO21, GMSZ22, HPBL21, LSTY21, LS24, MK20, MU22, MST22, Otá22]. Control-Based [CLR23]. Convected [MNO21]. Convection [CCHHK20, CJZ22, MST22, WX20, WLF23, ZSH20, ZRZ23]. Convergence [AP21b, ALW21, AB20, ABCS24, AKM23, BR20a, BL23, BLM21, BCJQ24, BR20b, BK21, BJ22, BCF23, BSWW22, BN23, BO20, BH21, BCT21, BDG23, CL21, CAGD22, CN21, CP23, CZZ22, CLLZ24, CAS23, CHS21, DXY22, EEST22, EH22, EPRX20, FGK23, FKM22, FMMS23, FDvW21, FMR22, GPHHA20, GKMR22, GB20, GWY22, GLW22, GGC23, HP20, Hu21, HQY21, Hu23, HPBL21, JLL21, KR23a, KR23b, KR23c, KR20, LPW24, Li20, Li21, LS22, LYC23, LWWW22, LMR23, LZCZ22, LWXZ20, MS23, NSD22, ND22, Otá22, QW23, SGT20, SC23a, WYZ20a, WSH20, WR21, Xia21, XLZZ24, XZ22, YZZ20, YC21, ZWC20, Zha22, ZN23, Zha23a]. Convergent [BNY23, CG20, JZ21b, LWX20, LWXZ22, Met21, NRY22]. Converging [EPRX20]. Convex [BLV20, BC20, CT21, KR20, Par20, SZ20, Tra24]. Convexity [JSZ23]. Convexity-Preserving [JSZ23]. Convolution [LM22a, SC23b]. Cook [QW20]. Correct [EG23]. Correction [Lab20]. Corresponding [MPV20]. Corrigendum [ADM21]. Cost [AB24]. Countably [PSS22]. Coupled [CS22c, KLMZ21, LSXY24, vWR23]. Coupling [BJMO22, DFFS22, GHM21, KV22, ZCPZ21]. Covariances [BGNS20]. Crank [AGS21, JLY21, SR20]. Cross [BPS22, CG20, DNT22]. Cross-Diffusion [BPS22, CG20]. Cross-Points [DNT22]. Crowd [LMSS20]. Crystal [BNY23]. Crystalline [BO20]. Crystals [GWY22, NRY22]. Cube [MSS20]. Cubic [CG24]. Curl [CFV23, EG23, WL20, WX20]. Curse [RW24]. Curvature [BL23, DN21, DN23, EG22, Gaw20, Li21]. Curve [JSZ23, Li20, YC21]. Curved [AW20, AJH23, CGP24, DFFS22, GHM20, NO21, WM23]. Curves [BJL23, SJ21]. Cut [BHL24].

D [DFFS22, FLH23, HM22, Han23, KLMZ21]. Darcy [FA20, JLY21, SSZ23, ZCPZ21]. Data [CDD^{+}20, Col23, DH20, DSY21, DGL20, FHK22, GANT20, Gla20, HL23, LM22a, LLMR23, LWZ23, bLWZ21, LCS24, LS20b]. Data-Based [CDD^{+}20]. Data-Driven [Col23]. Davidson [LWX24]. Decay [WS23]. decaying [ALW21]. Decomposition [AFB^{+}21, BBHT24, CDW24, DNT22, GL24, GSZ20, Hu21, KRS^{+}21, LS20b, MS21, NR21, RS21a, SSZ23, Zha23b]. Deconvolution [Lab20]. Decoupled [FHW21, LMR23, ZRZ23]. Decoupling [LUZ20]. Decreasing [JSZ23]. Deep [AN21, DGYZ22, GHLY23, HPBL21, MS23, MR21]. Defect [BO20]. Defected [YHLR22]. Defined [DL20b, SJ21]. Deformations [RSS22]. Degenerate [CT21, DE20]. Degree [GGPS21]. Degrees [CKP23]. Delay [SV24]. Densities [GKS23]. Density [GHLY23, KN23]. Dependence [HHL24]. Dependent [GVMRBV23, GH24, HK20, KS20]. Derivative [BBH22, DGG20, EW21, GAJN23, KM20a, MNO21]. Descent [JLM22]. Design [MSD22]. Detailed [LWWW22]. Deterministic [GFSZ20]. Deviations [CHJS21]. DG [ZCPZ21]. Diagonalization [GW20]. Diagonalization-Based [GW20]. Difference [AF21, AGS21, AY20, DN21, lDzSW22, FL21, FLH23, GvR24, KRS^{+}21, MNO21, MSS20, RN20, SBL22, WK23, ZW22, Zou23]. Difference-Discontinuous [WK23]. Differential [ACM24, AB20, DK23, EEST22, GNS20, GHM20, GLS23, HHL24, HT20, HK20, LL22a, LZ23, lLLL23, MS23, OS24, OSV23, SW23, SZ21a, WZZ22]. Diffraction [HN21]. Diffusion [AHW23, BZ21, BJL23, BJ22, BPS22, BCT21, CK20, CG20, CCHHK20, CJZ22, CCL23, DEL21, DL20a, EEST22, GVMRBV23, HZ20, HLTW21, JZ21a, KVM22, KV22, KR23d, LMSS20, LTTF21, LWWW22, MZZ22, MST22, Mus20, NSD22, ND22, SWZ20, WR21, WX20, YZDCNC23, ZZZ22b, ZW20a, ZW20b, ZDMZ20]. Diffusions [GHLY23]. Diffusivity [LWXZ20]. Dimension [CHW22, CH22b, GK24, HO20, HO24]. Dimensional [ADBP22, BPP20, CCMP24, GV20, GK24, HHL24, HJHUT20, HK20, JLM22, KLMZ21, LWX20, LCS24, PS21, PSS22, QS21, RW24, ZWC20, Zha23a]. Dimensionality [RW24]. Dimensions [BLY21, HZZ22, KN23, Led21, SSX23, WHL21, IG21]. Diminishing [CG20]. Dirac [BCY21]. Direct [CP23]. Dirichlet [CMZ21, GL24, GMSZ22, JK24, Zha23b]. Dirichlet-to [Zha23b]. Discontinuities [ARSY20]. Discontinuous [AXSZ22, AM22, BR20a, BCF23, CDG23, CFV24, Che21, CDP22, CHSZ23, CS22c, DHS24, DMS21, DS20, DWW22, EG23, FHN23, GM22, HX20, KR23a, KR23b, KR23c, KK24, KK20, KR23d, LET20, LR20, LZM^{+}20, LLS21, MKR24, MMD24, Pap22, Spr24, WK23, WJS23, XSZ20]. Discontinuous-Galerkin-in-Time [CS22c]. Discovery [DGYZ22, KD21]. Discrepancy [MR21]. Discrete [AM22, ADBP22, BR20b, CHKL22, CRR24, DN23, DH24, GHM21, GZ21, GV20, GWY22, GGC23, Guo21, Kar21, KK24, KN20, LYC23, lLLL23, LZCZ22, NPT22, QW20, Tra24, XLZZ24, YC21, ZZZ22b].

Discrete-Time [LYC23]. Discretization [ABB^{+}21, ACM24, BGJ22, BFOQ24, BPS22, BGG23, Che21, CHS21, DL20a, GS21, HW22, JK24, LSTY21, LMSS20, MvdZ20, SV24, SR20, YZZ20, ZZZ22a]. Discretizations [CFV22, CFV24, DRV21, DMS21, GGPS21, MST20, RN20]. Discretized [BDE23, FM21, LPW24]. Disparate [JLL21]. Dispersion [WD24, ZW23]. Dissipative [GW20, SX20b]. Distance [XLZZ24]. Distances [CWW24]. Distributed [HMO22, bLWZ21, LS24]. Distribution [GKS23, HHL24]. Div [EG23, CH22b, WX20]. div- [CH22b]. divdiv [CH22b, HLM22]. divdiv-Conforming [CH22b]. Divergence [BDL20, BHL24, FK22, Gal21, Han23, KK20, LS22, NO21, VL23, WHL21]. Divergence-Conforming [BDL20, FK22, KK20]. Divergence-Free [BHL24, Han23, LS22, WHL21, NO21]. Domain [AFB^{+}21, AJH23, BK21, BBHT24, DNT22, GL24, GSZ20, RS21a, SSZ23, WYZ20a, vWR23]. Domains [AJH23, CC23, CFV20, DL20b, GHM20, HW22, KLMZ21, LWX20, LL22b, MZZ22, NO21, QS21, SJ21, SST^{+}20, ST22, ZW22]. Douglas [AK20b]. Downscaling [GANT20]. DP [BPP20]. DPG [FH24]. Drift [BJ22, BDG23, CCW22]. Driven [BH21, Col23, CHS21, NSD22, ND22, PSS22]. DtN [BLY21]. Dual [BBH22, CK20, CWW24, WW20]. Duality [AP24, Tra24]. Duality-Based [AP24]. Dynamic [MQ23, Met21]. Dynamical [Col23, DEL21, EHK24, Guo21]. Dynamics [BCF22, DGYZ22, FV23, HLS23, KD21, LPW24, LZCZ22, WZ21, ZZZ22a]. Dziuk [BL23, Li20, Li21, YC21].

Earth [GP22]. EC [CCMP24]. Edge [WYZ20b]. Effective [LWXZ20]. Efficiency [HP20]. Efficient [BS22, Met21, PSS22, SSX23, WL23]. Eigenspaces [DE24]. Eigenstates [ST22]. Eigenvalue [ACZ24, CZZ20, CP23, CLLZ24, CL24, GLS23, HMO22, HP20, HT20, Kam21, LX23, RN20, Sch23, SC23a]. Eigenvalues [DE24, GO23, GSW21, LWX24]. Einstein [HLM22]. Elastic [BLY21, Bar20, CHH23, QS21, ZW22]. Elasticity [AKNY20, YLZ22, ZQ23]. Electric [HY22]. Electromagnetic [WYZ20a]. Element [AKNY20, ABBV23, AH21, AJH23, BL23, BLY21, BZ21, BJL23, BPW23, BLV20, BGJ22, BPP20, BDL20, BGHY22, BFS23, BFS20, BHL24, CGO20, CK20, CSWY23, CZZ20, CM21, CFV22, CZZ22, CHH23, CHW24, CAS23, DN24, DFFS22, DL20b, DGL20, DWW22, ER20, ER21, FA20, FLM21, FLH23, GS23, GANT20, Geo21, GSW21, GLW22, Han23, HG22, HL23, HJHUT20, HLM22, HZZ22, JKVY23, JK24, JLWZ22, JSZ23, JZ21a, Kam21, LSTY21, Li20, LWX20, Li21, LAZ22, LZ23, LSXY24, LYL24, LS24, LL22b, MCL20, MZZ22, MSD22, Met21, Mu20, MRV23, NRY22, QW20, QZ20, RSS22, Spr24, Sto20, WW20, WYZ20b, WX20, WB21, WLF23, ZSH20, dDFH23]. Elements [AW20, BL23, BV22, CKP23, CHW22, CH22b, Gaw20, GMMP24, Li21, NO21, WHL21, dVDV24]. Elliptic [BIP22, BKR^{+}24, CM21, CL24, DHM23, DE20, EEST22, EH22, GHM20, GMP22, HL23, Hu23, JZ21a, LWX24, Mai21, MK20, MS23, QZ20, SC23a, Spr24, WYZ20b]. Embedded [LR20, NWZ22]. Energy [ALW21, ACYZ21, AHW23, BCJQ24, BLN21, CDCVV21, DH24, DWW22, EHK24, FLM21, GMSZ22, GKMR22, GWY22, HLTW21, LSTY21, lLTZ20, LMN23, MCL20, MQ23, PCQL21, RK20, SX20b, SWW22, WB21]. Energy-based [PCQL21]. Energy-conserving [FLM21]. Energy-decaying [ALW21]. Energy-Preserving [WB21].

Energy-Stable [GWY22]. Enhanced [DGG20, PCQL21]. Enhancing [MR21]. Enriched [BV22, YLZ22]. Ensemble [AHV22, BSWW22, CST20, JLY21, SSZ23]. Entropic [CHKL22]. Entropy [BPS22, CHKL22, CG20, DMM24]. Equation [ABB^{+}21, ALW21, AGS21, BLM21, BCY21, BCF22, BW24a, BW24b, BK21, BDL20, BBHT24, BS22, BDE21, BGG23, BDE23, CS22a, CKP23, CGP24, CFV20, CMZ21, CDP22, CHSZ23, CHS21, DFFS22, DNT22, DEL21, DWW22, FLM21, GM24, Geo21, GM22, GMMP24, GSZ20, GPR20, GH24, HQY21, JLWZ22, KK20, KLS20, KM20b, LY20, LWY23, lLTZ20, LLZZ22, LS24, MZZ22, MST22, MSS20, Mus20, NSD22, ND22, NWZ22, PR22, QW23, SX20a, Sto20, TC24, WK23, WYZ24, WB21, YZ20, YHLR22, ZW22, ZW20a, ZW23, ZW24, Zou23, QW20]. Equations [ACM24, AB20, ADM17, ADM21, Avi23, BR20b, BM22, CCHHK20, CJZ22, CFF20, CN21, CFV22, CFV24, Che21, DH22, DL20a, DE20, DS20, lDzSW22, DK23, DGL20, ER21, EEST22, EH22, FHW21, FL21, FLH23, FM21, GS22, GS23, GANT20, GHM20, GLS23, GO20, HM22, HHL24, Han23, HZ20, HX20, HS20, Hu21, HS21, HS23, HS24, HK20, IG21, JKVY23, JK24, Kam21, Kar21, KN20, KS20, KV22, Kop20, KR23d, LL22a, Led21, LWX20, LS20a, LWZ20, LAZ22, LM22a, LM22b, LWS22, LMS22, LLMR23, LWZ23, LWC23, MQ23, MS23, Met21, NPT22, OSV23, PS22, QS21, QZ20, SGT20, SV24, SW23, SZ21a, SX20b, SR20, Spr24, SWZ20, WZ20, WW20, WZZ22, WS23, WJS23, XSZ20, YZDCNC23, ZZZ22b, ZW20b, ZDMZ20]. Equilibrated [CFV23]. Equivalent [AKM23]. Ergodic [CL21]. Ergodicity [TS23]. Ericksen [NRY22]. Erratum [BL23]. Error [ALOS23, AXSZ22, AGS21, ACZ24, ADM17, AP24, ACW21, BCY21, BCF22, BW24b, BPW23, BLV20, BV22, BBH22, CSWY23, CDG23, CH22a, CGN24, CGP24, CFV22, CFV24, CMZ21, CGP20, Col22, DLZ20, DN21, DEL21, DGG20, DRV21, DMS21, DH24, DS20, DGYZ22, FV23, FHW21, FGP21, Gal21, GANT20, GGPS21, GO21, GK24, HW22, HZ20, HZW23, HS21, HS23, JZ21a, KS20, KK20, KRS^{+}21, KM20a, Kop20, KR23d, KM20b, LS20a, LWS22, LSXY24, MST22, MS20, MPV20, Mus20, NT21a, NR21, Otá22, QW20, WLPU21, WL20, WZ21, Wan23, WYZ24, WR21, WB21, XSZ20, ZZZ22a, ZW20a, ZDMZ20, Zou23, dVCN^{+}23, vWR23, ADM21, JK22]. Errors [BIP22]. Estimate [AXSZ22, BBH22, DRS24, FHW21, XSZ20, ZW20a]. Estimates [AM22, ACZ24, ADM17, ADM21, BW22, BPW23, BLV20, BV22, BLN21, CGO20, CSWY23, CDG23, CGP24, CFV22, CFV24, CMZ21, CGP20, DLZ20, DRV21, DMS21, GANT20, GO21, HW22, HZ20, HG22, KR23d, MST22, QW20, SR20, WZ21, ZZZ22a]. Estimation [BIP22, CDD^{+}20, DGYZ22, GHLY23, KN23, WL20, Zou23]. Estimator [GGPS21, KK20, MPV20]. Euler [AGS21, ADBP22, BJ22, KLS20, ZZZ22a]. Eulerian [HX20, LMN23, vWR23]. Evaluation [SWZ20]. Events [WLPU21]. Evolutionary [WS23]. Evolutions [MS20]. Evolving [LL22b]. Exactly [LS22]. Expanding [ST22]. Expansion [CGR20, FDvW21, HN21, SGT20]. Expansions [ZWC20]. Expected [FP20]. Explicit [AH21, BW24a, BGG23, CFV22, CFV24, CAS23, HM22, MST20, OSV23, RK20, AK20a, LWZ20]. Exponent [BDS23]. Exponential [BW24a, BW24b, DH22, FMMS23, HKYY21, JLQ22, LM22a, LMS22, LLMR23, LWZ23, MS23, NWZ22, OSV23, ZWC20, Zha22]. Exponential-Type [NWZ22]. Extension [HHM^{+}20]. Extensions [Hu21]. Exterior [DFFS22]. Extrapolation [ZEG20]. Extrinsic [CL20, CHW24].

Face [VL23]. Factor [KK24]. Factorization [ST22]. Family [HZZ22]. Fanbeam [BH21]. Fast [BN23, DL20a, QS21, SST^{+}20, SWZ20, ZDMZ20]. FBSDEs [YZZ20]. FEAST [HT20]. FEM [AP21a, AP21b, AP22, AJ24, BIP22, BNY23, CN21, DL20b, FKM22, FMMS23, FHW21, ZW23]. FEMs [Sun21]. FETI [BPP20]. FETI-DP [BPP20]. Fibrous [GKMR22]. Field [Avi23, CSZ24, CL21, FGOS20, HN21, Kar21, LS22, LJL^{+}21, OS24, WZ21]. Fifth [HKYY21]. Fifth-Order [HKYY21]. Film [FM21]. Filter [QWL22]. Filtered [bLWZ21, WLF23]. Filtering [ACM24]. Filters [FGK23]. Finite [AF21, AKNY20, AGS21, AY20, AH21, AJH23, BL23, BLY21, BZ21, BJL23, BPW23, BLV20, BDL20, BGHY22, BFS23, BHL24, BD23, CGO20, CK20, CSWY23, CG20, CCHHK20, CZZ20, CM21, CFV22, CH22b, CZZ22, CHH23, CAS23, DN21, DN24, lDzSW22, DL20b, DGL20, ER20, ER21, FA20, FLM21, FL21, FLH23, FMR22, GvR24, GS23, GANT20, Gaw20, Geo21, GSW21, GLW22, Han23, HL23, HLM22, HZZ22, HPBL21, JKVY23, JK24, JLWZ22, JSZ23, JZZ24, JZ21a, JZ21b, Kam21, KSG23, LSTY21, Li20, LWX20, Li21, LZ23, LSXY24, LYL24, LS24, MCL20, MZZ22, MSD22, MT20, MNO21, Met21, Mu20, MRV23, MSS20, NRY22, QW20, QZ20, RN20, RSS22, SBL22, Spr24, Sto20, WW20, WK23, WX20, WB21, WLF23, ZSH20, ZW22, Zou23, dVDV24, dDFH23, BGJ22, LL22b]. Finite-Difference [SBL22]. Finite-Element [AJH23]. Finite-Volume [JZ21b]. Finitely [KOO23]. First [CSZ24, EG23, GS22, LJL^{+}21, QZ20, WYZ24]. First-Order [EG23, LJL^{+}21, QZ20, WYZ24]. Fixed [BC22, BN23, EPRX20]. Fixed-Point [EPRX20]. Flow [AFB^{+}21, AP21a, AP21b, AP22, BL23, BW22, BHL24, CLLZ24, DN21, DN23, GLS20, GLW22, HP20, HD24, JSZ23, Li20, Li21, Sun21, YC21, GWY22]. Flows [CWW24, GAJN23, GMSZ22, JLQ22, Lab20, LWXZ20, LWXZ22, NR21, Rub20, ZEG20]. Fluid [ALOS23, FK22, GP22, ZSH20, ZRZ23]. Fluid-Fluid [ALOS23]. Fluid-Structure [FK22]. Fluids [FGOS20]. Flux [BR20a, BGHY22, CFV23, ER20]. Flux-Mortar [BGHY22]. Fluxes [LZM^{+}20]. Fokker [BS22, LLZZ22, WW20]. Followed [GKS23]. Forchheimer [FA20, ZCPZ21]. Forcing [HL23]. Form [DWW22, EG23, LZ23, QZ20, Spr24, WK23]. Formulae [KYB23]. Formulation [BHYZ23, BPS22, FGOS20, GS22, GVMRBV23, GLS20, Led21]. Fortin [FH24]. Forward [HHL24]. Fourier [CLL24, HQY21, LET20, PR22, PS21, SST^{+}20, ZL21]. Fourier-Based [PS21]. Fourier-like [SST^{+}20]. Fourier-Matching [ZL21]. Fourth [CGN24, MZZ22, XSZ20]. Fourth-Order [CGN24, MZZ22, XSZ20]. Fractal [JK22]. Fractional [ADS23, BFOQ24, BLN21, BG20, BGG23, lDzSW22, FKM22, FMMS23, GO21, HW22, HZ20, Kar21, KM20a, Kop20, KLS20, Mus20, NSD22, ND22, Otá22, SST^{+}20, SSX23, WS23, ZW20a, ZW20b]. Fractional-Derivative [KM20a]. Fracture [AFB^{+}21, GHM21]. Framework [CCMP24, CS22c, LMN23, MG23, RS21b, SV24, XZ22, vWR23]. Free [BCJQ24, BHL24, DN24, Han23, LS22, NRY22, PS22, SW23, WHL21, YZ20, YLZ22, dVCN^{+}23, LLS21, NO21]. Free-Congested [PS22]. Freedom [CKP23]. Frequency [AB24, CFV22, CFV24]. Frequency-Explicit [CFV22, CFV24]. Frequency-Independent [AB24]. Friedrichs [LZM^{+}20]. Frog [JLY21]. Front [LWXZ22]. Full [BGJ22, Che21, CHS21, GAJN23, WD24]. Full-Spectrum [WD24]. Fully [BR20b, CRR24, GWY22, Guo21, QW20, SSZ23, YC21, vWR23]. Function [ACW21, BE24, CC23, GNÖ23, Led21]. Functional [AB20]. Functionals [BCJQ24]. Functions [AJ24, AY20, BCKS21, CS23, GvR24, GKS23, Gla20, HHT23, HO20, HO24, Huy22, IG21, KOO23, KSG23, KM20a, MO21, PS21, PU22, RW24, SZ21a, Xia21].

Gain [OP24]. Galerkin [AXSZ22, AM22, ADS23, ADM17, ADM21, BCF23, BPS22, BS22, CDG23, CJZ22, CSZ24, CFV24, Che21, CDP22, CHSZ23, CS22c, DLZ20, DHS24, DRV21, DMS21, DE22, DS20, EG23, FA20, FHN23, GS23, GM22, HZ20, HJHUT20, HX20, HQY21, KR23a, KR23b, KR23c, KK24, KK20, KR23d, LET20, LR20, LZM^{+}20, LLS21, MCL20, MKR24, MST20, Mu20, MvdZ20, MMD24, Pap22, PR22, SST^{+}20, WW20, WK23, WJS23, XSZ20, YZ20, YLZ22]. Galerkin-Mixed [GS23]. Galerkin/Hermite [BCF23]. Game [CSZ24, OS24]. Games [CL21, LJL^{+}21]. Gamma [NRY22]. Gamma-Convergent [NRY22]. Gauge [LWZ23, MQ23]. Gauge-Transformed [LWZ23]. Gauss [FLM21, KSG23]. Gaussian [Gaw20, Huy22, NSD22, ND22, QS21]. General [DC24, Gla20, GNÖ23, HHL24, Mai21, RS21b, ZEG20]. Generalized [CDP22, CAS23, GNS20, GK24, IG21, JK24, JLQ22, LZM^{+}20, LWZ23, MSD22, MSS20]. Generate [MU22]. Genetic [CCW22]. Genus [DN21]. Genus-0 [DN21]. Geodesics [ZN23]. Geometric [ABBV23, LRK22, TS23, WZ23, WJS23]. Geometries [BPV20]. Geothermal [QWL22]. Gevrey [CL24]. Gibbs [CHKL22]. Ginzburg [DH24, GS23, MQ23]. Glaciology [dDFH23]. Global [BLV20, HP20, JZZ24]. Gordon [BCF22, CS22a, MCL20]. Grad [EG23, WX20, SGT20]. Graded [KM20a, Mus20, ZW20b]. Gradient [BHYZ23, BC20, CWW24, CHH23, CLLZ24, DRS24, DL20a, HP20, HK20, JLM22, JLQ22, LWY23, lLLL23, Par20, SZ20, ZEG20]. Gradient-Dependent [HK20]. Gradients [GZ21]. Green [GvR24, KN20, NPT22]. Grid [XZ22]. Grids [AKNY20, BGHY22]. Gross [CLLZ24, HP20]. Guaranteed [CP23].

Halton [OP24]. Hamilton [CFF20, FL21]. Hamiltonian [Ohs23, TS23]. Harmonic [BPW23, GLW22, KOO23]. Hasimoto [BMS24]. HDG [CMZ21, FK22, LRK22, ZW24]. Heat [BDE23, GMMP24, GLW22, KM20b]. Hellan [AW20, Led21]. Helmholtz [BDE21, CFV20, CAS23, DFFS22, DNT22, GSZ20, GH24, HM22, HS20, JLWZ22, LX23, MST20, ZW23, ZW24]. Hermite [AF21, BCF23, KSG23, SGT20]. Hermite-based [AF21]. Herrmann [AW20, Led21]. Heterogeneous [CFV20, CAS23]. Heuristic [KR20]. Hidden [ZW20a]. Hidden-Memory [ZW20a]. High [ALOS23, AXSZ22, AJ24, AH21, BL23, BDE21, BGG23, BDE23, CT21, CDW24, DE22, DHM23, DH24, DE20, EW21, FLM21, FdRS24, GvR24, Gaw20, Gla20, GGHS22, GMP22, GK24, HHL24, HS21, HK20, JLWZ22, KN23, Lab20, LET20, Li21, LWC23, LCS24, LZCZ22, Mai21, PS21, RW24, SJ21, SC23b, Tra24, WZ20, YHLR22, ZZZ22a, ZW22]. High- [DH24]. High-Accuracy [YHLR22]. High-Dimensional [GK24, HHL24, HK20, PS21, RW24]. High-Index [LZCZ22, ZZZ22a]. High-Order [AH21, BL23, BDE21, BDE23, CDW24, DE22, DHM23, GvR24, Gaw20, GMP22, HS21, LET20, LWC23, Mai21, SJ21, SC23b, Tra24, CT21, DE20, FLM21, Li21, WZ20]. Higher [CG24, DRV21, DE24, Zha23a].

Higher-Order [CG24, Zha23a]. Highly [WZ23]. Hilliard [CSWY23, LMR23, QW20, CHSZ23, CHS21, Met21]. Hiptmair [Hu21, Hu23]. Hodge [WB21]. Hölder [HW22]. Homogeneous [HQY21]. Homogenization [BKR^{+}24, FGP21, Spr24]. Homogenized [Spr24]. Homographic [DH21]. Hood [GS21]. Horizon [HPBL21]. hp [AP22]. hp-FEM [AP22]. Hurst [ND22]. Hybrid [BCT21, CT21, CFV20, DE22, ER20, GMP22, MS21, Tra24]. Hybrid-Mixed [CFV20]. Hybridizable [CHSZ23, DS20]. Hybridized [BDE21]. Hydrodynamic [LWS22, LWX20]. Hydrodynamics [MT20]. Hyperbolic [CDW24, DMM24, DN23, HX20, LYL24, LLS21, SBL22, XSZ20]. Hyperinterpolation [bLWZ21]. Hyperlens [LYL24]. Hypersingular [WM23]. Hypocoercivity [Geo21]. Hypocoercivity-compatible [Geo21]. Hysteresis [PS20].

Ideal [WJS23]. Identification [JZ21a, ZZZ22b]. ii [Hu23, AP21b, BCKS21, CS22b, JK22, KR23b]. III [AP22, KR23c]. Ill [GB20]. Ill-Posed [GB20]. IMEX [GGHS22, HS21, HS24, LWS22, WYZ24]. IMEX1 [PCQL21]. IMEX1-LDG [PCQL21]. Immersed [Guo21]. Impedance [GSZ20]. Implementation [AP22]. Implicit [BR20b, BGG23, CDP22, CCMP24, CRR24, GM24, GGHS22, Li20, LWZ20, LWC23, YC21, ZEG20, AK20a, LMS22]. Implicit-explicit [LWZ20, AK20a]. Implicitly [FGOS20, SJ21]. Importance [HZW23]. Imposed [GLS20]. Imposition [BFS20]. Improved [BCF22, DHM23, WR21]. Improves [EPRX20]. Improving [HKYY21]. Including [GAJN23]. Inclusions [OS24]. Incomplete [BBC^{+}24]. Incompressible [BHL24, FGOS20, GAJN23, Han23, NR21, Rub20, SR20, Sun21]. Incremental [MS20].

Indefinite [CM21]. Independent [AB24, Huy22, MS20]. Index [LZCZ22, ND22, ZZZ22a]. Induced [GKMR22]. Inequalities [WYZ20b]. Inequality [dDFH23]. Inexact [XZ22]. Inextensible [Bar20]. Inf [Pra20, Gal21]. inf-sup [Gal21]. Infinite [AF21, CS23, JLM22, SV24]. Infinite-Dimensional [JLM22]. Influence [AJ24]. Information [BBC^{+}24]. Informed [DMM24]. Inhibiting [DGG20]. Inhomogeneous [JK24]. Initial [LM22a, NH20]. Injection [LRK22]. Instationary [GS22]. Integral [FKM22, FMMS23, GO20, HW22, KYB23, LM22b, SST^{+}20, SSX23, WM23, YHLR22]. Integrals [BHLZ24]. Integration [BMS24, CLW21, CLS23, GKS23, GSM24, GK24, IKM22, LTTF21, MO21]. Integrator [BLM21, BW24a, BW24b, DEL21, JLQ22, KLS20, LMS22, LWZ23, NWZ22]. Integrators [CS22a, DH22, Ohs23, RS21b, WZ23]. Integro [LZ23, lLLL23, OSV23]. Integro-Differential [LZ23, lLLL23, OSV23]. Interacting [JLL21, LWXZ22]. Interaction [ALOS23, DXY22, FK22]. Interactions [LSXY24]. Interface [CS22c, Guo21, SSZ23]. Interface-Coupled [CS22c]. Interfaces [ZW22]. Interpolation [AK20b, Bör22, BE24, DC24, GFSZ20, LCS24]. Interpolatory [MG23, Pra20]. Interval [BE24]. Intervals [CS23]. Invariants [IKM22, Ohs23]. Inverse [ACL20, BLX20, CZZ22, GB20, GH24, Hel24, JLM22, MZ24]. Inversion [AHV22, BSWW22, CST20]. Involution [EG23]. Involution-Preserving [EG23]. Involving [HX20]. Isogeometric [BPV20, MST22]. Isometric [RSS22]. Isoparametric [LL22b]. Itô [GHLY23]. Iterated [Bör22]. Iteration [BN23]. Iterations [AHW23, BCJQ24]. Iterative [LX23, MPV20].

Jacobi [CFF20, FL21, LWX24]. Johnson [AW20, Led21]. Johnson-like [Led21]. Joseph [SSZ23]. Jump [BCT21, Hu21, Hu23]. Jump-Diffusion [BCT21].

Kacanov [BDS23]. Kalman [AHV22, BSWW22, CST20]. Kármán [CN21]. Kawasaki [JZ21b]. KdV [LWZ23]. Keller [SX20b]. Kernel [CCL23, GB20, LTTF21, LCS24, MZ24]. Kernel-Based [CCL23]. Kernels [BG20]. Kind [LM22b, WYZ20b]. Kinetic [BS22, LPW24, SGT20]. Klein [MCL20, BCF22, CS22a]. Kolmogorov [Geo21]. Korobov [KN23]. Korteweg [CDP22]. KPP [LWXZ22]. Krasnosel’skiı [BN23]. Krylov [AKM23, BCKS21]. Kutta [AXSZ22, BKSS24, GGHS22, OSV23, RK20, SZ20, SWW22, XSZ20].

L2 [MG23, QW23]. L2-Optimal [MG23]. L2-Type [QW23]. Lacking [Pra20]. Lagrange [BHL24, CSZ24, CS22b, DLZ20, KLMZ21, LWX20]. Lagrangian [CFF20, HX20, LMSS20, LMN23, MT20]. Landau [AGS21, BCJQ24, DH24, GS23, MQ23]. Langevin [LPW24]. Laplace [BHL21, BG20]. Laplacian [ADS23, BDS23, BFOQ24, BLN21, CP23, CDCVV21, FKM22, FMMS23, HW22, HMO22, SST^{+}20, SSX23]. Large [BDS23, BM22, CCHHK20, CHJS21, HS20, RSS22, ZW24]. Large-Scale [BM22, RSS22]. Lattice [GvR24, LYC23]. Lavrentiev [BGJ22]. Law [PS20]. Laws [AXSZ22, BR20a, CDW24, DMM24, FR21, FMR22, LZM^{+}20, LLS21, SWW22]. Lax [LZM^{+}20]. Layer [JLWZ22]. Layered [HN21, ZWC20]. Layers [BK21, ZSH20, Zha22, Zha23a, ZRZ23]. LBB [Sto20]. LDG [PCQL21]. Leap [JLY21]. Leap-Frog [JLY21]. Leapfrog [CHS20, CH22a, HLS23, ZN23].

Leapfrog-Chebyshev [CHS20]. Leapfrog-Type [CH22a]. Learning [BHYZ23, BKR^{+}24, CL21, DGYZ22, GHLY23, GB20, HHL24, Hel24, MZ24, MR21]. Least [CCL23, DC24, GS22, Hel24, HS20, LY20, QZ20, Sto20]. Least-Squares [CCL23, DC24, QZ20, Sto20]. Leffler [KLS20]. LES-C [ALOS23]. Level [DL20b, LX23, LWX24, FA20]. Level-Sets [DL20b]. Lifshitz [AGS21]. Lifting [PS24]. Liftings [AP20]. like [Led21, RS21a, SST^{+}20]. Limit [AF21, AHV22, BSWW22, CS22a, DEL21, DXY22]. Limited [MST22]. Line [GKN21, ZDMZ20]. Linear [AB24, AFB^{+}21, BIP22, CHJS21, DHS24, DEL21, DGYZ22, FHW21, FK22, GM22, Hel24, IKM22, KD21, LMSS20, LYC23, MvdZ20, Ni21, SGT20, Sch23, SWW22, XSZ20, YLZ22, ZQ23]. Linearized [HLM22, KN20, LWX20, dVDV24]. Linearizing [LUZ20]. Linearly [EPRX20, Huy22, Li20, YC21]. Lines [LMN23]. Lipschitz [AY20, HW22]. Liquid [BNY23, GWY22, NRY22]. Local [BLV20, BLN21, CFV23, DXY22, FKM22, GSZ20, JK22, KR23a, KR23b, KR23c, KV22, LZM^{+}20, MS20, MRV23, ZWC20, ZDMZ20]. Localization [NT21a, Wan23]. Localized [GM24, MS21]. Locally [MSD22, YHLR22]. Locking [YLZ22]. Locking-Free [YLZ22]. Log [CS23, NT21b]. Logarithmatic [Xia21]. Logarithmic [KK24, WYZ24]. Lognormal [GK24]. Long [BCF22, Geo21, LWZ20, WS23]. Long-Time [BCF22, WS23, Geo21, LWZ20]. Loosely [LSXY24]. Lossy [MST20]. Low [BW24a, BW24b, BCKS21, CS22a, DEL21, EHK24, LMS22, MR21, NWZ22, RS21b]. Low-Discrepancy [MR21]. Low-Rank [BCKS21, DEL21, EHK24]. Low-Regularity [LMS22, NWZ22]. Lower [AH21, CZZ20, CP23, CHH23, GO23, Kop20]. Lower-Order [CHH23]. Lower-Triangular [AH21]. Lowest [Sun21]. Lowest-Order [Sun21]. lumped [DE20]. Lumping [ER21]. Lyapunov [SZ21a].

MAC [LS20a]. Machine [BHYZ23, CL21, MZ24]. Magnetic [BCY21, LS22, WZ21]. Magneto [LWX20, LWS22]. Magneto-Hydrodynamic [LWS22, LWX20]. Magnetohydrodynamics [BD23, dVDV24]. Magnetostatic [GGPS21]. Main [Hu23]. Manifolds [CL20, CCMP24, TS23]. Mann [BN23]. Mapped [SST^{+}20]. Maps [BMS24, BPW23, GLW22, WR21, Zha23b]. Marching [MMD24]. Marini [AK20b]. Markov [BCT21]. Maruyama [BJ22]. Mass [AP21a, AP21b, AP22, AH21, CCW22, DE20, ER21, FLM21, GLS20, TC24]. Mass- [FLM21]. Mass-lumped [DE20]. Matched [BK21, JLWZ22, Zha22, Zha23a]. Matching [ZL21]. Material [CDCVV21]. Matérn [BGNS20]. Matrix [AH21, BCKS21, GHM21, GLS23]. Maximal [AM22, KK24]. Maximization [CDCVV21]. Maximum [HG22, lLTZ20, MQ23]. Maximum-Principle [lLTZ20]. Maxwell [ACZ24, AJH23, CFV22, CFV24, Che21, DS20, ER21, GO23, GSW21, HY22, Hu21, HLTW21, LX23, MCL20]. McKean [HHL24]. Mean [BL23, CSZ24, CL21, DN21, HO20, HO24, Li21, LJL^{+}21, OS24]. Means [LLMR23]. Measurable [BJ22]. Measure [Col23, PSS22]. Measure-Preserving [Col23]. Media [DL20a, FA20, Sun21, ZWC20]. Median [GSM24]. Medium [ACL20, FM21]. Member [Lab20]. Membrane [BNY23]. Memory [ZW20a]. Mesh [ACL20, CAGD22]. Meshes [CDG23, CJZ22, CKP23, CGP24, ER20, Guo21, Kam21, KM20a, Mus20, QW23, YZ20, ZW20b]. Meshfree [HW21, LTTF21].

Meshless [CL20]. Metamaterials [LYL24]. Method [AB24, AXSZ22, ALW21, ACYZ21, AKNY20, ACZ24, ADS23, ABBV23, ABCS24, AW20, BL23, BLY21, BZ21, BJL23, BHLZ24, BGJ22, BPP20, BDL20, BFS20, BDE21, BDE23, BHL24, CGO20, CK20, CHKL22, CLL24, CJZ22, CCW22, CZZ20, CT21, CFV20, CDP22, CCL23, CHH23, CHSZ23, CAS23, CLR23, DHS24, DN24, DFFS22, DHM23, DSS24, DL20a, DL20b, DWW22, ER21, EH22, FA20, FL21, FLH23, FA22, FGP21, FDvW21, FMR22, FHN23, GANT20, GM22, GMP22, GO20, HZ20, HW21, HX20, HN21, HS20, HQY21, JKVY23, JLY21, JLWZ22, JSZ23, JLL21, KVM22, KK24, LL22a, LR20, Led21, Li20, LWX20, LY20, LUZ20, Li21, LAZ22, LLMR23, LX23, LWX24, lLLL23, LYL24, LL22b, LLS21, LMN23, LWXZ22, MZZ22, MU22, MRV23, Pap22, PCQL21, QS21, QW23, RS21a, SC23a, SSX23, Sto20, TC24, WW20]. Method [WK23, WYZ20a, WLF23, WM23, YZ20, YLZ22, ZW22, Zha23b, ZCPZ21, ZQ23, ZL21, ZW24]. Method-of-Lines [LMN23]. Methods [AM22, AY20, AB20, AH21, Avi23, BGNS20, BCY21, BCF22, BCKS21, BCF23, BM22, BKSS24, BGHY22, BFS23, BBHT24, BS22, BPV20, BGG23, CDG23, CH22a, CWW24, CM21, CL20, CHJS21, CMZ21, CHW24, DNT22, DE22, DS20, DGYZ22, ER20, EPRX20, FLM21, FR21, FdRS24, GL24, Geo21, GPHHA20, GGHS22, GB20, GLW22, HLS23, Hal21, HG22, HJHUT20, JLZ24, KD21, KK20, KR23d, LET20, LZM^{+}20, LWZ20, LM22b, LlL22, LZ23, LWC23, LJL^{+}21, LS24, LRK22, LMN23, MCL20, MSD22, MKR24, Mu20, MvdZ20, NRY22, NR21, OSV23, PR22, Par20, PS21, QW20, QZ20, RK20, Sch23, SZ20, SST^{+}20, SWW22, Tra24, WL23, WX20, WB21, XSZ20, XZ22, YDS21, AK20a]. Metric [CAGD22]. MHD [LS22, WJS23]. Mindlin [FHN23, GS21].

Minimal [BPV20, DC24]. Minimization [CT21, GKMR22, MS20, MvdZ20, MRV23, TWZZ23, Tra24]. Minimizations [CFV23]. Minimizers [DH24]. Minimizing [BCJQ24]. Minimum [TWZZ23]. Minimum-Surface [TWZZ23]. MINRES [FHK22]. Miscible [Sun21]. Mittag [KLS20]. Mittag-Leffler [KLS20]. Mixed [AP21a, AP21b, AP22, AKNY20, ACZ24, BGHY22, BFS23, CM21, CFV20, CHH23, ER20, FA20, GS23, GKN21, GLS20, JKVY23, KYB23, KLMZ21, MvdZ20, MRV23, SSZ23, Sun21, WB21, ZQ23]. Mixed-Dimensional [KLMZ21]. mKdV [NWZ22]. Mobility [CWW24]. Model [BNY23, BCT21, BG20, CHH23, CDD^{+}20, DIL^{+}20, DHS24, FA20, FK22, GV20, GGC23, HD24, JLY21, LZ23, NRY22, Pra20, RS21a, SSZ23, ZSH20, ZRZ23, vWR23]. Modeling [ALOS23, GHM21, MG23]. Models [AFB^{+}21, BG20, DXY22, GHM21, KS20, Lab20, lLLL23, MU22, SC23b, WLPU21, XLZZ24]. Modified [ACW21, GSW21]. Molecular [FV23]. Moment [CLL24, PR22]. Moment-Preserving [CLL24]. Moments [KYB23]. Monge [BBHT24]. Monolithic [FK22]. Monotone [BR20a, EG22, HW22, MvdZ20]. Monte [TS23, CG24, FA22, GKS23, GSM24, HAST22, HZW23, KVM22, LWY23, SSX23]. Morley [CN21, CKP23]. Mortar [BGHY22, JKVY23]. Motion [EG22, FLO20, LMSS20]. Moving [Guo21, SJ21, vWR23]. mpEDMD [Col23]. Multidegree [HHM^{+}20]. MultiDerivative [GGHS22]. MultiDimensional [MK20]. Multigrid [ABBV23, LRK22]. Multilevel [GHM20, HAST22, HK20, MPV20]. Multiple [LWX24, SSX23]. Multiplicity [DE24]. Multiplier [BHL24, CS22b]. Multipliers [KLMZ21]. Multipoint [AKNY20, BFS23, ER20]. Multipole [FDvW21, ZWC20]. Multirate [CH22a, CS22c]. Multiscale [CFV20, CAS23, DHM23, JKVY23, MS21, Mai21, QS21]. Multistage [MMD24]. Multistep [DGYZ22, KD21]. Multivariate [LL22a, PU22].

Naghdi [KN20, NPT22]. Narrow [FL21]. Narrow-stencil [FL21]. Natural [WLF23]. Navier [CSWY23, LMR23, FLH23, GANT20, Han23, HS21, HS23, JK24, KR23a, KR23b, KR23c, KS20, LS20a, LMS22, PS22, SR20]. Near [WZ23]. Nearly [BHLZ24]. Nédélec [GGPS21]. Neighborhoods [DXY22]. Nematic [NRY22]. Nernst [TC24]. Networks [AN21, AY20, ABCS24, BNY23, CFF20, CLW21, CLS23, DMM24, FGK23, FMR22, HPBL21, MS23]. Neumann [GL24, Zha23b]. Neural [ABCS24, Avi23, DMM24, HPBL21, LLZZ22]. Newton [MU22, SC23a, TWZZ23]. Nicolson [AGS21, JLY21, SR20]. Nodal [AJ24, BPW23]. Noise [BJ22, CHS21, KLS20, NSD22, ND22]. Noisy [DSY21, bLWZ21]. Non [AY20, CM21, GS23, LWZ20]. Non-Lipschitz [AY20]. Non-self-adjoint [CM21, LWZ20]. Non-Uniform [GS23]. Nonautonomous [RK20]. Noncoercive [AKM23]. Nonconforming [BDE23, CKP23, WHL21, ZW22]. Nonconvex [LWX20]. Nondivergence [QZ20, Spr24]. Nondivergence-Form [Spr24]. Nonintrusive [Ni21]. Nonlinear [AFB^{+}21, AXSZ22, BCY21, BCF22, BW24a, BW24b, BM22, CCHHK20, CWW24, CCMP24, DRV21, DE20, EH22, FLM21, FR21, JLWZ22, KV22, LZM^{+}20, LUZ20, LM22a, lLLL23, MvdZ20, MMD24, RK20, Sch23, WL23, WS23, Zou23]. Nonlinearities [HK20]. Nonlinearity [BCF22, BW24a, BW24b, CGN24]. Nonlocal [BG20, DXY22, FR21, HD24, KVM22, KV22, LTTF21, LJL^{+}21, YZDCNC23, ZDMZ20]. NonMatching [BGHY22]. Nonmonotone [GMP22]. Nonoverlapping [EH22, YDS21]. Nonregular [BDG23]. Nonrelativistic [BCY21, CS22a]. NonReversible [SZ21b]. Nonseparable [Ohs23]. Nonsmooth [BC20, BC22, LWX20, LM22a, DGL20, LM22b]. Nonuniform [QW23]. Norm [CMZ21, GPHHA20, HG22, PU22, QW23]. Normal [BBH22]. Normalization [WR21]. Note [PU22]. Novel [ACZ24, HS20, LZ23, MSD22]. Number [BGG23, GH24, JLWZ22, Lab20, NH20, ZW24]. Numbers [ALOS23, HS20]. Numerical [AHW23, ABCS24, ADBP22, BMS24, Bar20, BCT21, CL21, CDCVV21, CLR23, CCMP24, DDO20, DK23, FGOS20, FGP21, FR21, GP22, GWY22, GK24, HY22, HN21, HD24, JLZ24, LL22a, LZM^{+}20, LS22, LM22b, LYC23, MO21, OS24, Pap22, PS22, Rub20, SV24, SW23, WL23, WS23, ZRZ23, ZW20a, ZW20b, ZL21]. Nystrom [FA22].

Observable [FV23]. Observation [MST22, ZZZ22b]. Obstacle [HY22]. ODEs [AB24, CH22a]. One [DFFS22, GPHHA20]. One-Equation [DFFS22]. One-Stage [GPHHA20]. Operator [AKM23, CGR20, EW21, HJHUT20, KV22, LWWW22, MS23]. Operators [BKR^{+}24, EG23, FP20, FH24, GNS20, GNÖ23, Kar21, LR20, LWZ20, LWX24, LRK22, WD24, WM23]. Optimal [AP21b, ACL20, AGS21, BW24b, BPW23, BHYZ23, BFOQ24, BDG23, CC23, CSWY23, CCW22, CN21, CM21, CP23, CLR23, CDD^{+}20, CDW24, DH21, FLO20, GL24, GS23, GO21, HZ20, HG22, HJHUT20, KRS^{+}21, LSTY21, LSXY24, LS24, MSD22, MU22, MST22, MG23, MSS20, Otá22, ST22, WS23, WSH20, XLZZ24, ZW20b]. Optimal-Order [MSS20, ZW20b]. Optimality [CSSS23, GFSZ20, KSG23, Otá22]. Optimization [CAGD22, GZ21, GLS23, HW21, LWY23, LYC23, Pap22, Par20, SZ21a]. Optimized [BJMO22, DH21, DNT22]. Options [DLZ20]. Order [AF21, AB24, AXSZ22, AJ24, ARSY20, AK20b, AH21, BL23, BKSS24, BG20, BDE21, BDE23, CH22a, CSZ24, CM21, CGN24, CHH23, CG24, CDW24, DIL^{+}20, DRV21, DE22, DHM23, lDzSW22, DWW22, ER20, ER21, EW21, EG23, EG22, FR21, FdRS24, FHK22, GvR24, GS22, GAJN23, Gaw20, GNS20, Gla20, GGHS22, GMP22, HKYY21, Han23, HS21, HS23, KS20, KLS20, LET20, LUZ20, LWC23, lLTZ20, lLLL23, LJL^{+}21, MZZ22, Mai21, MNO21, MG23, MSS20, Mus20, NSD22, OS24, Pra20, QZ20, SJ21, SC23b, SR20, Sun21, Tra24, WK23, WYZ24, XSZ20, ZW22, Zha23a, ZW20a, ZW20b, ZDMZ20, ALW21, CT21, DE20, FLM21, Li21, WZ20]. Oriented [VL23]. Orthogonal [CS23, GO20, KRS^{+}21, LS20b, NR21, Xia21]. Orthogonalization [CCW21]. Oscillation [LLS21]. Oscillation-free [LLS21]. Oscillator [CHJS21]. Oscillatory [AB24, WZ23]. Oseen [ABB^{+}21, BGG23, KK20]. Outer [BBH22]. Oversampling [DC24].

P [CCMP24]. P- [CCMP24]. Parabolic [BR20b, BW20, BJMO22, DRV21, GL24, GO21, GPHHA20, Guo21, HS24, HK20, JKVY23, JZ21a, KM20a, Kop20, LSTY21, LWZ20, LUZ20, LAZ22, LWC23, vWR23]. Parameter [BIP22, KR20]. Parameter-Errors [BIP22]. Parametric [BZ21, BJL23, CL24, GHM20, JSZ23, Li20, LLZZ22, Pra20]. Parametrized [BG20, RSS22]. Parareal [BW20, BJMO22, GW20]. Part [AP21a, AP21b, AP22, KR23a, KR23b, KR23c]. Partial [EEST22, GHM20, LL22a, MS23, OS24]. Particle [FGK23, HLS23, LWXZ22, WZ21]. Particles [JLL21]. Parts [EW21, GNÖ23, RN20, WD24]. Patch [LY20]. PDE [GO21, GZ21, HW21, LS20b, NT21b, vWR23]. PDE-Constrained [HW21, GZ21]. PDEs [BBC^{+}24, BW20, CM21, CL20, FHK22, KLMZ21, LL22b, SST^{+}20]. Penalization [CDP22]. Penalty [BC20, DN24]. Penalty-Free [DN24]. Perfectly [BK21, JLWZ22, Zha22, Zha23a]. Perform [SJ21]. Peridynamics [CCMP24]. Perimeter [JSZ23]. Perimeter-Decreasing [JSZ23]. Periodic [AB20, BLX20, HN21, HS21, KKS20, PS21, ST22, YHLR22, Zha22]. Periodization [BGNS20]. Periodization-Based [BGNS20]. Perturbed [CK20, KR23d]. Petersson [Pap22]. Petrov [BS22, CJZ22, FHN23, MvdZ20, MMD24]. Phase [CGO20, CDCVV21, GKMR22, Kar21, MO21]. Phase-Field [Kar21]. Phenomenon [FM21]. Physics [DMM24]. Picard [HK20]. Piecewise [WHL21]. Pitaevskii [CLLZ24, HP20]. Pixel [BH21]. Pixel-Driven [BH21]. Planar [NT21b, SJ21]. Planck [BS22, LLZZ22, TC24, WW20]. Plane [DN23, IG21]. Plate [GS21]. Plates [FHN23]. PML [LZ23, WYZ20a, YHLR22]. POD [GAJN23, KS20, MU22, Rub20]. POD-ROM [Rub20]. POD-ROMs [GAJN23]. Point [BC22, BN23, EPRX20, GVMRBV23]. Points [AXSZ22, DH20, DNT22, JZZ24, OP24]. Pointwise [BLV20, DRS24, HW22, Han23, KRS^{+}21, KR23d, KM20b, Sch23]. Poisson [BCF23, CGP24, CMZ21, DDO20, GKN21, LY20, PSS22, TC24]. Pollution [ZRZ23]. Polygonal [CDG23, CKP23, GHM20]. Polygons [FMMS23]. Polyhedra [BLV20]. Polyhedral [CDG23]. Polymer [BNY23]. Polynomial [AP20, BE24, FLO20, GGPS21, NT21a, PS24]. Polynomial-Degree-Robust [GGPS21].

Polynomials [DH20, GO20]. Polytopal [YZ20]. Population [JZ21b]. Poroelasticity [BFS23, GVMRBV23]. Porous [DL20a, FA20, FM21, Sun21, ZSH20, ZRZ23]. Posed [GS22, GB20]. Posedness [FMR22, ZSH20]. Positive [CLL24, GPR20, IKM22, WJS23]. Positivity [FHW21, HLTW21, Kar21, TC24]. Positivity-Preserving [FHW21, HLTW21]. Posteriori [BW22, BBH22, CDG23, CGN24, CFV22, CFV23, CFV24, CGP20, DRV21, DMS21, Gal21, GGPS21, KK20, KR23d, KM20b, WL20, dVCN^{+}23]. Potential [BCY21, BW24a, BW24b, ZZZ22b]. Preconditioned [LX23, LWX24]. Preconditioner [Hu21, Hu23, ST22, TWZZ23]. Preconditioners [AJ24, AKM23]. Preconditioning [AP22, HJHUT20, MST22]. Preintegration [GKS23, LO23]. Presence [CRR24]. Preservation [Ohs23]. Preserving [BZ21, BPS22, CLL24, CWW24, CHJS21, CS22b, Col23, CDW24, DHS24, EHK24, EG23, FHW21, GV20, GGHS22, GPR20, HLTW21, IKM22, JSZ23, JZ21b, lLTZ20, LWXZ20, MQ23, PR22, SX20b, TC24, WL23, WD24, WB21]. Pressure [ABB^{+}21, KR23c, KS20, LR20, SR20]. Pressure-Robust [ABB^{+}21, LR20]. Primal [CWW24, WW20]. Primal-Dual [WW20]. Principle [lLTZ20, MQ23]. Principles [CHJS21]. Priori [CP23, CGP24, FGP21, MS20]. Probabilistic [YZZ20, YZDCNC23]. Probabilities [HAST22, WLPU21]. Problem [AP24, BLY21, BLV20, BGJ22, BHYZ23, BDE23, CGO20, CK20, CFV23, CMZ21, CLLZ24, CGP20, CRR24, DH21, DN24, DMS21, DE22, DDO20, GKN21, HP20, HW21, KM20a, LR20, Pap22, SC23a, Spr24, WL20, WHL21, vWR23]. Problems [ALOS23, ACL20, AHW23, ACZ24, AJH23, BIP22, BC22, BJMO22, CC23, CT21, CGN24, CZZ22, CL24, CAS23, CLR23, CS22c, DH21, DFFS22, DRV21, DHM23, GW20, GL24, GGPS21, GPHHA20, GB20, GMP22, Guo21, GH24, Hal21, Hal22, HMO22, Hel24, HL23, HY22, HT20, Hu23, HPBL21, JLM22, JZ21a, KVM22, LSTY21, LX23, MS21, Mai21, MU22, MST20, MZ24, Mu20, MvdZ20, MMD24, NH20, Pra20, RK20, SWW22, TWZZ23, Tra24, WL23, WYZ20a, WX20, WLF23, Zha22, Zha23a, ZCPZ21]. Procedure [BPV20]. Process [PSS22]. Processes [BJ22, KYB23]. Progressive [ARSY20]. Projection [AGS21, Avi23, DST21, DRS24, NRY22, Rub20, TC24, Xia21]. Projection-Based [Rub20]. Projection-Free [NRY22]. Proof [EPRX20, HQY21]. Propagation [GW20, JK22]. Proper [KRS^{+}21, LS20b, NR21]. Properties [GFSZ20, GGHS22, HHM^{+}20]. Property [CSSS23, PCQL21, dVCN^{+}23]. Provably [WJS23]. Proximal [BC20]. Pseudo [AH21]. Pseudo-Mass [AH21]. Pseudoinverse [EW21]. Pseudospectral [SV24].

Quadratic [BIP22, CGN24, Ohs23]. Quadratically [EPRX20]. Quadrature [Gla20, GHM20, Huy22, KSG23, LM22a, SJ21, SC23b, WM23]. Quantification [DE24, KKS20]. Quantum [FV23]. Quantum-Classical [FV23]. Quasi [BPW23, GFSZ20, GKS23, GSM24, HZW23, IG21, KM20a, LYC23, ST22]. Quasi-Graded [KM20a]. Quasi-Interpolation [GFSZ20]. Quasi-Linear [LYC23]. Quasi-Monte [GKS23, GSM24, HZW23]. Quasi-Optimal [BPW23, ST22]. Quasi-Trefftz [IG21]. Quasilinear [DH22, GMP22]. Quasilinearization [WJS23]. Quasiperiodic [JZZ24, JLZ24]. Quotients [KRS^{+}21].

Radial [Hal21, Hal22, HO24]. Radiation [GPR20]. Radiative [GP22, GV20, LWY23]. Radon [BH21]. Random [BLX20, CCW22, JLL21, KKS20, LWXZ20, PSS22, SX20a, SSZ23]. Randomized [DC24, HZW23]. Rank [BCKS21, CLS23, DEL21, EHK24, GLS23]. Rank-1 [GLS23]. Rank-Adaptive [CLS23]. Rapidly [MO21]. Rare [WLPU21]. Rate [AY20, BN23, BO20, BCT21, BDG23, EPRX20, HZW23, MS20, WS23]. Rate-Independent [MS20]. Rates [BR20a, CN21, CM21, CP23, GB20, KR23b, KR23c, LPW24, WSH20, Xia21]. Rational [BCKS21, HHT23, Pra20]. Reaction [AHW23, AY20, CK20, FGK23, HZ20, KR23d, LWWW22, MST22, Mus20]. Real [ZDMZ20]. Reciprocal [NT21b]. Reciprocal-Log [NT21b]. Reconstructed [LY20]. Reconstruction [BdBEO21, CFV23, LR20]. Recover [JZZ24]. Recovery [KS20, PU22]. Rectangular [CJZ22]. Rectifier [AN21]. Reduced [AFB^{+}21, CDD^{+}20, DIL^{+}20, KS20, MG23]. Reduced-Order [KS20, MG23]. Reducing [ALOS23]. Reduction [AHW23, BG20, Pra20]. Regge [Gaw20]. Regime [BCY21, CS22a, DH24]. Regression [BC20]. Regular [Hu21]. Regularities [Xia21]. Regularity [AM22, BW24a, BW24b, CS22a, CL24, HW22, HZ20, Kam21, KK24, LMS22, NWZ22, RS21b]. Regularization [CST20, GFSZ20, GB20, KR20, LSTY21, WSH20]. Regularized [CZZ22, FGK23, HL23]. Reissner [FHN23, GS21]. Relation [WD24]. Relativistic [HLS23]. Relaxation [BJMO22, Zou23]. Relaxed [BDS23]. Remap [VL23]. Renormalized [GLW22]. Reproducing [LTTF21]. Residual [BM22, DMS21, MvdZ20, MRV23].

Residual-Based [DMS21]. Resolution [HHT23]. Resonance [Hal21, Hal22]. Resonances [ZL21]. Response [Ni21]. Results [CS23, Hu23, KV22]. Retarded [AB20]. Revisitation [EW21]. Reynolds [ALOS23, BGG23, Lab20]. Ribbons [Bar20]. Ridge [HO20]. Rigorous [HN21, PCQL21]. Ritz [DRS24]. RKHS [KN23]. Robin [DH21, EH22]. Robust [ABB^{+}21, AFB^{+}21, AKM23, CFV23, CHH23, FH24, GGPS21, KK20, LR20, MST22, MPV20, Mu20, dVDV24]. ROM [Rub20]. ROMs [GAJN23]. Rotating [ALW21, MO21]. Rotation [BFS23]. Rotation-Based [BFS23]. Rough [DK23, LLMR23, LWZ23]. Rounding [MK20]. Rule [BHLZ24, KSG23]. Rules [ACW21, Gla20, Huy22, KR20]. Runge [AXSZ22, BKSS24, GGHS22, OSV23, RK20, SZ20, SWW22, XSZ20].

Saddle [GVMRBV23, LZCZ22, ZZZ22a]. Saddle-Point [GVMRBV23]. Sample [ABCS24]. Sample-Wise [ABCS24]. Sampling [BGNS20, HZW23, PU22, SZ21b]. Satisfying [FdRS24]. SAV [FLM21, JLQ22, LS20a, LWS22]. SAV-Exponential [JLQ22]. Scalar [AXSZ22, CLR23, Hal21, Hal22, IG21, LZM^{+}20, LLS21, PS20]. Scale [BM22, RSS22, WZ23]. Scaling [Hal21, Hal22]. Scattered [Gla20, LCS24]. Scattering [ACL20, BLX20, BLY21, MZ24, WYZ20a, YHLR22, Zha22, Zha23a]. Scheme [AP22, ADBP22, BDS23, BJ22, BD23, CG20, CFF20, CSZ24, DN21, EG22, FV23, FK22, GV20, GWY22, GGC23, HKYY21, Han23, HS23, HLTW21, JZ21b, LS20a, LSXY24, lLTZ20, LWWW22, LMR23, MQ23, MT20, Met21, MMD24, NSD22, WYZ24, YZZ20, YZDCNC23, YC21, Zou23]. Schemes [AF21, AFB^{+}21, ACM24, AGS21, ABBV23, BR20a, BCT21, CSWY23, CCHHK20, CHS20, CS22b, CCMP24, CDW24, DGG20, DE20, lDzSW22, HS21, HD24, HS24, IKM22, JLQ22, LWS22, LWXZ20, MS20, PS22, SBL22, SZ21b, SX20b, WZ20, WZ21, WZZ22, WJS23, ZEG20]. Schrödinger [BMS24, BW24a, BW24b, FLM21, GM22, LAZ22, ST22, WYZ24, Zou23]. Schur [SC23a]. Schwarz [BJMO22, DH21, Par20, RS21a, YDS21]. Science [GP22]. Scott [NO21]. Scrambled [OP24]. Screening [RS21a]. Screens [HJHUT20]. SDEs [FdRS24, PSS22]. Secant [BM22]. Second [AB24, ALW21, CH22a, CM21, lDzSW22, DWW22, ER20, ER21, EW21, EG22, FR21, FHK22, GNS20, HS23, LUZ20, LM22b, lLTZ20, lLLL23, MNO21, Mus20, OS24, QZ20, SR20, WK23, WYZ20b, ZDMZ20]. Second-Order [AB24, CH22a, CM21, lDzSW22, ER20, ER21, FR21, FHK22, HS23, LUZ20, lLTZ20, lLLL23, MNO21, Mus20, OS24, QZ20, ZDMZ20, ALW21]. Segel [SX20b]. Segments [BE24]. self [CM21, LWZ20]. Semi [BR20b, CFF20, CS23, LMS22]. Semi-implicit [BR20b, LMS22]. Semi-Infinite [CS23]. Semi-Lagrangian [CFF20]. Semicoercive [dDFH23]. Semiconductor [DHS24]. Semidiscrete [BL23, Li21, QW20]. Semiexplicit [Ohs23]. Semigroups [Col22]. Semilinear [AHW23, BLM21, BW20, CH22a, CGN24, HK20, Kop20, KM20b, LLMR23, MU22, OSV23, Otá22, WZ20, YZDCNC23]. Seminegative [SWW22]. Sequences [MR21]. Sequential [ACM24, BM22, LY20]. Serre [ADM17, ADM21]. Sets [DL20b, Huy22]. Setting [DDO20, GK24]. Shadowing [Ni21]. Shallow [ALW21]. Shape [GZ21]. Sharp [BO20, HM22, Kam21]. Sharper [BV22]. Shielding [HY22]. Shifted [DWW22]. Shigesada [JZ21b]. Shortening [JSZ23, Li20, YC21]. Sign [CLR23].

Sign-Changing [CLR23]. Signorini [AP24, BFS20]. Simplex [WX20]. Simplex-Averaged [WX20]. Simplicial [AKNY20]. Simplified [MU22]. Simulation [Bar20]. Simulations [JLY21]. Sinc [WZZ22, ADS23]. Sinc- [WZZ22]. sinc-Galerkin [ADS23]. Singular [BHLZ24, BR20b, CGR20, DDO20, FHK22, SC23b, WM23]. Singularities [HHT23, HX20]. Singularity [BHLZ24]. Singularly [CK20, KR23d]. Six [ACYZ21]. Six-Step [ACYZ21]. Size [ACL20]. Skeletal [CZZ20]. Slab [ZL21]. Slits [ZL21]. Slow [DL20a]. Smallest [Kam21]. Smooth [AXSZ22]. Smoothing [BC20]. Smoothness [AN21, KSG23]. Snapshots [GAJN23, MU22]. Sobolev [CLLZ24, DST21, HP20]. Solution [CL21, FP20, GAJN23, HMO22, LL22a, LS20b, SW23]. Solutions [AXSZ22, AB20, CZZ22, DMM24, FL21, Kop20, LS22, LM22b, MSS20, SX20a, WS23]. Solver [MPV20, YHLR22]. Solvers [LRK22, NT21b]. Solving [BBC^{+}24, BM22, Guo21, LZ23, MZZ22]. Some [SZ21a, WZ21]. Sonic [AXSZ22]. Source [CZZ22, GH24, MZ24, SC23b]. Sources [DDO20, GKN21, ZWC20]. Space [BPS22, BHL21, BDE23, DRV21, GM22, GMMP24, GMSZ22, HKYY21, JKVY23, JLM22, LSTY21, Led21, LY20, LS24, WL23, ZW20a, ZW20b]. Space-fractional [ZW20b]. Space-Time [BPS22, BHL21, BDE23, DRV21, GM22, GMMP24, JKVY23, LSTY21, LS24, ZW20a]. Spaces [DC24, FH24, GNÖ23, Guo21, KN23, MvdZ20, WSH20]. Sparse [FP20, GO20, KVM22]. Spatially [FM21, HQY21]. SPDEs [BDG23]. Species [JLL21]. Spectral [AB24, ACZ24, BCF23, CLL24, DWW22, GB20, GO20, HZ20, HQY21, LAZ22, LLMR23, MSD22, MS21, PR22, SST^{+}20, Wan23, WR21, Xia21, YDS21, Zha23b, ZQ23]. Spectral-Galerkin [SST^{+}20]. Spectrally [EG23]. Spectrum [GNS20, WD24]. Speeds [LWXZ22]. Sphere [bLWZ21]. Spline [LL22a]. Splines [HHM^{+}20]. Splitting [BCY21, BCF22, CSSS23, FV23, FdRS24, HS23, LWWW22, LMR23, WZ21]. Squares [CCL23, DH20, DC24, GS22, Hel24, HS20, LY20, QZ20, Sto20]. Stability [AP21a, BK21, CGO20, DST21, GGHS22, HM22, HQY21, HS21, HS23, IKM22, KV22, LWS22, LlL22, PCQL21, Pra20, QW23, RK20, SV24, Sch23, SBL22, SWW22, WL23, YZZ20, ZDMZ20, AK20a]. Stability-enhanced [PCQL21]. Stability-Preserving [WL23]. Stabilization [ABB^{+}21, AHV22, BPV20, CHW24, LUZ20, dVCN^{+}23]. Stabilization-Free [dVCN^{+}23]. Stabilized [NR21, Rub20]. Stabilizer [YZ20]. Stable [AP20, BPW23, BS22, CSWY23, DWW22, EHK24, FHW21, Gla20, GWY22, HLTW21, LWC23, lLTZ20, LMN23, MQ23, PS24]. Stage [BKSS24, GPHHA20]. Staggered [MT20, PS22, ZCPZ21]. State [AHW23, CDD^{+}20]. Static [DS20]. Stationary [GHLY23, OS24]. Statistical [GB20, Hel24, JLM22]. Steady [AHW23, FLH23, JK24]. Steady-State [AHW23]. Stefan [CRR24, HLTW21]. Stein [JLM22]. Stekloff [GSW21]. stencil [FL21]. Stencils [GvR24]. Step [ACYZ21, BGG23, LlL22, lLLL23, QWL22]. Stepping [GM24, KK24, WZ20, WLF23]. Steps [lLTZ20]. Stochastic [ACM24, ABCS24, BW22, BLM21, BSWW22, BW20, BHL21, CHJS21, Che21, CZZ22, CHS21, EEST22, FGK23, FGP21, GFSZ20, HHL24, HPBL21, KYB23, KLS20, LWXZ20, SW23, WZZ22]. Stokes [AP21a, AP21b, AP22, BLV20, BDL20, CGO20, CSWY23, DN24, DGL20, FHW21, FLH23, GS22, GANT20, GMSZ22, GLS20, Han23, HZZ22, HS21, HS23, JK24, JLY21, KR23a, KR23b, KR23c, KS20, LR20, Led21, LS20a, LMS22, LMR23, PS22, SSZ23, SR20, Sto20, WHL21, ZCPZ21, dDFH23]. Strain [CHH23]. Strang [CSSS23]. Stratification [CG24]. Stratified [CC23, GP22]. Stream [Led21]. Stress [AKNY20, GVMRBV23, GLS20]. Stress-Dependent [GVMRBV23]. Strong [BLM21, BSWW22, CHS21, GGHS22, NSD22, SBL22, WZ21]. Structure [BLX20, BZ21, BKSS24, BPS22, CWW24, CS22b, FK22, JZ21b, LSXY24, lLLL23, LWXZ20]. Structure-Preserving [BZ21, JZ21b, LWXZ20]. Structured [CCW21, GLS23]. Structures [HN21]. Study [ZRZ23]. Subdiffusion [LM22a, LLMR23, QW23, SC23b, WZ20]. Subject [BDE21]. Submanifolds [SZ21b]. Suboptimality [KSG23]. Subspace [LX23, LO23]. Substructuring [AJ24]. Subwavelength [ZL21]. Sum [DH20, MK20]. Sum-Up [MK20]. Summation [EW21, GNÖ23, RN20, WD24]. Summation-by-Parts [EW21, GNÖ23, RN20, WD24]. Sup [Pra20, Gal21]. Supercell [BO20]. Superconductivity [GS23]. Superconvergence [CJZ22, CHSZ23]. Superconvergent [MRV23]. Supercooled [CRR24]. Superlinear [AKM23]. Superposed [ZSH20, ZRZ23]. Supremizer [KS20]. Surface [BZ21, BJL23, BDL20, CCL23, DN24, TWZZ23, YHLR22, Zha23a]. Surfaces [BL23, DN21, Li21, RSS22, WM23, Zha22]. Swapping [BHLZ24]. Swept [VL23]. Symmetric [BW24a, CH22b, SC23a]. Symmetrized [BJL23, LWZ20]. Symmetry [GLS20]. Symplectic [CHJS21, Che21, Ohs23]. System [ADBP22, BCF23, CSWY23, CG20, GS22, HLM22, HLTW21, JZ21b, KR23a, KR23b, KR23c, KN20, LS22, LYC23, LWWW22, LMR23, MCL20, QWL22, QZ20, WZ23]. Systems [BM22, BPS22, BHL21, CHKL22, CSZ24, Col23, JZZ24, JLZ24, LUZ20, MK20, MKR24, Ohs23, SZ20, SBL22].

Tangential [DN24]. Taylor [GS21]. Tchebycheffian [HHM^{+}20]. Technique [ACYZ21]. Temporal [lDzSW22, GAJN23, KV22, MQ23]. Tensor [CLW21, CLS23, GWY22]. Tensors [CH22b]. Teramoto [JZ21b]. Term [SC23b]. Terminal [ZZZ22b]. Test [FH24]. Tetrahedra [AK20b, ER21]. Their [CZZ22, ZWC20, ACM24]. Theorem [VL23]. Theorems [SJ21]. Theoretical [XZ22]. Theory [LS20b]. Thermodynamically [BD23]. Thin [FM21, LSXY24]. Thin-Structure [LSXY24]. Those [EPRX20]. Three [BLY21, BPP20, HZZ22, Led21, QS21, Zha23a]. Three-Dimensional [QS21, Zha23a]. Tikhonov [CST20, KR20, WSH20]. Time [AM22, AH21, BCY21, BCF22, BK21, BSWW22, BPS22, BHL21, BGG23, BDE23, CCHHK20, CLW21, CLS23, CS22c, DRV21, lDzSW22, FV23, FM21, GM24, GL24, GANT20, GM22, GMMP24, GGC23, IKM22, JKVY23, Kar21, KK24, KS20, KRS^{+}21, Kop20, LSTY21, LYC23, LS24, MZZ22, MNO21, MMD24, Mus20, NPT22, QWL22, RN20, SBL22, WZ20, WS23, WYZ20a, WLF23, ZW20a, Geo21, LWZ20]. Time-Dependent [KS20]. Time-Discrete [AM22, NPT22]. Time-Domain [BK21, WYZ20a]. Time-Fractional [lDzSW22, Kar21, Kop20, WS23]. Time-Graded [Mus20]. Time-Marching [MMD24]. Time-Splitting [BCY21, BCF22, FV23]. Time-Stepping [KK24, WLF23]. Time-Varying [MZZ22]. Topology [Pap22]. Tori [KOO23]. Total [BW22, BJ22, TWZZ23]. Traces [PS24]. Traffic [HD24]. Training [ABCS24, MR21]. Transfer [GP22, GV20]. Transform [BMS24]. Transformed [HN21, LWZ23].

Transforms [BH21]. Transitions [GKMR22]. Translations [ZWC20]. Transmission [ACZ24, CLR23, DNT22]. Transparent [JK22, KN20]. Transport [CCW22, CWW24, GPR20, LWY23, SJ21, XLZZ24]. Trapezoidal [ACW21, BHLZ24, KSG23, WM23]. Tree [CLW21, CLS23]. Trees [JK22]. Trefftz [GM22, IG21]. Triangle [GO20]. Triangles [AP20, AJ24, AK20b, PS24]. Triangular [AH21]. Trimmed [BPV20]. Truly [GV20]. Truncation [GK24]. TVD [FR21]. Two [ADBP22, CGO20, CDCVV21, CCMP24, DGG20, FA20, GV20, Led21, LWX20, LZ23, LX23, LWX24, WZ23, XZ22, ZWC20, IG21]. Two-Derivative [DGG20]. Two-Dimensional [ADBP22, CCMP24, GV20, LWX20, ZWC20]. Two-Grid [XZ22]. Two-Level [LX23, LWX24, FA20]. Two-Phase [CGO20, CDCVV21]. Two-Scale [WZ23]. Twofold [GVMRBV23]. Type [BHL21, CH22a, DH22, GZ21, GMP22, HS24, JLQ22, KV22, NWZ22, QW23, WW20].

Unbalanced [XLZZ24]. Unbounded [SST^{+}20]. Uncertainty [DE24, KKS20]. Unconditionally [CSWY23, FHW21, IKM22, SX20b]. Unfitted [CGO20, Guo21, LL22b, MZZ22]. Unified [BGNS20, DS20, ND22, WM23, YZZ20]. Uniform [AP21a, AJ24, BCY21, BCF22, BO20, GS23, GANT20, PU22, Pra20, TWZZ23, WZ23, ZW20b]. Uniformly [CS22a, Mu20]. Unifying [MG23]. Unique [BDE21, BDE23]. Universal [GSM24]. Unstabilized [CT21]. Unsteady [FGOS20]. Unstructured [Mai21]. Updates [BCKS21]. Upper [Kop20, MNO21]. Upper-Convected [MNO21]. Ups [CRR24]. Using [ABB^{+}21, ACW21, BBH22, GHLY23, GO20, HKYY21, KKS20, KD21, Kop20, LY20, LLMR23, SJ21, HL23, KM20a].

Valuation [DLZ20]. Value

[AHW23, BHYZ23, CC23, CGR20, NH20]. Value-Gradient [BHYZ23]. Variable [lDzSW22, LlL22, lLTZ20, lLLL23, QWL22, ZW20a, ZW20b]. Variable-Order [lDzSW22, ZW20a, ZW20b]. Variable-Step [LlL22, lLLL23]. Variables [KKS20]. Variance [AY20]. Variation [BW22, BJ22, LM22b, TWZZ23]. Variational [AJH23, JLM22, LWWW22, WYZ20b, ZEG20, dDFH23]. Varying [MZZ22]. Vector [WL23]. Velocity [ADBP22, KR23b]. VEM [dVCN^{+}23]. Ventcel [CGP24]. Verlet [BLM21]. Version [BV22]. via [BMS24, CAGD22, CDP22, DGYZ22, LO23, WJS23]. Virtual [ABBV23, BV22, BPP20, CKP23, CHW22, CHW24, DFFS22, GMMP24, HG22, WHL21]. Viscosity [FL21]. Viscous [ALW21]. Vlasov [BCF23, HHL24]. Vogelius [NO21]. Volterra [GO20]. Volume [BD23, CG20, CCHHK20, FMR22, JZ21b, MT20, YZDCNC23]. Vorticity [ABB^{+}21]. VP [LLMR23]. Vries [CDP22].

W [GPHHA20]. Waiting [FM21]. Walk [SX20a]. Wasserstein [XLZZ24]. Water [ALW21]. Wave [BLM21, BLY21, BW24a, BW24b, BK21, DH22, lDzSW22, DWW22, GM24, GW20, GH24, HS20, JLWZ22, JK22, LAZ22, LS24, QS21, WK23, WB21, YHLR22, ZWC20, ZW22, ZW24]. Wave-Number-Dependent [GH24]. Wave-Type [DH22]. Waveform [BJMO22]. Waveguides [BK21, Zha23b]. Wavenumber [CAS23, HM22, MST20]. Wavenumber-Explicit [HM22]. Waves [BdBEO21, IG21, Sch23]. Weak [BCF22, BKSS24, BFS20, DMM24, DE22, MT20, Mu20, Tra24, WW20, YZ20]. Weakly [GLS20]. Weight [Gla20]. Weighted [DDO20]. Weights [BBH22, JLL21]. Well [FMR22, GS22, ZSH20]. Well-Posed [GS22]. Well-Posedness [FMR22, ZSH20]. WENO [ARSY20, HKYY21]. WENO-2 [ARSY20]. Wiener [KYB23, PSS22]. Wigner [SX20a]. Wise [ABCS24]. within [CST20]. Without [CHW24, AXSZ22, BCY21, Kam21, KK24]. wPINNs [DMM24].

Xu [Hu21, Hu23].

Ziolkowski [LZ23].

References

Ando:2020:CAC

[AB20] Alessia Andò and Dimitri Breda. Convergence analysis of collocation methods for computing periodic solutions of retarded functional differential equations. SIAM Journal on Numerical Analysis, 58(5): 3010–3039, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Agocs:2024:ASM

[AB24] Fruzsina J. Agocs and Alex H.

Barnett. An adaptive spectral method for oscillatory second-order linear ODEs with frequency-independent cost. SIAM Journal on Numerical Analysis, 62(1):295–321, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ahmed:2021:PRD

[ABB^{+}21] Naveed Ahmed, Gabriel R. Barrenechea, Erik Burman, Johnny Guzmán, Alexander Linke, and Christian Merdon. A pressure-robust discretization of Oseen’s equation using stabilization in the vorticity equation. SIAM Journal on Numerical Analysis, 59(5): 2746–2774, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Antonietti:2023:ABG

[ABBV23] Paola F. Antonietti, Stefano Berrone, Martina Busetto, and Marco Verani. Agglomerationbased geometric multigrid schemes for the virtual element method. SIAM Journal on Numerical Analysis, 61(1):223–249, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466864.

Archibald:2024:NAC

[ABCS24] Richard Archibald, Feng Bao, Yanzhao Cao, and Hui Sun. Numerical analysis for convergence of a sample-wise backpropagation method for training stochastic neural networks. SIAM Journal on Numerical Analysis, 62(2):593–621, March 2024. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ammari:2020:OMS

[ACL20] Habib Ammari, Yat Tin Chow, and Keji Liu. Optimal mesh size for inverse medium scattering problems. SIAM Journal on Numerical Analysis, 58 (1):733–756, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Akyildiz:2024:SDS

[ACM24] Ö. Deniz Akyildiz, Dan Crisan, and Joaquin Miguez. Sequential discretization schemes for a class of stochastic differential equations and their application to Bayesian filtering. SIAM Journal on Numerical Analysis, 62(2):946– 973, April 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Azah:2021:CCE

[ACW21] Mohammad Al Azah and Simon N. Chandler-Wilde. Computation of the complex error function using modified trapezoidal rules. SIAM Journal on Numerical Analysis, 59(5): 2346–2367, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Akrivis:2021:ETS

[ACYZ21] Georgios Akrivis, Minghua Chen, Fan Yu, and Zhi Zhou. The energy technique for the six-step BDF method. SIAM Journal on Numerical Analysis, 59(5):2449–2472, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

An:2024:NMS

[ACZ24] Jing An, Waixiang Cao, and Zhimin Zhang. A novel mixed spectral method and error estimates for Maxwell transmission eigenvalue problems. SIAM Journal on Numerical Analysis, 62(3):1039–1066, May 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aregba-Driollet:2022:DVN

[ADBP22] Denise Aregba-Driollet, Stéphane Brull, and Corentin Prigent. A discrete velocity numerical scheme for the two-dimensional bitemperature Euler system. SIAM Journal on Numerical Analysis, 60(1):28–51, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1407185.

Antonopoulos:2017:EEG

[ADM17] D. C. Antonopoulos, V. A. Dougalis, and D. E. Mitsotakis. Error estimates for Galerkin approximations of the Serre equations. SIAM Journal on Numerical Analysis, 55(2):841–868, ???? 2017. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See corrigendum [ADM21].

Antonopoulos:2021:CEE

[ADM21] D. C. Antonopoulos, V. A.

Dougalis, and D. E. Mitsotakis. Corrigendum: Error estimates for Galerkin approximations of the Serre equations. SIAM Journal on Numerical Analysis, 59(6):3098– 3101, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See [ADM17].

Antil:2023:ASG

[ADS23] Harbir Antil, Patrick W.

Dondl, and Ludwig Striet. Analysis of a sinc-Galerkin method for the fractional Laplacian. SIAM Journal on Numerical Analysis, 61(6): 2967–2993, December 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Abrahamsen:2021:IOL

[AF21] Dylan Abrahamsen and Bengt Fornberg. On the infinite order limit of Hermite-based finite difference schemes. SIAM Journal on Numerical Analysis, 59(4):1857–1874, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ahmed:2021:RLD

[AFB^{+}21] Elyes Ahmed, Alessio Fumagalli, Ana Budisa, Eirik Keilegavlen, Jan M. Nordbotten, and Florin A. Radu. Robust linear domain decomposition schemes for reduced nonlinear fracture flow models. SIAM Journal on Numerical Analysis, 59(1):583–612, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

An:2021:OEA

[AGS21] Rong An, Huadong Gao, and Weiwei Sun. Optimal error analysis of Euler and Crank– Nicolson projection finite difference schemes for Landau– Lifshitz equation. SIAM Journal on Numerical Analysis, 59 (3):1639–1662, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Appleton:2021:HOL

[AH21] Jay Miles Appleton and Brian T. Helenbrook. A highorder lower-triangular pseudomass matrix for explicit time advancement of hp triangular finite element methods. SIAM Journal on Numerical Analysis, 59(3):1618–1638, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Armbruster:2022:SCL

[AHV22] Dieter Armbruster, Michael Herty, and Giuseppe Visconti. A stabilization of a continuous limit of the ensemble Kalman inversion. SIAM Journal on Numerical Analysis, 60(3):1494–1515, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1414000.

Amrein:2023:NER

[AHW23] Mario Amrein, Pascal Heid, and Thomas P. Wihler. A numerical energy reduction approach for semilinear diffusion–reaction boundary value problems based on steady-state iterations. SIAM Journal on Numerical Analysis, 61(2):755–783, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1478586.

Ainsworth:2024:USP

[AJ24] Mark Ainsworth and Shuai Jiang. Uniform substructuring preconditioners for high order FEM on triangles and the influence of nodal basis functions. SIAM Journal on Numerical Analysis, 62 (4):1465–1491, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1561920.

Aylwin:2023:FED

[AJH23] Rubén Aylwin and Carlos Jerez-Hanckes. Finite-element domain approximation for Maxwell variational problems on curved domains. SIAM Journal on Numerical Analysis, 61(3):1139–1171, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1468772.

Akrivis:2020:ASC

[AK20a] Georgios Akrivis and Emmanouil Katsoprinakis. An analogue to the A(ϑ)-stability concept for implicit-explicit BDF methods. SIAM Journal on Numerical Analysis, 58(6): 3475–3503, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Apel:2020:BDM

[AK20b] Thomas Apel and Volker Kempf. Brezzi–Douglas– Marini interpolation of any order on anisotropic triangles and tetrahedra. SIAM Journal on Numerical Analysis, 58(3): 1696–1718, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Axelsson:2023:RSK

[AKM23] Owe Axelsson, János Karátson, and Frédéric Magoulès. Robust superlinear Krylov convergence for complex noncoercive compact-equivalent operator preconditioners. SIAM Journal on Numerical Analysis, 61(2):1057–1079, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466955.

Ambartsumyan:2020:MSM

[AKNY20] Ilona Ambartsumyan, Eldar Khattatov, Jan M. Nordbotten, and Ivan Yotov. A multipoint stress mixed finite element method for elasticity on simplicial grids. SIAM Journal on Numerical Analysis, 58 (1):630–656, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aggul:2023:FFI

[ALOS23] Mustafa Aggul, Alexander E.

Labovsky, Eda Onal, and Kyle J. Schwiebert. Fluid-fluid interaction problems at high Reynolds numbers: Reducing the modeling error with LES-C. SIAM Journal on Numerical Analysis, 61(2):707–732, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1494269.

Akrivis:2021:CSO

[ALW21] Georgios Akrivis, Buyang Li, and Jilu Wang. Convergence of a second-order energydecaying method for the viscous rotating shallow water equation. SIAM Journal on Numerical Analysis, 59(1): 265–288, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Akrivis:2022:MRE

[AM22] Georgios Akrivis and Charalambos Makridakis. On maximal regularity estimates for discontinuous Galerkin timediscrete methods. SIAM Journal on Numerical Analysis, 60(1):180–194, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1383781.

Ali:2021:ASC

[AN21] Mazen Ali and Anthony Nouy. Approximation of smoothness classes by deep rectifier networks. SIAM Journal on Numerical Analysis, 59(6):3032– 3051, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ainsworth:2020:SPL

[AP20] Mark Ainsworth and Charles Parker. H^{2}-stable polynomial liftings on triangles. SIAM Journal on Numerical Analysis, 58(3):1867–1892, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ainsworth:2021:MCMa

[AP21a] Mark Ainsworth and Charles Parker. Mass conserving mixed hp-FEM approximations to Stokes flow. Part I: Uniform stability. SIAM Journal on Numerical Analysis, 59 (3):1218–1244, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ainsworth:2021:MCMb

[AP21b] Mark Ainsworth and Charles Parker. Mass conserving mixed hp-FEM approximations to Stokes flow. Part II: Optimal convergence. SIAM Journal on Numerical Analysis, 59(3):1245–1272, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ainsworth:2022:MCM

[AP22] Mark Ainsworth and Charles Parker. A mass conserving mixed hp-FEM scheme for Stokes flow. Part III: Implementation and preconditioning. SIAM Journal on Numerical Analysis, 60(3):1574–1606, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1433927.

Ashby:2024:DBE

[AP24] Ben S. Ashby and Tristan Pryer. Duality-based error control for the Signorini problem. SIAM Journal on Numerical Analysis, 62 (4):1687–1712, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1534791.

Amat:2020:NWA

[ARSY20] Sergio P. Amat, Juan Ruiz, Chi-Wang Shu, and Dionisio F. Yáñez. A new WENO-2r algorithm with progressive order of accuracy close to discontinuities. SIAM Journal on Numerical Analysis, 58(6): 3448–3474, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Avitabile:2023:PMN

[Avi23] Daniele Avitabile. Projection methods for neural field equations. SIAM Journal on Numerical Analysis, 61 (2):562–591, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1463768.

Arnold:2020:HHJ

[AW20] Douglas N. Arnold and Shawn W. Walker. The Hellan–Herrmann– Johnson method with curved elements. SIAM Journal on Numerical Analysis, 58(5): 2829–2855, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ai:2022:EES

[AXSZ22] Jingqi Ai, Yuan Xu, Chi-Wang Shu, and Qiang Zhang. L^{2} error estimate to smooth solutions of high order Runge– Kutta discontinuous Galerkin method for scalar nonlinear conservation laws with and without sonic points. SIAM Journal on Numerical Analysis, 60(4):1741–1773, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1435495.

Anderson:2020:VFD

[AY20] David F. Anderson and Chaojie Yuan. Variance of finite difference methods for reaction networks with non-Lipschitz rate functions. SIAM Journal on Numerical Analysis, 58(6): 3125–3143, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bartels:2020:NSI

[Bar20] Sören Bartels. Numerical simulation of inextensible elastic ribbons. SIAM Journal on Numerical Analysis, 58(6):3332– 3354, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Binev:2024:SPI

[BBC^{+}24] Peter Binev, Andrea Bonito, Albert Cohen, Wolfgang Dahmen, Ronald DeVore, and Guergana Petrova. Solving PDEs with incomplete information. SIAM Journal on Numerical Analysis, 62(3):1278–1312, June 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1546671.

Bertoluzza:2022:PEE

[BBH22] Silvia Bertoluzza, Erik Burman, and Cuiyu He. An a posteriori error estimate of the outer normal derivative using dual weights. SIAM Journal on Numerical Analysis, 60(1):475–501, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1358219.

Boubendir:2024:DDM

[BBHT24] Yassine Boubendir, Jake Brusca, Brittany F. Hamfeldt, and Tadanaga Takahashi. Domain decomposition methods for the Monge–Ampère equation. SIAM Journal on Numerical Analysis, 62 (4):1979–2003, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1576839.

Bian:2020:SPG

[BC20] Wei Bian and Xiaojun Chen. A smoothing proximal gradient algorithm for nonsmooth convex regression with cardinality penalty. SIAM Journal on Numerical Analysis, 58 (1):858–883, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bian:2022:AAN

[BC22] Wei Bian and Xiaojun Chen.

Anderson acceleration for nonsmooth fixed point problems. SIAM Journal on Numerical Analysis, 60(5):2565–2591, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1475983.

Bao:2022:IUE

[BCF22] Weizhu Bao, Yongyong Cai, and Yue Feng. Improved uniform error bounds on timesplitting methods for longtime dynamics of the nonlinear Klein–Gordon equation with weak nonlinearity. SIAM Journal on Numerical Analysis, 60(4):1962–1984, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1449774.

Bessemoulin-Chatard:2023:CDG

[BCF23] Marianne Bessemoulin-Chatard and Francis Filbet. On the convergence of discontinuous Galerkin/Hermite spectral methods for the Vlasov– Poisson system. SIAM Journal on Numerical Analysis, 61(4):1664–1688, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1518232.

Bao:2024:CAB

[BCJQ24] Chenglong Bao, Chang Chen, Kai Jiang, and Lingyun Qiu. Convergence analysis for Bregman iterations in minimizing a class of Landau free energy functionals. SIAM Journal on Numerical Analysis, 62(1): 476–499, February 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Beckermann:2021:LRU

[BCKS21] Bernhard Beckermann, Alice Cortinovis, Daniel Kressner, and Marcel Schweitzer. Low-rank updates of matrix functions II: Rational Krylov methods. SIAM Journal on Numerical Analysis, 59(3): 1325–1347, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Briani:2021:CRM

[BCT21] Maya Briani, Lucia Caramellino, and Giulia Terenzi. Convergence rate of Markov chains and hybrid numerical schemes to jump-diffusion with application to the Bates model. SIAM Journal on Numerical Analysis, 59(1):477–502, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bao:2021:UEB

[BCY21] Weizhu Bao, Yongyong Cai, and Jia Yin. Uniform error bounds of time-splitting methods for the nonlinear Dirac equation in the nonrelativistic regime without magnetic potential. SIAM Journal on Numerical Analysis, 59(2):1040– 1066, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Busto:2023:NTC

[BD23] Saray Busto and Michael Dumbser. A new thermodynamically compatible finite volume scheme for magnetohydrodynamics. SIAM Journal on Numerical Analysis, 61(1):343–364, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M147815X.

Baudouin:2021:CBR

[BdBEO21] Lucie Baudouin, Maya de Buhan, Sylvain Ervedoza, and Axel Osses. Carleman-based reconstruction algorithm for waves. SIAM Journal on Numerical Analysis, 59(2):998–1039, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burman:2021:HHO

[BDE21] Erik Burman, Guillaume Delay, and Alexandre Ern. A hybridized high-order method for unique continuation subject to the Helmholtz equation. SIAM Journal on Numerical Analysis, 59(5):2368– 2392, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burman:2023:UCP

[BDE23] Erik Burman, Guillaume Delay, and Alexandre Ern. The unique continuation problem for the heat equation discretized with a highorder space-time nonconforming method. SIAM Journal on Numerical Analysis, 61(5): 2534–2557, October 2023. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Butkovsky:2023:ORC

[BDG23] Oleg Butkovsky, Konstantinos Dareiotis, and Máté Gerencsér. Optimal rate of convergence for approximations of SPDEs with nonregular drift. SIAM Journal on Numerical Analysis, 61(2):1103–1137, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1454213.

Bonito:2020:DCF

[BDL20] Andrea Bonito, Alan Demlow, and Martin Licht. A divergence-conforming finite element method for the surface Stokes equation. SIAM Journal on Numerical Analysis, 58 (5):2764–2798, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Balci:2023:RAS

[BDS23] Anna Kh. Balci, Lars Diening, and Johannes Storn. Relaxed kačanov scheme for the p-Laplacian with large exponent. SIAM Journal on Numerical Analysis, 61(6):2775– 2794, November 2023. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bruno:2024:PIF

[BE24] Ludovico Bruni Bruno and Wolfgang Erb. Polynomial interpolation of function averages on interval segments. SIAM Journal on Numerical Analysis, 62 (4):1759–1781, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1598271.

Bersetche:2024:BOC

[BFOQ24] Francisco Bersetche, Francisco Fuica, Enrique Otárola, and Daniel Quero. Bilinear optimal control for the fractional Laplacian: Analysis and discretization. SIAM Journal on Numerical Analysis, 62(3):1344–1371, June 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M154947X.

Burman:2020:WIS

[BFS20] Erik Burman, Stefan Frei, and Matthew W. Scroggs. Weak imposition of Signorini boundary conditions on the boundary element method. SIAM Journal on Numerical Analysis, 58(4):2334–2350, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boon:2023:MMF

[BFS23] Wietse M. Boon, Alessio Fumagalli, and Anna Scotti. Mixed and multipoint finite element methods for rotationbased poroelasticity. SIAM Journal on Numerical Analysis, 61(5):2485–2508, October 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burkovska:2020:AAP

[BG20] Olena Burkovska and Max Gunzburger. Affine approximation of parametrized kernels and model order reduction for nonlocal and fractional Laplace models. SIAM Journal on Numerical Analysis, 58 (3):1469–1494, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burman:2023:IET

[BGG23] Erik Burman, Deepika Garg, and Johnny Guzman. Implicit– explicit time discretization for Oseen’s equation at high Reynolds number with application to fractional step methods. SIAM Journal on Numerical Analysis, 61(6):2859– 2886, November 2023. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Boon:2022:FMM

[BGHY22] Wietse M. Boon, Dennis Gläser, Rainer Helmig, and Ivan Yotov. Flux-mortar mixed finite element methods on NonMatching grids. SIAM Journal on Numerical Analysis, 60(3):1193–1225, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1361407.

BenBelgacem:2022:FDC

[BGJ22] Faker Ben Belgacem, Vivette Girault, and Faten Jelassi. Full discretization of Cauchy’s problem by Lavrentiev–Finite element method. SIAM Journal on Numerical Analysis, 60(2):558–584, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1401310.

Bachmayr:2020:UAP

[BGNS20] Markus Bachmayr, Ivan G.

Graham, Van Kien Nguyen, and Robert Scheichl. Unified analysis of periodization-based sampling methods for Matérn covariances. SIAM Journal on Numerical Analysis, 58(5): 2953–2980, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bredies:2021:CAP

[BH21] Kristian Bredies and Richard Huber. Convergence analysis of pixel-driven Radon and fanbeam transforms. SIAM Journal on Numerical Analysis, 59 (3):1399–1432, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Breit:2021:STA

[BHL21] Dominic Breit, Martina Hofmanová, and Sébastien Loisel. Space-time approximation of stochastic p-Laplace-type systems. SIAM Journal on Numerical Analysis, 59(4):2218– 2236, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burman:2024:CFE

[BHL24] Erik Burman, Peter Hansbo, and Mats Larson. Cut finite element method for divergencefree approximation of incompressible flow: a Lagrange multiplier approach. SIAM Journal on Numerical Analysis, 62(2):893–918, April 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bao:2024:SSM

[BHLZ24] Gang Bao, Wenmao Hua, Jun Lai, and Jinrui Zhang. Singularity swapping method for nearly singular integrals based on trapezoidal rule. SIAM Journal on Numerical Analysis, 62(2):974–997, April 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bensoussan:2023:VGB

[BHYZ23] Alain Bensoussan, Jiayue Han, Sheung Chi Phillip Yam, and Xiang Zhou. Value-gradient based formulation of optimal control problem and machine learning algorithm. SIAM Journal on Numerical Analysis, 61(2):973–994, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1442838.

Becker:2022:AFP

[BIP22] Roland Becker, Michael Innerberger, and Dirk Praetorius. Adaptive FEM for parameter-errors in elliptic linear–quadratic parameter estimation problems. SIAM Journal on Numerical Analysis, 60(3):1450–1471, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458077.

Bencheikh:2022:CTV

[BJ22] Oumaima Bencheikh and Benjamin Jourdain. Convergence in total variation of the Euler– Maruyama scheme applied to diffusion processes with measurable drift coefficient and additive noise. SIAM Journal on Numerical Analysis, 60(4):1701–1740, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1371774.

Bao:2023:SPF

[BJL23] Weizhu Bao, Wei Jiang, and Yifei Li. A symmetrized parametric finite element method for anisotropic surface diffusion of closed curves. SIAM Journal on Numerical Analysis, 61(2):617–641, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1472851.

Bui:2022:CPO

[BJMO22] Duc Quang Bui, Caroline Japhet, Yvon Maday, and Pascal Omnes. Coupling parareal with optimized Schwarz waveform relaxation for parabolic problems. SIAM Journal on Numerical Analysis, 60 (3):913–939, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1419428.

Becache:2021:SCA

[BK21] Eliane Bécache and Maryna Kachanovska. Stability and convergence analysis of timedomain perfectly matched layers for the wave equation in waveguides. SIAM Journal on Numerical Analysis, 59(4): 2004–2039, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bhattacharya:2024:LHE

[BKR^{+}24] Kaushik Bhattacharya, Nikola B.

Kovachki, Aakila Rajan, Andrew M. Stuart, and Margaret Trautner. Learning homogenization for elliptic operators. SIAM Journal on Numerical Analysis, 62 (4):1844–1873, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1585015.

Biswas:2024:ASW

[BKSS24] Abhijit Biswas, David Ketcheson, Benjamin Seibold, and David Shirokoff. Algebraic structure of the weak stage order conditions for Runge– Kutta methods. SIAM Journal on Numerical Analysis, 62 (1):48–72, January 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bai:2023:ECD

[BL23] Genming Bai and Buyang Li. Erratum: Convergence of Dziuk’s semidiscrete finite element method for mean curvature flow of closed surfaces with high-order finite elements. SIAM Journal on Numerical Analysis, 61 (3):1609–1612, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1521791.

Banjai:2021:SCV

[BLM21] Lehel Banjai, Gabriel Lord, and Jeta Molla. Strong convergence of a Verlet integrator for the semilinear stochastic wave equation. SIAM Journal on Numerical Analysis, 59 (4):1976–2003, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Borthagaray:2021:LEE

[BLN21] Juan Pablo Borthagaray, Dmitriy Leykekhman, and Ricardo H. Nochetto. Local energy estimates for the fractional Laplacian. SIAM Journal on Numerical Analysis, 59 (4):1918–1947, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Behringer:2020:GLP

[BLV20] Niklas Behringer, Dmitriy Leykekhman, and Boris Vexler.

Global and local pointwise error estimates for finite element approximations to the Stokes problem on convex polyhedra. SIAM Journal on Numerical Analysis, 58(3):1531–1555, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lin:2021:DFH

[bLWZ21] Shao bo Lin, Yu Guang Wang, and Ding-Xuan Zhou. Distributed filtered hyperinterpolation for noisy data on the sphere. SIAM Journal on Numerical Analysis, 59(2):634– 659, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bao:2020:ISR

[BLX20] Gang Bao, Yiwen Lin, and Xiang Xu. Inverse scattering by a random periodic structure. SIAM Journal on Numerical Analysis, 58(5):2934– 2952, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bao:2021:AFE

[BLY21] Gang Bao, Peijun Li, and Xiaokai Yuan. An adaptive finite element DtN method for the elastic wave scattering problem in three dimensions. SIAM Journal on Numerical Analysis, 59(6):2900– 2925, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Birgin:2022:SAS

[BM22] Ernesto G. Birgin and J. M. MartÍnez. Secant acceleration of sequential residual methods for solving large-scale nonlinear systems of equations. SIAM Journal on Numerical Analysis, 60(6):3145–3180, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1388024.

Banica:2024:NIS

[BMS24] Valeria Banica, Georg Maierhofer, and Katharina Schratz. Numerical integration of Schrödinger maps via the Hasimoto transform. SIAM Journal on Numerical Analysis, 62(1):322– 352, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bo:2023:FKS

[BN23] Radu Ioan Bo and Dang-Khoa Nguyen. Fast Krasnosel’skiĭ– Mann algorithm with a convergence rate of the fixed point iteration of o (\frac{1}{k} ). SIAM Journal on Numerical Analysis, 61(6):2813–2843, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bouck:2023:CFM

[BNY23] Lucas Bouck, Ricardo H. Nochetto, and Shuo Yang. Convergent FEM for a membrane model of liquid crystal polymer networks. SIAM Journal on Numerical Analysis, 61(6): 2887–2916, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Braun:2020:SUC

[BO20] Julian Braun and Christoph Ortner. Sharp uniform convergence rate of the supercell approximation of a crystalline defect. SIAM Journal on Numerical Analysis, 58(1):279–297, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Borm:2022:II

[Bör22] Steffen Börm. On iterated interpolation. SIAM Journal on Numerical Analysis, 60(6):3124–3144, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1510649.

Bertoluzza:2020:FDT

[BPP20] Silvia Bertoluzza, Micol Pennacchio, and Daniele Prada. FETI-DP for the three dimensional virtual element method. SIAM Journal on Numerical Analysis, 58(3):1556–1591, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Braukhoff:2022:ESP

[BPS22] Marcel Braukhoff, Ilaria Perugia, and Paul Stocker.

An entropy structure preserving space-time formulation for cross-diffusion systems: Analysis and Galerkin discretization. SIAM Journal on Numerical Analysis, 60(1):364–395, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1360086.

Buffa:2020:MSP

[BPV20] A. Buffa, R. Puppi, and R. Vázquez. A minimal stabilization procedure for isogeometric methods on trimmed geometries. SIAM Journal on Numerical Analysis, 58(5): 2711–2735, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bartels:2023:QOE

[BPW23] Sören Bartels, Christian Palus, and Zhangxian Wang. Quasioptimal error estimates for the finite element approximation of stable harmonic maps with nodal constraints. SIAM Journal on Numerical Analysis, 61(4):1819–1834, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1524497.

Badwaik:2020:CRM

[BR20a] Jayesh Badwaik and Adrian M. Ruf. Convergence rates of monotone schemes for conservation laws with discontinuous flux. SIAM Journal on Numerical Analysis, 58(1):607–629, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bartels:2020:CFD

[BR20b] Sören Bartels and Michael Ruzicka. Convergence of fully discrete implicit and semiimplicit approximations of singular parabolic equations. SIAM Journal on Numerical Analysis, 58(1):811–833, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Brunken:2022:SEP

[BS22] Julia Brunken and Kathrin Smetana. Stable and efficient Petrov–Galerkin methods for a kinetic Fokker– Planck equation. SIAM Journal on Numerical Analysis, 60(1):157–179, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1374857.

Blomker:2022:CTL

[BSWW22] Dirk Blömker, Claudia Schillings, Philipp Wacker, and Simon Weissmann. Continuous time limit of the stochastic ensemble Kalman inversion: Strong convergence analysis. SIAM Journal on Numerical Analysis, 60(6):3181–3215, ????

2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1437561.

BeiraodaVeiga:2022:SEE

[BV22] L. Beirão da Veiga and G. Vacca. Sharper error estimates for virtual elements and a bubble-enriched version. SIAM Journal on Numerical Analysis, 60(4):1853–1878, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1411275.

Brehier:2020:PAS

[BW20] Charles-Edouard Brehier and Xu Wang. On parareal algorithms for semilinear parabolic stochastic PDEs. SIAM Journal on Numerical Analysis, 58 (1):254–278, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Banas:2022:PES

[BW22] L’ubomír Banas and André Wilke. A posteriori estimates for the stochastic total variation flow. SIAM Journal on Numerical Analysis, 60 (5):2657–2680, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1447982.

Bao:2024:ESE

[BW24a] Weizhu Bao and Chushan Wang. An explicit and symmetric exponential wave integrator for the nonlinear Schrödinger equation with low regularity potential and nonlinearity. SIAM Journal on Numerical Analysis, 62 (4):1901–1928, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1615656.

Bao:2024:OEB

[BW24b] Weizhu Bao and Chushan Wang. Optimal error bounds on the exponential wave integrator for the nonlinear Schrödinger equation with low regularity potential and nonlinearity. SIAM Journal on Numerical Analysis, 62(1):93– 118, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bao:2021:SPP

[BZ21] Weizhu Bao and Quan Zhao. A structure-preserving parametric finite element method for surface diffusion. SIAM Journal on Numerical Analysis, 59 (5):2775–2799, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carson:2022:CAM

[CAGD22] Hugh A. Carson, Steven Allmaras, Marshall Galbraith, and David Darmofal. Convergence of anisotropic mesh adaptation via metric optimization. SIAM Journal on Numerical Analysis, 60 (3):1281–1306, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1338721.

Chupeng:2023:WEC

[CAS23] Ma Chupeng, Christian Alber, and Robert Scheichl. Wavenumber explicit convergence of a multiscale generalized finite element method for heterogeneous Helmholtz problems. SIAM Journal on Numerical Analysis, 61 (3):1546–1584, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466748.

Cacace:2023:AVF

[CC23] Simone Cacace and Fabio Camilli. Approximation of the value function for optimal control problems on stratified domains. SIAM Journal on Numerical Analysis, 61(3):1172–1194, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1509370.

Cances:2020:LTB

[CCHHK20] Clément Cancès, Claire Chainais-Hillairet, Maxime Herda, and Stella Krell. Large time behavior of nonlinear finite volume schemes for convection– diffusion equations. SIAM Journal on Numerical Analysis, 58(5):2544–2571, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:2023:KBL

[CCL23] Meng Chen, Ka Chun Cheung, and Leevan Ling. A kernelbased least-squares collocation method for surface diffusion. SIAM Journal on Numerical Analysis, 61(3):1386–1404, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1444369.

Coclite:2024:NFN

[CCMP24] Alessandro Coclite, Giuseppe M.

Coclite, Francesco Maddalena, and Tiziano Politi. A numerical framework for nonlinear peridynamics on twodimensional manifolds based on implicit P-(EC)^{k} schemes. SIAM Journal on Numerical Analysis, 62(2):622–645, March 2024. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:2021:AAS

[CCW21] Pengwen Chen, Chung-Kuan Cheng, and Xinyuan Wang. Arnoldi algorithms with structured orthogonalization. SIAM Journal on Numerical Analysis, 59(1):370–400, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carrillo:2022:OMT

[CCW22] José A. Carrillo, Lin Chen, and Qi Wang. An optimal mass transport method for random genetic drift. SIAM Journal on Numerical Analysis, 60(3):940–969, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1389431.

Casado-Diaz:2021:NML

[CDCVV21] Juan Casado-Díaz, Carlos Conca, and Donato Vásquez-Varas. Numerical maximization of the p-Laplacian energy of a two-phase material. SIAM Journal on Numerical Analysis, 59(6):3077–3097, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cohen:2020:ORM

[CDD^{+}20] Albert Cohen, Wolfgang Dahmen, Ronald DeVore, Jalal Fadili, Olga Mula, and James Nichols. Optimal reduced model algorithms for databased state estimation. SIAM Journal on Numerical Analysis, 58(6):3355–3381, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cangiani:2023:PEE

[CDG23] Andrea Cangiani, Zhaonan Dong, and Emmanuil H. Georgoulis. A posteriori error estimates for discontinuous Galerkin methods on polygonal and polyhedral meshes. SIAM Journal on Numerical Analysis, 61(5):2352–2380, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1516701.

Chen:2022:NCD

[CDP22] Yanlai Chen, Bo Dong, and Rebecca Pereira. A new conservative discontinuous Galerkin method via implicit penalization for the generalized Korteweg–de Vries equation. SIAM Journal on Numerical Analysis, 60 (6):3078–3098, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1470827.

Cui:2024:OCA

[CDW24] Shumo Cui, Shengrong Ding, and Kailiang Wu. On optimal cell average decomposition for high-order boundpreserving schemes of hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 62(2):775–810, March 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carlini:2020:SLS

[CFF20] E. Carlini, A. Festa, and N. Forcadel. A semi-Lagrangian scheme for Hamilton– Jacobi–Bellman equations on networks. SIAM Journal on Numerical Analysis, 58(6): 3165–3196, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chaumont-Frelet:2020:MHM

[CFV20] Théophile Chaumont-Frelet and Frédéric Valentin. A multiscale hybrid-mixed method for the Helmholtz equation in heterogeneous domains. SIAM Journal on Numerical Analysis, 58(2):1029–1067, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chaumont-Frelet:2022:FEP

[CFV22] Théophile Chaumont-Frelet and Patrick Vega. Frequencyexplicit a posteriori error estimates for finite element discretizations of Maxwell’s equations. SIAM Journal on Numerical Analysis, 60 (4):1774–1798, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1421805.

Chaumont-Frelet:2023:REF

[CFV23] Théophile Chaumont-Frelet and Martin Vohralík. p-robust equilibrated flux reconstruction in H(curl) based on local minimizations: Application to a posteriori analysis of the curl–curl problem. SIAM Journal on Numerical Analysis, 61(4):1783–1818, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M141909X.

Chaumont-Frelet:2024:FEP

[CFV24] Théophile Chaumont-Frelet and Patrick Vega. Frequencyexplicit a posteriori error estimates for discontinuous Galerkin discretizations of Maxwell’s equations. SIAM Journal on Numerical Analysis, 62(1):400–421, February 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cances:2020:CED

[CG20] Clément Cancès and Benoît Gaudeul. A convergent entropy diminishing finite volume scheme for a crossdiffusion system. SIAM Journal on Numerical Analysis, 58 (5):2684–2710, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chopin:2024:HOM

[CG24] Nicolas Chopin and Mathieu Gerber. Higher-order Monte Carlo through cubic stratification. SIAM Journal on Numerical Analysis, 62(1):229– 247, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:2024:PEC

[CGN24] Carsten Carstensen, Benedikt Gräßle, and Neela Nataraj. A posteriori error control for fourth-order semilinear problems with quadratic nonlinearity. SIAM Journal on Numerical Analysis, 62(2):919– 945, April 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Caceres:2020:NSE

[CGO20] Ernesto Cáceres, Johnny Guzmán, and Maxim Olshanskii. New stability estimates for an unfitted finite element method for two-phase Stokes problem. SIAM Journal on Numerical Analysis, 58(4): 2165–2192, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chrysafinos:2020:PEE

[CGP20] Konstantinos Chrysafinos, Emmanuil H. Georgoulis, and Dimitra Plaka. A posteriori error estimates for the Allen– Cahn problem. SIAM Journal on Numerical Analysis, 58(5): 2662–2683, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Caubet:2024:PEE

[CGP24] Fabien Caubet, Joyce Ghantous, and Charles Pierre. A priori error estimates of a Poisson equation with Ventcel boundary conditions on curved meshes. SIAM Journal on Numerical Analysis, 62 (4):1929–1955, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1582497.

Crane:2020:ASV

[CGR20] Daniel K. Crane, Mark S.

Gockenbach, and Matthew J. Roberts. Approximating the singular value expansion of a compact operator. SIAM Journal on Numerical Analysis, 58(2):1295–1318, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carle:2022:EAM

[CH22a] Constantin Carle and Marlis Hochbruck. Error analysis of multirate leapfrog-type methods for second-order semilinear ODEs. SIAM Journal on Numerical Analysis, 60 (5):2897–2924, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1427255.

Chen:2022:FED

[CH22b] Long Chen and Xuehai Huang.

Finite elements for div- and divdiv-conforming symmetric tensors in arbitrary dimension. SIAM Journal on Numerical Analysis, 60(4):1932–1961, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1433708.

Chen:2021:SDG

[Che21] Chuchu Chen. A symplectic discontinuous Galerkin full discretization for stochastic Maxwell equations. SIAM Journal on Numerical Analysis, 59(4):2197–2217, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:2023:RLO

[CHH23] Mingqing Chen, Jianguo Huang, and Xuehai Huang. A robust lower-order mixed finite element method for a strain gradient elastic model. SIAM Journal on Numerical Analysis, 61(5):2237–2260, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M151649X.

Chen:2021:APL

[CHJS21] Chuchu Chen, Jialing Hong, Diancong Jin, and Liying Sun. Asymptoticallypreserving large deviations principles by stochastic symplectic methods for a linear stochastic oscillator. SIAM Journal on Numerical Analysis, 59(1):32–59, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cai:2022:EMD

[CHKL22] Zhenning Cai, Jingwei Hu, Yang Kuang, and Bo Lin. An entropic method for discrete systems with Gibbs entropy. SIAM Journal on Numerical Analysis, 60(4):2345–2371, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1429023.

Carle:2020:LCS

[CHS20] Constantin Carle, Marlis Hochbruck, and Andreas Sturm. On leapfrog-Chebyshev schemes. SIAM Journal on Numerical Analysis, 58(4): 2404–2433, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cui:2021:SCF

[CHS21] Jianbo Cui, Jialin Hong, and Liying Sun. Strong convergence of full discretization for stochastic Cahn–Hilliard equation driven by additive noise. SIAM Journal on Numerical Analysis, 59(6):2866–2899, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:2023:SHD

[CHSZ23] Gang Chen, Daozhi Han, John R. Singler, and Yangwen Zhang. On the superconvergence of a hybridizable discontinuous Galerkin method for the Cahn–Hilliard equation. SIAM Journal on Numerical Analysis, 61(1):83–109, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1437780.

Chen:2022:CVE

[CHW22] Chunyu Chen, Xuehai Huang, and Huayi Wei. H^{m}- conforming virtual elements in arbitrary dimension. SIAM Journal on Numerical Analysis, 60(6):3099–3123, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1440323.

Chen:2024:VEM

[CHW24] Chunyu Chen, Xuehai Huang, and Huayi Wei. Virtual element methods without extrinsic stabilization. SIAM Journal on Numerical Analysis, 62(1):567–591, February 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cao:2022:CPG

[CJZ22] Waixiang Cao, Lueling Jia, and Zhimin Zhang. A C^{1} conforming Petrov–Galerkin method for convection–diffusion equations and superconvergence analysis over rectangular meshes. SIAM Journal on Numerical Analysis, 60(1):274–311, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1391006.

Cai:2020:DFE

[CK20] Zhiqiang Cai and JaEun Ku.

A dual finite element method for a singularly perturbed reaction–diffusion problem. SIAM Journal on Numerical Analysis, 58(3):1654–1673, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:2023:NVE

[CKP23] Carsten Carstensen, Rekha Khot, and Amiya K. Pani. Nonconforming virtual elements for the biharmonic equation with Morley degrees of freedom on polygonal meshes. SIAM Journal on Numerical Analysis, 61(5):2460– 2484, October 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:2020:EMC

[CL20] Meng Chen and Leevan Ling.

Extrinsic meshless collocation methods for PDEs on manifolds. SIAM Journal on Numerical Analysis, 58(2):988– 1007, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carmona:2021:CAM

[CL21] René Carmona and Mathieu Laurière. Convergence analysis of machine learning algorithms for the numerical solution of mean field control and games I: The ergodic case. SIAM Journal on Numerical Analysis, 59 (3):1455–1485, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chernov:2024:AGC

[CL24] Alexey Chernov and Tùng Lê. Analytic and Gevrey class regularity for parametric elliptic eigenvalue problems and applications. SIAM Journal on Numerical Analysis, 62(4):1874–1900, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1596296.

Cai:2024:PMP

[CLL24] Zhenning Cai, Bo Lin, and Meixia Lin. A positive and moment-preserving Fourier spectral method. SIAM Journal on Numerical Analysis, 62(1):273–294, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:2024:CSG

[CLLZ24] Ziang Chen, Jianfeng Lu, Yulong Lu, and Xiangxiong Zhang. On the convergence of Sobolev gradient flow for the Gross–Pitaevskii eigenvalue problem. SIAM Journal on Numerical Analysis, 62 (2):667–691, March 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ciarlet:2023:OCB

[CLR23] Patrick Ciarlet, Jr., David Lassounon, and Mahran Rihani. An optimal controlbased numerical method for scalar transmission problems with sign-changing coefficients. SIAM Journal on Numerical Analysis, 61 (3):1316–1339, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1495998.

Ceruti:2023:RAT

[CLS23] Gianluca Ceruti, Christian Lubich, and Dominik Sulz. Rankadaptive time integration of tree tensor networks. SIAM Journal on Numerical Analysis, 61(1):194–222, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1473790.

Ceruti:2021:TIT

[CLW21] Gianluca Ceruti, Christian Lubich, and Hanna Walach. Time integration of tree tensor networks. SIAM Journal on Numerical Analysis, 59(1):289– 313, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:2021:AMFb

[CM21] Carsten Carstensen and Rui Ma. Adaptive mixed finite element methods for non-self-adjoint indefinite second-order elliptic PDEs with optimal rates. SIAM Journal on Numerical Analysis, 59(2):955– 982, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:2021:NEE

[CMZ21] Gang Chen, Peter B. Monk, and Yangwen Zhang. L^{∞} norm error estimates for HDG methods applied to the Poisson equation with an application to the Dirichlet boundary control problem. SIAM Journal on Numerical Analysis, 59 (2):720–745, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:2021:AMFa

[CN21] Carsten Carstensen and Neela Nataraj. Adaptive Morley FEM for the von Kármán equations with optimal convergence rates. SIAM Journal on Numerical Analysis, 59(2): 696–719, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Colbrook:2022:CSE

[Col22] Matthew J. Colbrook. Computing semigroups with error control. SIAM Journal on Numerical Analysis, 60(1):396–422, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1398616.

Colbrook:2023:MAD

[Col23] Matthew J. Colbrook. The mpEDMD algorithm for datadriven computations of measurepreserving dynamical systems. SIAM Journal on Numerical Analysis, 61(3):1585–1608, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1521407.

Carstensen:2023:DGL

[CP23] Carsten Carstensen and Sophie Puttkammer. Direct guaranteed lower eigenvalue bounds with optimal a priori convergence rates for the bi-Laplacian. SIAM Journal on Numerical Analysis, 61(2):812–836, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M139921X.

Cuchiero:2024:IFD

[CRR24] Christa Cuchiero, Christoph Reisinger, and Stefan Rigger. Implicit and fully discrete approximation of the supercooled Stefan problem in the presence of blow-ups. SIAM Journal on Numerical Analysis, 62(3): 1145–1170, May 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Calvo:2022:UAL

[CS22a] María Cabrera Calvo and Katharina Schratz. Uniformly accurate low regularity integrators for the Klein–Gordon equation from the classical to NonRelativistic limit regime. SIAM Journal on Numerical Analysis, 60(2):888–912, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1415030.

Cheng:2022:NLM

[CS22b] Qing Cheng and Jie Shen. A new Lagrange multiplier approach for constructing structure preserving schemes, II. Bound preserving. SIAM Journal on Numerical Analysis, 60(3):970–998, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M144877X.

Connors:2022:MDG

[CS22c] Jeffrey M. Connors and Kenneth C. Sockwell. A multirate discontinuous-Galerkin-in-time framework for interfacecoupled problems. SIAM Journal on Numerical Analysis, 60(5):2373–2404, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1461149.

Chen:2023:LOF

[CS23] Sheng Chen and Jie Shen. Log orthogonal functions in semiinfinite intervals: Approximation results and applications. SIAM Journal on Numerical Analysis, 61(1):110–134, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466840.

Casas:2023:NOP

[CSSS23] Fernando Casas, Jesús María Sanz-Serna, and Luke Shaw. A new optimality property of Strang’s splitting. SIAM Journal on Numerical Analysis, 61(3):1369–1385, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1528690.

Chada:2020:TRW

[CST20] Neil K. Chada, Andrew M.

Stuart, and Xin T. Tong.

Tikhonov regularization within ensemble Kalman inversion. SIAM Journal on Numerical Analysis, 58(2):1263–1294, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cai:2023:OEE

[CSWY23] Wentao Cai, Weiwei Sun, Jilu Wang, and Zongze Yang. Optimal L^{2} error estimates of unconditionally stable finite element schemes for the Cahn– Hilliard–Navier–Stokes system. SIAM Journal on Numerical Analysis, 61(3):1218–1245, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1486844.

Carlini:2024:LGS

[CSZ24] Elisabetta Carlini, Francisco J. Silva, and Ahmad Zorkot. A Lagrange–Galerkin scheme for first order mean field game systems. SIAM Journal on Numerical Analysis, 62(1):167– 198, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:2021:UHH

[CT21] Carsten Carstensen and Tien Tran. Unstabilized hybrid high-order method for a class of degenerate convex minimization problems. SIAM Journal on Numerical Analysis, 59(3):1348–1373, ???? 2021. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Carrillo:2024:SPP

[CWW24] José A. Carrillo, Li Wang, and Chaozhen Wei. Structure preserving primal dual methods for gradient flows with nonlinear mobility transport distances. SIAM Journal on Numerical Analysis, 62(1):376– 399, February 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Carstensen:2020:SFE

[CZZ20] Carsten Carstensen, Qilong Zhai, and Ran Zhang. A skeletal finite element method can compute lower eigenvalue bounds. SIAM Journal on Numerical Analysis, 58(1):109– 124, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Chen:2022:SCR

[CZZ22] Zhiming Chen, Wenlong Zhang, and Jun Zou. Stochastic convergence of regularized solutions and their finite element approximations to inverse source problems. SIAM Journal on Numerical Analysis, 60(2):751–780, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1409779.

Dolbeault:2024:RLS

[DC24] Matthieu Dolbeault and Moulay Abdellah Chkifa. Randomized least-squares with minimal oversampling and interpolation in general spaces. SIAM Journal on Numerical Analysis, 62(4):1515–1538, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M160178X.

deDiego:2023:FEA

[dDFH23] Gonzalo G. de Diego, Patrick E. Farrell, and Ian J. Hewitt. On the finite element approximation of a semicoercive Stokes variational inequality arising in glaciology. SIAM Journal on Numerical Analysis, 61(1):1–25, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1437640.

Drelichman:2020:WSN

[DDO20] Irene Drelichman, Ricardo G. Durán, and Ignacio Ojea. A weighted setting for the numerical approximation of the Poisson problem with singular sources. SIAM Journal on Numerical Analysis, 58(1): 590–606, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Droniou:2020:HOM

[DE20] Jerome Droniou and Robert Eymard. High-order masslumped schemes for nonlinear degenerate elliptic equations. SIAM Journal on Numerical Analysis, 58(1):153– 188, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dong:2022:HHO

[DE22] Zhaonan Dong and Alexandre Ern. Hybrid high-order and weak Galerkin methods for the biharmonic problem. SIAM Journal on Numerical Analysis, 60(5):2626–2656, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1408555.

Dolz:2024:UQE

[DE24] Jürgen Dölz and David Ebert.

On uncertainty quantification of eigenvalues and eigenspaces with higher multiplicity. SIAM Journal on Numerical Analysis, 62(1):422–451, February 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ding:2021:DLR

[DEL21] Zhiyan Ding, Lukas Einkemmer, and Qin Li. Dynamical low-rank integrator for the linear Boltzmann equation: Error analysis in the diffusion limit. SIAM Journal on Numerical Analysis, 59(4):2254– 2285, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Desiderio:2022:CCV

[DFFS22] Luca Desiderio, Silvia Falletta, Matteo Ferrari, and Letizia Scuderi. On the coupling of the curved virtual element method with the one-equation boundary element method for 2D exterior Helmholtz problems. SIAM Journal on Numerical Analysis, 60(4):2099–2124, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1460776.

Ditkowski:2020:TDE

[DGG20] Adi Ditkowski, Sigal Gottlieb, and Zachary J. Grant. Two-derivative error inhibiting schemes and enhanced error inhibiting schemes. SIAM Journal on Numerical Analysis, 58(6):3197–3225, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duran:2020:AFE

[DGL20] Ricardo Durán, Lucia Gastaldi, and Ariel Lombardi. Analysis of finite element approximations of Stokes equations with NonSmooth data. SIAM Journal on Numerical Analysis, 58 (6):3309–3331, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Du:2022:DDL

[DGYZ22] Qiang Du, Yiqi Gu, Haizhao Yang, and Chao Zhou. The discovery of dynamics via linear multistep methods and deep learning: Error estimation. SIAM Journal on Numerical Analysis, 60 (4):2014–2045, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M140691X.

Despres:2020:CSS

[DH20] Bruno Després and Maxime Herda. Computation of sum of squares polynomials from data points. SIAM Journal on Numerical Analysis, 58(3): 1719–1743, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Delourme:2021:CHB

[DH21] Bérangère Delourme and Laurence Halpern. A complex homographic best approximation problem. application to optimized Robin–Schwarz algorithms, and optimal control problems. SIAM Journal on Numerical Analysis, 59(3): 1769–1810, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dorich:2022:EIQ

[DH22] Benjamin Dörich and Marlis Hochbruck. Exponential integrators for quasilinear wavetype equations. SIAM Journal on Numerical Analysis, 60(3):1472–1493, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1410579.

Dorich:2024:EBD

[DH24] Benjamin Dörich and Patrick Henning. Error bounds for discrete minimizers of the Ginzburg–Landau energy in the high-κ regime. SIAM Journal on Numerical Analysis, 62(3):1313–1343, June 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1560938.

Dong:2023:IHO

[DHM23] Zhaonan Dong, Moritz Hauck, and Roland Maier. An improved high-order method for elliptic multiscale problems. SIAM Journal on Numerical Analysis, 61(4):1918–1937, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M153392X.

DeCaria:2024:APD

[DHS24] Victor P. DeCaria, Cory D. Hauck, and Stefan R. Schnake.

An asymptotic preserving discontinuous Galerkin method for a linear Boltzmann semiconductor model. SIAM Journal on Numerical Analysis, 62 (3):1067–1097, May 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

DeCaria:2020:ACR

[DIL^{+}20] Victor DeCaria, Traian Iliescu, William Layton, Michael McLaughlin, and Michael Schneier. An artificial compression reduced order model. SIAM Journal on Numerical Analysis, 58(1):565–589, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duc:2023:NAR

[DK23] Luu Hoang Duc and Peter Kloeden. Numerical attractors for rough differential equations. SIAM Journal on Numerical Analysis, 61(5):2381– 2407, October 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Droniou:2020:GDM

[DL20a] Jérôme Droniou and Kim-Ngan Le. The gradient discretization method for slow and fast diffusion porous media equations. SIAM Journal on Numerical Analysis, 58(3): 1965–1992, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Duprez:2020:FFE

[DL20b] Michel Duprez and Alexei Lozinski. ϕ-FEM: a finite element method on domains defined by level-sets. SIAM Journal on Numerical Analysis, 58 (2):1008–1028, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dautov:2020:EEL

[DLZ20] R. Z. Dautov, A. V. Lapin, and S. Zhang. Error estimates for Lagrange–Galerkin approximation of American options valuation. SIAM Journal on Numerical Analysis, 58(1):48– 65, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

DeRyck:2024:WWP

[DMM24] Tim De Ryck, Siddhartha Mishra, and Roberto Molinaro. wPINNs: Weak physics informed neural networks for approximating entropy solutions of hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 62(2): 811–841, March 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dong:2021:RBP

[DMS21] Zhaonan Dong, Lorenzo Mascotto, and Oliver J. Sutton. Residual-based a posteriori error estimates for hpdiscontinuous Galerkin discretizations of the biharmonic problem. SIAM Journal on Numerical Analysis, 59(3): 1273–1298, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Deckelnick:2021:EAF

[DN21] Klaus Deckelnick and Robert Nürnberg. Error analysis for a finite difference scheme for axisymmetric mean curvature flow of genus-0 surfaces. SIAM Journal on Numerical Analysis, 59(5):2698– 2721, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Deckelnick:2023:DHC

[DN23] Klaus Deckelnick and Robert Nürnberg. Discrete hyperbolic curvature flow in the plane. SIAM Journal on Numerical Analysis, 61(4):1835–1857, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1493112.

Demlow:2024:TPF

[DN24] Alan Demlow and Michael Neilan. A tangential and penalty-free finite element method for the surface Stokes problem. SIAM Journal on Numerical Analysis, 62(1): 248–272, January 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Despres:2022:OTC

[DNT22] B. Després, A. Nicolopoulos, and B. Thierry. Optimized transmission conditions in domain decomposition methods with cross-points for Helmholtz equation. SIAM Journal on Numerical Analysis, 60(5):2482–2507, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1421210.

Diening:2024:PGE

[DRS24] Lars Diening, Julian Rolfes, and Abner J. Salgado. Pointwise gradient estimate of the Ritz projection. SIAM Journal on Numerical Analysis, 62 (3):1212–1225, May 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dolejsi:2021:PEE

[DRV21] Vít Dolejsí, Filip Roskovec, and Miloslav Vlasák. A posteriori error estimates for higher order space-time Galerkin discretizations of nonlinear parabolic problems. SIAM Journal on Numerical Analysis, 59(3):1486–1509, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Du:2020:UEA

[DS20] Shukai Du and Francisco-Javier Sayas. A unified error analysis of hybridizable discontinuous Galerkin methods for the static Maxwell equations. SIAM Journal on Numerical Analysis, 58(2):1367– 1391, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dozsa:2024:BM

[DSS24] Tamás Dózsa, Ferenc Schipp, and Alexandros Soumelidis. On Bernoulli’s method. SIAM Journal on Numerical Analysis, 62(3):1259–1277, May 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Diening:2021:SSP

[DST21] Lars Diening, Johannes Storn, and Tabea Tscherpel. On the Sobolev and L^{p}-stability of the L^{2}-projection. SIAM Journal on Numerical Analysis, 59(5): 2571–2607, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dong:2021:AND

[DSY21] Bin Dong, Zuowei Shen, and Jianbin Yang. Approximation from noisy data. SIAM Journal on Numerical Analysis, 59 (5):2722–2745, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

daVeiga:2023:AVS

[dVCN^{+}23] L. Beirão da Veiga, C. Canuto, R. H. Nochetto, G. Vacca, and M. Verani. Adaptive VEM: Stabilization-free a posteriori error analysis and contraction property. SIAM Journal on Numerical Analysis, 61(2):457–494, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458740.

daVeiga:2024:RFE

[dVDV24] L. Beirão da Veiga, F. Dassi, and G. Vacca. Robust finite elements for linearized magnetohydrodynamics. SIAM Journal on Numerical Analysis, 62(4):1539–1564, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1582783.

Duru:2022:CES

[DWW22] Kenneth Duru, Siyang Wang, and Kenny Wiratama. A conservative and energy stable discontinuous spectral element method for the shifted wave equation in second order form. SIAM Journal on Numerical Analysis, 60(4):1631–1664, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1432922.

Du:2022:CLL

[DXY22] Qiang Du, Hehu Xie, and Xiaobo Yin. On the convergence to local limit of nonlocal models with approximated interaction neighborhoods. SIAM Journal on Numerical Analysis, 60(4):2046–2068, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1448227.

Eigel:2022:CAS

[EEST22] Martin Eigel, Oliver G. Ernst, Björn Sprungk, and Lorenzo Tamellini. On the convergence of adaptive stochastic collocation for elliptic partial differential equations with affine diffusion. SIAM Journal on Numerical Analysis, 60(2):659–687, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1364722.

Esedoglu:2022:MSO

[EG22] Selim Esedoğlu and Jiajia Guo. A monotone, second order accurate scheme for curvature motion. SIAM Journal on Numerical Analysis, 60(5):2435–2447, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466050.

Ern:2023:DGA

[EG23] Alexandre Ern and Jean-Luc Guermond. The discontinuous Galerkin approximation of the grad–div and curl— curl operators in first-order form is involution-preserving and spectrally correct. SIAM Journal on Numerical Analysis, 61(6):2940–2966, December 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Engstrom:2022:CAN

[EH22] Emil Engström and Eskil Hansen. Convergence analysis of the nonoverlapping Robin– Robin method for nonlinear elliptic equations. SIAM Journal on Numerical Analysis, 60(2):585–605, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1414942.

Einkemmer:2024:APE

[EHK24] Lukas Einkemmer, Jingwei Hu, and Jonas Kusch. Asymptoticpreserving and energy stable dynamical low-rank approximation. SIAM Journal on Numerical Analysis, 62(1):73– 92, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Evans:2020:PAA

[EPRX20] Claire Evans, Sara Pollock, Leo G. Rebholz, and Mengying Xiao. A proof that Anderson acceleration improves the convergence rate in linearly converging fixed-point methods (but not in those converging quadratically). SIAM Journal on Numerical Analysis, 58 (1):788–810, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Egger:2020:SOM

[ER20] Herbert Egger and Bogdan Radu. On a second-order multipoint flux mixed finite element methods on hybrid meshes. SIAM Journal on Numerical Analysis, 58(3):1822– 1844, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Egger:2021:SOF

[ER21] Herbert Egger and Bogdan Radu. A second-order finite element method with mass lumping for Maxwell’s equations on tetrahedra. SIAM Journal on Numerical Analysis, 59(2):864–885, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Eriksson:2021:SPA

[EW21] Sofia Eriksson and Siyang Wang. Summation-by-parts approximations of the second derivative: Pseudoinverse and revisitation of a high order accurate operator. SIAM Journal on Numerical Analysis, 59 (5):2669–2697, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fairag:2020:TLG

[FA20] Faisal A. Fairag and Johnson D. Audu. Two-level Galerkin mixed finite element method for Darcy– Forchheimer model in porous media. SIAM Journal on Numerical Analysis, 58(1):234– 253, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Feppon:2022:AMC

[FA22] Florian Feppon and Habib Ammari. Analysis of a Monte– Carlo Nystrom method. SIAM Journal on Numerical Analysis, 60(3):1226–1250, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1432338.

Foster:2024:HOS

[FdRS24] James M. Foster, Gonçalo dos Reis, and Calum Strange. High order splitting methods for SDEs satisfying a commutativity condition. SIAM Journal on Numerical Analysis, 62(1): 500–532, February 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fitzpatrick:2021:CME

[FDvW21] Brian Fitzpatrick, Enzo De Sena, and Toon van Waterschoot. On the convergence of the multipole expansion method. SIAM Journal on Numerical Analysis, 59(5):2473– 2499, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fang:2023:CRP

[FGK23] Zhou Fang, Ankit Gupta, and Mustafa Khammash. Convergence of regularized particle filters for stochastic reaction networks. SIAM Journal on Numerical Analysis, 61(2):399–430, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1453025.

Farrell:2020:NAU

[FGOS20] P. E. Farrell, P. A. Gazca-Orozco, and E. Süli. Numerical analysis of unsteady implicitly constituted incompressible fluids: 3-field formulation. SIAM Journal on Numerical Analysis, 58(1):757– 787, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fischer:2021:PEA

[FGP21] Julian Fischer, Dietmar Gallistl, and Daniel Peterseim. A priori error analysis of a numerical stochastic homogenization method. SIAM Journal on Numerical Analysis, 59 (2):660–674, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fuhrer:2024:RDT

[FH24] Thomas Führer and Norbert Heuer. Robust DPG test spaces and Fortin operators — the H^{1} and h(div) cases. SIAM Journal on Numerical Analysis, 62(2):718–748, March 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fuhrer:2022:MSO

[FHK22] Thomas Führer, Norbert Heuer, and Michael Karkulik. MINRES for second-order PDEs with singular data. SIAM Journal on Numerical Analysis, 60(3):1111–1135, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1457023.

Fuhrer:2023:DPG

[FHN23] Thomas Führer, Norbert Heuer, and Antti H. Niemi. A discontinuous Petrov–Galerkin method for Reissner–Mindlin plates. SIAM Journal on Numerical Analysis, 61(2):995– 1017, ???? 2023. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1498838.

Feng:2021:EEU

[FHW21] Xinlong Feng, Xueling Huang, and Kun Wang. Error estimate of unconditionally stable and decoupled linear positivity-preserving FEM for the chemotaxis-Stokes equations. SIAM Journal on Numerical Analysis, 59(6):3052– 3076, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fu:2022:MDC

[FK22] Guosheng Fu and Wenzheng Kuang. A monolithic divergence-conforming HDG scheme for a linear fluidstructure interaction model. SIAM Journal on Numerical Analysis, 60(2):631–658, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1385950.

Faustmann:2022:LCF

[FKM22] Markus Faustmann, Michael Karkulik, and Jens Markus Melenk. Local convergence of the FEM for the integral fractional Laplacian. SIAM Journal on Numerical Analysis, 60(3):1055–1082, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1343853.

Feng:2021:NSF

[FL21] Xiaobing Feng and Thomas Lewis. A narrow-stencil finite difference method for approximating viscosity solutions of Hamilton–Jacobi–Bellman equations. SIAM Journal on Numerical Analysis, 59(2): 886–924, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Feng:2023:DFE

[FLH23] Xinlong Feng, Xiaoli Lu, and Yinnian He. Difference finite element method for the 3D steady Navier– Stokes equations. SIAM Journal on Numerical Analysis, 61(1):167–193, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1450872.

Feng:2021:HOM

[FLM21] Xiaobing Feng, Buyang Li, and Shu Ma. High-order massand energy-conserving SAV– Gauss collocation finite element methods for the nonlinear Schrödinger equation. SIAM Journal on Numerical Analysis, 59(3):1566–1591, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Foster:2020:OPA

[FLO20] James Foster, Terry Lyons, and Harald Oberhauser. An optimal polynomial approximation of Brownian motion. SIAM Journal on Numerical Analysis, 58(3):1393–1421, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fischer:2021:WTP

[FM21] Julian Fischer and Daniel Matthes. The waiting time phenomenon in spatially discretized porous medium and thin film equations. SIAM Journal on Numerical Analysis, 59(1):60–87, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Faustmann:2023:ECH

[FMMS23] Markus Faustmann, Carlo Marcati, Jens Markus Melenk, and Christoph Schwab. Exponential convergence of hp-FEM for the integral fractional Laplacian in polygons. SIAM Journal on Numerical Analysis, 61(6):2601–2622, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fjordholm:2022:WPC

[FMR22] Ulrik S. Fjordholm, Markus Musch, and Nils H. Risebro. Well-posedness and convergence of a finite volume method for conservation laws on networks. SIAM Journal on Numerical Analysis, 60(2):606–630, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M145001X.

Feischl:2020:SCE

[FP20] Michael Feischl and Daniel Peterseim. Sparse compression of expected solution operators. SIAM Journal on Numerical Analysis, 58(6):3144– 3164, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fjordholm:2021:SOA

[FR21] Ulrik S. Fjordholm and Adrian M. Ruf. Secondorder accurate TVD numerical methods for nonlocal nonlinear conservation laws. SIAM Journal on Numerical Analysis, 59 (3):1167–1194, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fang:2023:OEB

[FV23] Di Fang and Albert Tres Vilanova. Observable error bounds of the time-splitting scheme for quantum-classical molecular dynamics. SIAM Journal on Numerical Analysis, 61(1):26–44, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462349.

Garcia-Archilla:2023:PRI

[GAJN23] Bosco García-Archilla, Volker John, and Julia Novo. POD-ROMs for incompressible flows including snapshots of the temporal derivative of the full order solution. SIAM Journal on Numerical Analysis, 61(3):1340–1368, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1503853.

Gallistl:2021:PEA

[Gal21] Dietmar Gallistl. A posteriori error analysis of the infsup constant for the divergence. SIAM Journal on Numerical Analysis, 59(1):249– 264, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Garcia-Archilla:2020:UTE

[GANT20] Bosco García-Archilla, Julia Novo, and Edriss S. Titi. Uniform in time error estimates for a finite element method applied to a downscaling data assimilation algorithm for the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 58(1): 410–429, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gawlik:2020:HOA

[Gaw20] Evan S. Gawlik. High-order approximation of Gaussian curvature with Regge finite elements. SIAM Journal on Numerical Analysis, 58(3):1801– 1821, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guastavino:2020:CRS

[GB20] Sabrina Guastavino and Federico Benvenuto. Convergence rates of spectral regularization methods: a comparison between ill-posed inverse problems and statistical kernel learning. SIAM Journal on Numerical Analysis, 58(6): 3504–3529, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Georgoulis:2021:HCF

[Geo21] Emmanuil H. Georgoulis. Hypocoercivity-compatible finite element methods for the long-time computation of Kolmogorov’s equation. SIAM Journal on Numerical Analysis, 59(1):173–194, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gao:2020:ORP

[GFSZ20] Wenwu Gao, Gregory E. Fasshauer, Xingping Sun, and Xuan Zhou. Optimality and regularization properties of quasi-interpolation: Deterministic and stochastic approaches. SIAM Journal on Numerical Analysis, 58(4): 2059–2078, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guillen-Gonzalez:2023:CTD

[GGC23] Francisco Guillén-González and André Luiz Corrêa Vianna Filho. Convergence of a time discrete scheme for a chemotaxis-consumption model. SIAM Journal on Numerical Analysis, 61(5):2509– 2533, October 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gottlieb:2022:HOS

[GGHS22] Sigal Gottlieb, Zachary J.

Grant, Jingwei Hu, and Ruiwen Shu. High order strong stability preserving Multi-Derivative implicit and IMEX Runge–Kutta methods with asymptotic preserving properties. SIAM Journal on Numerical Analysis, 60(1):423–449, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1403175.

Gedicke:2021:PDR

[GGPS21] Joscha Gedicke, Sjoerd Geevers, Ilaria Perugia, and Joachim Schöberl. A polynomialdegree-robust a posteriori error estimator for Nédélec discretizations of magnetostatic problems. SIAM Journal on Numerical Analysis, 59(4): 2237–2253, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guo:2024:IWN

[GH24] Hongxia Guo and Guanghui Hu. Inverse wave-number-dependent source problems for the Helmholtz equation. SIAM Journal on Numerical Analysis, 62(3):1372–1393, June 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1572696.

Gu:2023:SDE

[GHLY23] Yiqi Gu, John Harlim, Senwei Liang, and Haizhao Yang. Stationary density estimation of Itô diffusions using deep learning. SIAM Journal on Numerical Analysis, 61(1):45–82, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1445363.

Griebel:2020:MQE

[GHM20] Michael Griebel, Helmut Harbrecht, and Michael D. Multerer. Multilevel quadrature for elliptic parametric partial differential equations in case of polygonal approximations of curved domains. SIAM Journal on Numerical Analysis, 58 (1):684–705, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gander:2021:MAC

[GHM21] Martin J. Gander, Julian Hennicker, and Roland Masson. Modeling and analysis of the coupling in discrete fracture matrix models. SIAM Journal on Numerical Analysis, 59 (1):195–218, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guth:2024:GDT

[GK24] Philipp A. Guth and Vesa Kaarnioja. Generalized dimension truncation error analysis for high-dimensional numerical integration: Lognormal setting and beyond. SIAM Journal on Numerical Analysis, 62(2): 872–892, March 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Grekas:2022:AEM

[GKMR22] Georgios Grekas, Konstantinos Koumatos, Charalambos Makridakis, and Phoebus Rosakis. Approximations of energy minimization in cell-induced phase transitions of fibrous biomaterials: Γ-convergence analysis. SIAM Journal on Numerical Analysis, 60(2):715–750, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M137286X.

Gjerde:2021:MAP

[GKN21] Ingeborg G. Gjerde, Kundan Kumar, and Jan M. Nordbotten. A mixed approach to the Poisson problem with line sources. SIAM Journal on Numerical Analysis, 59(2):1117– 1139, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gilbert:2023:APF

[GKS23] Alexander D. Gilbert, Frances Y.

Kuo, and Ian H. Sloan. Analysis of preintegration followed by quasi-Monte Carlo integration for distribution functions and densities. SIAM Journal on Numerical Analysis, 61(1):135–166, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M146658X.

Gander:2024:NTD

[GL24] Martin J. Gander and Liu-Di Lu. New time domain decomposition methods for parabolic optimal control problems I: Dirichlet– Neumann and Neumann– Dirichlet algorithms. SIAM Journal on Numerical Analysis, 62(4):2048–2070, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1584502.

Glaubitz:2020:SHO

[Gla20] Jan Glaubitz. Stable high order quadrature rules for scattered data and general weight functions. SIAM Journal on Numerical Analysis, 58(4): 2144–2164, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gopalakrishnan:2020:MCM

[GLS20] Jay Gopalakrishnan, Philip L. Lederer, and Joachim Schöberl.

A mass conserving mixed stress formulation for Stokes flow with weakly imposed stress symmetry. SIAM Journal on Numerical Analysis, 58 (1):706–732, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guglielmi:2023:RMD

[GLS23] Nicola Guglielmi, Christian Lubich, and Stefano Sicilia. Rank-1 matrix differential equations for structured eigenvalue optimization. SIAM Journal on Numerical Analysis, 61(4):1737–1762, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1498735.

Gui:2022:CRF

[GLW22] Xinping Gui, Buyang Li, and Jilu Wang. Convergence of renormalized finite element methods for heat flow of harmonic maps. SIAM Journal on Numerical Analysis, 60(1):312–338, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1402212.

Gomez:2022:STT

[GM22] Sergio Gómez and Andrea Moiola. A space-time Trefftz discontinuous Galerkin method for the linear Schrödinger equation. SIAM Journal on Numerical Analysis, 60 (2):688–714, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1426079.

Gallistl:2024:LIT

[GM24] Dietmar Gallistl and Roland Maier. Localized implicit time stepping for the wave equation. SIAM Journal on Numerical Analysis, 62 (4):1589–1608, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1582618.

Gomez:2024:STV

[GMMP24] Sergio Gomez, Lorenzo Mascotto, Andrea Moiola, and Ilaria Perugia. Space-time virtual elements for the heat equation. SIAM Journal on Numerical Analysis, 62(1): 199–228, January 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gudi:2022:HHO

[GMP22] Thirupathi Gudi, Gouranga Mallik, and Tamal Pramanick. A hybrid high-order method for quasilinear elliptic problems of nonmonotone type. SIAM Journal on Numerical Analysis, 60(4):2318–2344, ???? 2022. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1412050.

Gong:2022:AAD

[GMSZ22] Wei Gong, Mariano Mateos, John Singler, and Yangwen Zhang. Analysis and approximations of Dirichlet boundary control of Stokes flows in the energy space. SIAM Journal on Numerical Analysis, 60(1):450–474, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1406799.

Glaubitz:2023:SPO

[GNÖ23] Jan Glaubitz, Jan Nordström, and Philipp Öffner. Summation-by-parts operators for general function spaces. SIAM Journal on Numerical Analysis, 61(2):733–754, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1470141.

Gergelits:2020:GSS

[GNS20] Tomás Gergelits, Bjørn Fredrik Nielsen, and Zdenek Strakos. Generalized spectrum of second order differential operators. SIAM Journal on Numerical Analysis, 58(4):2193– 2211, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gutleb:2020:SSM

[GO20] Timon S. Gutleb and Sheehan Olver. A sparse spectral method for Volterra integral equations using orthogonal polynomials on the triangle. SIAM Journal on Numerical Analysis, 58(3):1993–2018, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Glusa:2021:EEO

[GO21] Christian Glusa and Enrique Otárola. Error estimates for the optimal control of a parabolic fractional PDE. SIAM Journal on Numerical Analysis, 59(2):1140–1165, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gallistl:2023:CLB

[GO23] Dietmar Gallistl and Vladislav Olkhovskiy. Computational lower bounds of the Maxwell eigenvalues. SIAM Journal on Numerical Analysis, 61(2):539–561, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1461447.

Golse:2022:SRT

[GP22] François Golse and Olivier R. Pironneau. Stratified radiative transfer in a fluid and numerical applications to Earth science. SIAM Journal on Numerical Analysis, 60(5):2963–3000, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1459009.

Gonzalez-Pinto:2020:CNO

[GPHHA20] S. González-Pinto, E. Hairer, and D. Hernandez-Abreu. Convergence in ℓ_{2} and ℓ_{∞} norm of one-stage AMF-W-methods for parabolic problems. SIAM Journal on Numerical Analysis, 58(2):1117– 1137, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guermond:2020:PAP

[GPR20] Jean-Luc Guermond, Bojan Popov, and Jean Ragusa. Positive and asymptotic preserving approximation of the radiation transport equation. SIAM Journal on Numerical Analysis, 58(1):519–540, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gallistl:2021:THD

[GS21] Dietmar Gallistl and Mira Schedensack. Taylor–Hood discretization of the Reissner– Mindlin plate. SIAM Journal on Numerical Analysis, 59(3): 1195–1217, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gantner:2022:WPF

[GS22] Gregor Gantner and Rob Stevenson. A well-posed first order system least squares formulation of the instationary Stokes equations. SIAM Journal on Numerical Analysis, 60(3):1607–1629, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1432600.

Gao:2023:OAN

[GS23] Huadong Gao and Weiwei Sun. Optimal analysis of nonuniform Galerkin-mixed finite element approximations to the Ginzburg–Landau equations in superconductivity. SIAM Journal on Numerical Analysis, 61(2):929–951, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1483670.

Goda:2024:UMQ

[GSM24] Takashi Goda, Kosuke Suzuki, and Makoto Matsumoto. A universal median quasi-Monte Carlo integration. SIAM Journal on Numerical Analysis, 62 (1):533–566, February 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gong:2021:FEA

[GSW21] Bo Gong, Jiguang Sun, and Xinming Wu. Finite element approximation of the modified Maxwell’s Stekloff eigenvalues. SIAM Journal on Numerical Analysis, 59(5):2430– 2448, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Graham:2020:DDL

[GSZ20] Ivan G. Graham, Euan A.

Spence, and Jun Zou. Domain decomposition with local impedance conditions for the Helmholtz equation with absorption. SIAM Journal on Numerical Analysis, 58(5): 2515–2543, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Guo:2021:SPM

[Guo21] Ruchi Guo. Solving parabolic moving interface problems with dynamical immersed spaces on unfitted meshes: Fully discrete analysis. SIAM Journal on Numerical Analysis, 59(2):797–828, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gosse:2020:TTD

[GV20] Laurent Gosse and Nicolas Vauchelet. A truly two-dimensional, asymptoticpreserving scheme for a discrete model of radiative transfer. SIAM Journal on Numerical Analysis, 58(2):1092–1116, ???? 2020. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Gomez-Vargas:2023:TSP

[GVMRBV23] Bryan Gómez-Vargas, Kent-André Mardal, Ricardo Ruiz-Baier, and Vegard Vinje. Twofold saddle-point formulation of Biot poroelasticity with stress-dependent diffusion. SIAM Journal on Numerical Analysis, 61 (3):1449–1481, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1449695.

Gabbard:2024:LGF

[GvR24] James Gabbard and Wim M. van Rees. Lattice Green’s functions for high-order finite difference stencils. SIAM Journal on Numerical Analysis, 62 (1):25–47, January 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gander:2020:DBP

[GW20] Martin J. Gander and Shu-Lin Wu. A diagonalizationbased parareal algorithm for dissipative and wave propagation problems. SIAM Journal on Numerical Analysis, 58(5): 2981–3009, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gudibanda:2022:CAF

[GWY22] Varun M. Gudibanda, Franziska Weber, and Yukun Yue. Convergence analysis of a fully discrete energy-stable numerical scheme for the Q-Tensor Flow of liquid crystals. SIAM Journal on Numerical Analysis, 60(4):2150–2181, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1383550.

Gong:2021:DSG

[GZ21] Wei Gong and Shengfeng Zhu.

On discrete shape gradients of boundary type for PDEconstrained shape optimization. SIAM Journal on Numerical Analysis, 59(3):1510– 1541, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Halla:2021:ARC

[Hal21] Martin Halla. Analysis of radial complex scaling methods: Scalar resonance problems. SIAM Journal on Numerical Analysis, 59(4):2054– 2074, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Halla:2022:RCS

[Hal22] Martin Halla. Radial complex scaling for anisotropic scalar resonance problems. SIAM Journal on Numerical Analysis, 60(5):2713–2730, ????

2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1455747.

Hanot:2023:AOP

[Han23] Marien-Lorenzo Hanot. An arbitrary order and pointwise divergence-free finite element scheme for the incompressible 3D Navier–Stokes equations. SIAM Journal on Numerical Analysis, 61(2):784–811, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443686.

Haji-Ali:2022:AMM

[HAST22] Abdul-Lateef Haji-Ali, Jonathan Spence, and Aretha L. Teckentrup. Adaptive multilevel Monte Carlo for probabilities. SIAM Journal on Numerical Analysis, 60(4):2125–2149, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1447064.

Huang:2024:ACC

[HD24] Kuang Huang and Qiang Du. Asymptotic compatibility of a class of numerical schemes for a nonlocal traffic flow model. SIAM Journal on Numerical Analysis, 62(3):1119–1144, May 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Helin:2024:LSA

[Hel24] Tapio Helin. Least squares approximations in linear statistical inverse learning problems. SIAM Journal on Numerical Analysis, 62(4):2025– 2047, August 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1538600.

He:2022:OMN

[HG22] Wen-Ming He and Hailong Guo. Optimal maximum norm estimates for virtual element methods. SIAM Journal on Numerical Analysis, 60(3):1251–1280, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1420186.

Han:2024:LHD

[HHL24] Jiequn Han, Ruimeng Hu, and Jihao Long. Learning highdimensional McKean–Vlasov forward–backward stochastic differential equations with general distribution dependence. SIAM Journal on Numerical Analysis, 62(1):1–24, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hiemstra:2020:TEM

[HHM^{+}20] René R. Hiemstra, Thomas J. R. Hughes, Carla Manni, Hendrik Speleers, and Deepesh Toshniwal. A Tchebycheffian extension of multidegree Bsplines: Algorithmic computation and properties. SIAM Journal on Numerical Analysis, 58(2):1138–1163, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Herremans:2023:RSR

[HHT23] Astrid Herremans, Daan Huybrechs, and Lloyd N. Trefethen. Resolution of singularities by rational functions. SIAM Journal on Numerical Analysis, 61(6):2580– 2600, November 2023. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hiptmair:2020:OOP

[HJHUT20] Ralf Hiptmair, Carlos Jerez-Hanckes, and Carolina Urzúa-Torres. Optimal operator preconditioning for Galerkin boundary element methods on 3-dimensional screens. SIAM Journal on Numerical Analysis, 58(1):834–857, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hutzenthaler:2020:MPA

[HK20] Martin Hutzenthaler and Thomas Kruse. Multilevel Picard approximations of high-dimensional semilinear parabolic differential equations with gradient-dependent nonlinearities. SIAM Journal on Numerical Analysis, 58(2): 929–961, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ha:2021:IAF

[HKYY21] Youngsoo Ha, Chang Ho Kim, Hyoseon Yang, and Jungho Yoon. Improving accuracy of the fifth-order WENO scheme by using the exponential approximation space. SIAM Journal on Numerical Analysis, 59(1):143–172, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Heltai:2023:AFE

[HL23] Luca Heltai and Wenyu Lei.

Adaptive finite element approximations for elliptic problems using regularized forcing data. SIAM Journal on Numerical Analysis, 61 (2):431–456, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1455991.

Hu:2022:CFE

[HLM22] Jun Hu, Yizhou Liang, and Rui Ma. Conforming finite element DIVDIV complexes and the application for the linearized Einstein–Bianchi system. SIAM Journal on Numerical Analysis, 60(3):1307–1330, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1404235.

Hairer:2023:LMR

[HLS23] Ernst Hairer, Christian Lubich, and Yanyan Shi. Leapfrog methods for relativistic chargedparticle dynamics. SIAM Journal on Numerical Analysis, 61(6):2844–2858, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huo:2021:ESP

[HLTW21] Xiaokai Huo, Hailiang Liu, Athanasios E. Tzavaras, and Shuaikun Wang. An energy stable and positivitypreserving scheme for the Maxwell–Stefan diffusion system. SIAM Journal on Numerical Analysis, 59(5):2321–2345, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Han:2022:SWE

[HM22] Bin Han and Michelle Michelle. Sharp wavenumber-explicit stability bounds for 2D Helmholtz equations. SIAM Journal on Numerical Analysis, 60(4):1985–2013, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1446186.

Hannukainen:2022:DSL

[HMO22] Antti Hannukainen, Jarmo Malinen, and Antti Ojalammi.

Distributed solution of Laplacian eigenvalue problems. SIAM Journal on Numerical Analysis, 60(1):76–103, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1342653.

Hong:2021:RNA

[HN21] Youngjoon Hong and David Nicholls. A rigorous numerical analysis of the transformed field expansion method for diffraction by periodic, layered structures. SIAM Journal on Numerical Analysis, 59 (1):456–476, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoyt:2020:MDR

[HO20] Christopher R. Hoyt and Art B. Owen. Mean dimension of ridge functions. SIAM Journal on Numerical Analysis, 58 (2):1195–1216, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoyt:2024:MDR

[HO24] Christopher Hoyt and Art B.

Owen. Mean dimension of radial basis functions. SIAM Journal on Numerical Analysis, 62(3):1191–1211, May 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Henning:2020:SGF

[HP20] Patrick Henning and Daniel Peterseim. Sobolev gradient flow for the Gross–Pitaevskii eigenvalue problem: Global convergence and computational efficiency. SIAM Journal on Numerical Analysis, 58 (3):1744–1772, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hure:2021:DNN

[HPBL21] Côme Huré, Huyên Pham, Achref Bachouch, and Nicolas Langrené. Deep neural networks algorithms for stochastic control problems on finite horizon: Convergence analysis. SIAM Journal on Numerical Analysis, 59(1):525–557, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hu:2021:NSC

[HQY21] Jingwei Hu, Kunlun Qi, and Tong Yang. A new stability and convergence proof of the Fourier–Galerkin spectral method for the spatially homogeneous Boltzmann equation. SIAM Journal on Numerical Analysis, 59(2):613– 633, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hu:2020:NLS

[HS20] Qiya Hu and Rongrong Song. A novel least squares method for Helmholtz equations with large wave numbers. SIAM Journal on Numerical Analysis, 58(5):3091–3123, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huang:2021:SEA

[HS21] Fukeng Huang and Jie Shen.

Stability and error analysis of a class of high-order IMEX schemes for Navier– Stokes equations with periodic boundary conditions. SIAM Journal on Numerical Analysis, 59(6):2926–2954, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huang:2023:SEA

[HS23] Fukeng Huang and Jie Shen.

Stability and error analysis of a second-order consistent splitting scheme for the Navier– Stokes equations. SIAM Journal on Numerical Analysis, 61 (5):2408–2433, October 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huang:2024:NCB

[HS24] Fukeng Huang and Jie Shen.

On a new class of BDF and IMEX schemes for parabolic type equations. SIAM Journal on Numerical Analysis, 62 (4):1609–1637, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1612986.

Horning:2020:FDE

[HT20] Andrew Horning and Alex Townsend. FEAST for differential eigenvalue problems. SIAM Journal on Numerical Analysis, 58(2):1239–1262, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hu:2021:CHX

[Hu21] Qiya Hu. Convergence of the Hiptmair–Xu preconditioner for Maxwell’s equations with jump coefficients (i): Extensions of the regular decomposition. SIAM Journal on Numerical Analysis, 59(5):2500– 2535, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hu:2023:CHX

[Hu23] Qiya Hu. Convergence of the Hiptmair–Xu preconditioner for H(curl)-elliptic problems with jump coefficients (ii): Main results. SIAM Journal on Numerical Analysis, 61(5): 2434–2459, October 2023. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Huybrechs:2022:CGQ

[Huy22] Daan Huybrechs. On the computation of Gaussian quadrature rules for Chebyshev sets of linearly independent functions.

SIAM Journal on Numerical Analysis, 60(3):1168–1192, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1456935.

Hoff:2021:MMP

[HW21] Daniel Hoff and Holger Wendland. A meshfree method for a PDE-Constrained optimization problem. SIAM Journal on Numerical Analysis, 59(4): 1896–1917, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Han:2022:MDI

[HW22] Rubing Han and Shuonan Wu.

A monotone discretization for integral fractional Laplacian on bounded Lipschitz domains: Pointwise error estimates under Hölder regularity. SIAM Journal on Numerical Analysis, 60(6):3052–3077, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1448239.

Hong:2020:ALE

[HX20] Xue Hong and Yinhua Xia. Arbitrary Lagrangian–Eulerian discontinuous Galerkin method for hyperbolic equations involving δ-singularities. SIAM Journal on Numerical Analysis, 58(1):125–152, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hensel:2022:NAM

[HY22] Maurice Hensel and Irwin Yousept. Numerical analysis for Maxwell obstacle problems in electric shielding. SIAM Journal on Numerical Analysis, 60(3):1083–1110, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1427693.

Hao:2020:ORE

[HZ20] Zhaopeng Hao and Zhongqiang Zhang. Optimal regularity and error estimates of a spectral Galerkin method for fractional advection–diffusion– reaction equations. SIAM Journal on Numerical Analysis, 58(1):211–233, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

He:2023:ERI

[HZW23] Zhijian He, Zhan Zheng, and Xiaoqun Wang. On the error rate of importance sampling with randomized quasi-Monte Carlo. SIAM Journal on Numerical Analysis, 61(2):515–538, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1510121.

Hu:2022:FFE

[HZZ22] Kaibo Hu, Qian Zhang, and Zhimin Zhang. A family of finite element Stokes complexes in three dimensions. SIAM Journal on Numerical Analysis, 60(1):222–243, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1358700.

Imbert-Gerard:2021:ABG

[IG21]  Lise-Marie  Imbert-Gerard.

Amplitude-based generalized plane waves: New quasi-Trefftz functions for scalar equations in two dimensions. SIAM Journal on Numerical Analysis, 59(3):1663–1686, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Izgin:2022:SUP

[IKM22] Thomas Izgin, Stefan Kopecz, and Andreas Meister. On the stability of unconditionally positive and linear invariants preserving time integration schemes. SIAM Journal on Numerical Analysis, 60(6):3029–3051, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1480318.

Joly:2022:LTB

[JK22] Patrick Joly and Maryna Kachanovska. Local transparent boundary conditions for wave propagation in fractal trees (II). Error and complexity analysis. SIAM Journal on Numerical Analysis, 60(2):529–557, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1357524.

Jessberger:2024:FED

[JK24] Julius Jeßberger and Alex Kaltenbach. Finite element discretization of the steady, generalized Navier– Stokes equations with inhomogeneous Dirichlet boundary conditions. SIAM Journal on Numerical Analysis, 62 (4):1660–1686, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1607398.

Jayadharan:2023:STM

[JKVY23] Manu Jayadharan, Michel Kern, Martin Vohralík, and Ivan Yotov. A space-time multiscale mortar mixed finite element method for parabolic equations. SIAM Journal on Numerical Analysis, 61 (2):675–706, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1447945.

Jin:2021:CRB

[JLL21] Shi Jin, Lei Li, and Jian-Guo Liu. Convergence of the random batch method for interacting particles with disparate species and weights. SIAM Journal on Numerical Analysis, 59(2):746–768, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jia:2022:SVG

[JLM22] Junxiong Jia, Peijun Li, and Deyu Meng. Stein variational gradient descent on infinitedimensional space and applications to statistical inverse problems. SIAM Journal on Numerical Analysis, 60(4):2225–2252, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1440773.

Ju:2022:GSE

[JLQ22] Lili Ju, Xiao Li, and Zhonghua Qiao. Generalized SAVexponential integrator schemes for Allen–Cahn type gradient flows. SIAM Journal on Numerical Analysis, 60 (4):1905–1931, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1446496.

Jiang:2022:FEM

[JLWZ22] Run Jiang, Yonglin Li, Haijun Wu, and Jun Zou. Finite element method for a nonlinear perfectly matched layer Helmholtz equation with high wave number. SIAM Journal on Numerical Analysis, 60(5):2866–2896, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1459381.

Jiang:2021:ACC

[JLY21] Nan Jiang, Ying Li, and Huanhuan Yang. An artificial compressibility Crank– Nicolson leap-frog method for the Stokes–Darcy model and application in ensemble simulations. SIAM Journal on Numerical Analysis, 59(1):401– 428, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jiang:2024:NMA

[JLZ24] Kai Jiang, Shifeng Li, and Pingwen Zhang. Numerical methods and analysis of computing quasiperiodic systems. SIAM Journal on Numerical Analysis, 62(1):353– 375, February 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jiang:2023:CPP

[JSZ23] Wei Jiang, Chunmei Su, and Ganghui Zhang. A convexitypreserving and perimeterdecreasing parametric finite element method for the areapreserving curve shortening flow. SIAM Journal on Numerical Analysis, 61 (4):1989–2010, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1514404.

Jin:2021:EAF

[JZ21a] Bangti Jin and Zhi Zhou. Error analysis of finite element approximations of diffusion coefficient identification for elliptic and parabolic problems. SIAM Journal on Numerical Analysis, 59(1):119–142, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jungel:2021:CSP

[JZ21b] Ansgar Jüngel and Antoine Zurek. A convergent structure-preserving finitevolume scheme for the Shigesada– Kawasaki–Teramoto population system. SIAM Journal on Numerical Analysis, 59(4): 2286–2309, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jiang:2024:ARG

[JZZ24] Kai Jiang, Qi Zhou, and Pingwen Zhang. Accurately recover global quasiperiodic systems by finite points. SIAM Journal on Numerical Analysis, 62(4):1713–1735, August 2024. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1620247.

Kamenski:2021:SBS

[Kam21] Lennard Kamenski. Sharp bounds on the smallest eigenvalue of finite element equations with arbitrary meshes without regularity assumptions. SIAM Journal on Numerical Analysis, 59(2):983– 997, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Karaa:2021:PDT

[Kar21] Samir Karaa. Positivity of discrete time-fractional operators with applications to phasefield equations. SIAM Journal on Numerical Analysis, 59(4): 2040–2053, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Keller:2021:DDU

[KD21] Rachael T. Keller and Qiang Du. Discovery of dynamics using linear multistep methods. SIAM Journal on Numerical Analysis, 59(1):429–455, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Khan:2020:RPE

[KK20] Arbaz Khan and Guido Kanschat. A robust a posteriori error estimator for divergence-conforming discontinuous Galerkin methods for the Oseen equation. SIAM Journal on Numerical Analysis, 58(1):492–518, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kashiwabara:2024:DMR

[KK24] Takahito Kashiwabara and Tomoya Kemmochi. Discrete maximal regularity for the discontinuous Galerkin timestepping method without logarithmic factor. SIAM Journal on Numerical Analysis, 62 (4):1638–1659, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1580802.

Kaarnioja:2020:UQU

[KKS20] V. Kaarnioja, F. Y. Kuo, and I. H. Sloan. Uncertainty quantification using periodic random variables. SIAM Journal on Numerical Analysis, 58(2): 1068–1091, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kuchta:2021:AAM

[KLMZ21] Miroslav Kuchta, Federica Laurino, Kent-Andre Mardal, and Paolo Zunino. Analysis and approximation of mixeddimensional PDEs on 3D– 1D domains coupled with Lagrange multipliers. SIAM Journal on Numerical Analysis, 59(1):558–582, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kovacs:2020:MLE

[KLS20] Mihály Kovács, Stig Larsson, and Fardin Saedpanah. Mittag-Leffler Euler integrator for a stochastic fractional order equation with additive noise. SIAM Journal on Numerical Analysis, 58(1):66–85, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kopteva:2020:EAF

[KM20a] Natalia Kopteva and Xiangyun Meng. Error analysis for a fractional-derivative parabolic problem on quasigraded meshes using barrier functions. SIAM Journal on Numerical Analysis, 58(2): 1217–1238, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kyza:2020:PPE

[KM20b] Irene Kyza and Stephen Metcalfe. Pointwise a posteriori error bounds for blow-up in the semilinear heat equation. SIAM Journal on Numerical Analysis, 58(5):2609– 2631, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kazakova:2020:DTB

[KN20] Maria Kazakova and Pascal Noble. Discrete transparent boundary conditions for the linearized Green–Naghdi system of equations. SIAM Journal on Numerical Analysis, 58 (1):657–683, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kazashi:2023:DER

[KN23] Yoshihito Kazashi and Fabio Nobile. Density estimation in RKHS with application to Korobov spaces in high dimensions. SIAM Journal on Numerical Analysis, 61 (2):1080–1102, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M147476X.

Kao:2023:HFF

[KOO23] Chiu-Yen Kao, Braxton Osting, and Édouard Oudet. Harmonic functions on finitely connected tori. SIAM Journal on Numerical Analysis, 61(6): 2795–2812, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kopteva:2020:EAT

[Kop20] Natalia Kopteva. Error analysis for time-fractional semilinear parabolic equations using upper and lower solutions. SIAM Journal on Numerical Analysis, 58(4):2212– 2234, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kindermann:2020:CHP

[KR20] Stefan Kindermann and Kemal Raik. Convergence of heuristic parameter choice rules for convex Tikhonov regularization. SIAM Journal on Numerical Analysis, 58(3): 1773–1800, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kaltenbach:2023:LDGa

[KR23a] Alex Kaltenbach and Michael Ruzicka. A local discontinuous Galerkin approximation for the p-Navier–Stokes system, Part I: Convergence analysis. SIAM Journal on Numerical Analysis, 61(4):1613–1640, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M151474X.

Kaltenbach:2023:LDGb

[KR23b] Alex Kaltenbach and Michael Ruzicka. A local discontinuous Galerkin approximation for the p-Navier–Stokes system, Part II: Convergence rates for the velocity. SIAM Journal on Numerical Analysis, 61(4):1641–1663, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1514751.

Kaltenbach:2023:LDGc

[KR23c] Alex Kaltenbach and Michael Ruzicka. A local discontinuous Galerkin approximation for the p-Navier–Stokes system, Part III: Convergence rates for the pressure. SIAM Journal on Numerical Analysis, 61(4):1763–1782, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1541472.

Kopteva:2023:PPE

[KR23d] Natalia Kopteva and Richard Rankin. Pointwise a posteriori error estimates for discontinuous Galerkin methods for singularly perturbed reaction– diffusion equations. SIAM Journal on Numerical Analysis, 61(4):1938–1961, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M149733X.

Koc:2021:OPT

[KRS^{+}21] Birgul Koc, Samuele Rubino, Michael Schneier, John Singler, and Traian Iliescu. On optimal pointwise in time error bounds and difference quotients for the proper orthogonal decomposition. SIAM Journal on Numerical Analysis, 59(4):2163–2196, ????

2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kean:2020:EAS

[KS20] Kiera Kean and Michael Schneier. Error analysis of supremizer pressure recovery for POD based reduced-order models of the time-dependent Navier–Stokes equations. SIAM Journal on Numerical Analysis, 58(4): 2235–2264, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Kazashi:2023:SGH

[KSG23] Yoshihito Kazashi, Yuya Suzuki, and Takashi Goda. Suboptimality of Gauss–Hermite quadrature and optimality of the trapezoidal rule for functions with finite smoothness. SIAM Journal on Numerical Analysis, 61(3):1426–1448, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1480276.

Khieu:2022:SRB

[KV22] Tran Thi Khieu and Hoang-Hung Vo. Stability results for backward nonlinear diffusion equations with temporal coupling operator of local and nonlocal type. SIAM Journal on Numerical Analysis, 60(4):1665–1700, ???? 2022. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421763.

Kaliuzhnyi-Verbovetskyi:2022:SMC

[KVM22] Dmitry Kaliuzhnyi-Verbovetskyi and Georgi S. Medvedev. Sparse Monte Carlo method for nonlocal diffusion problems. SIAM Journal on Numerical Analysis, 60(6): 3001–3028, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1308657.

Komori:2023:FMM

[KYB23] Yoshio Komori, Guoguo Yang, and Kevin Burrage. Formulae for mixed moments of Wiener processes and a stochastic area integral. SIAM Journal on Numerical Analysis, 61(4):1716–1736, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M152013X.

Labovsky:2020:ADC

[Lab20] Alexander E. Labovsky. Approximate deconvolution with correction: a member of a new class of models for high Reynolds number flows. SIAM Journal on Numerical Analysis, 58(5):3068–3090, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2022:ASE

[LAZ22] Hao Li, Daniel Appelö, and Xiangxiong Zhang. Accuracy of spectral element method for wave, parabolic, and Schrödinger equations. SIAM Journal on Numerical Analysis, 60(1):339–363, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1401760.

Lin:2024:KIH

[LCS24] Shao-Bo Lin, Xiangyu Chang, and Xingping Sun. Kernel interpolation of high dimensional scattered data. SIAM Journal on Numerical Analysis, 62(3):1098–1118, May 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Du:2022:TSO

[lDzSW22] Rui lian Du, Zhi zhong Sun, and Hong Wang. Temporal second-order finite difference schemes for variableorder time-fractional wave equations. SIAM Journal on Numerical Analysis, 60 (1):104–132, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/19M1301230.

Lederer:2021:HHJ

[Led21]  Philip L. Lederer.  A

Hellan–Herrmann–Johnsonlike method for the stream function formulation of the Stokes equations in two and three space dimensions. SIAM Journal on Numerical Analysis, 59(1):503–524, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

LeRoux:2020:FAH

[LET20] Daniel Y. Le Roux, Christopher Eldred, and Mark A. Taylor. Fourier analyses of high-order continuous and discontinuous Galerkin methods. SIAM Journal on Numerical Analysis, 58(3):1845–1866, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2020:CDL

[Li20] Buyang Li. Convergence of Dziuk’s linearly implicit parametric finite element method for curve shortening flow. SIAM Journal on Numerical Analysis, 58(4):2315–2333, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2021:CDS

[Li21] Buyang Li. Convergence of Dziuk’s semidiscrete finite element method for mean curvature flow of closed surfaces with high-order finite elements. SIAM Journal on Numerical Analysis, 59(3):1592– 1617, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:2021:CMF

[LJL^{+}21] Siting Liu, Matthew Jacobs, Wuchen Li, Levon Nurbekyan, and Stanley J. Osher. Computational methods for first-order nonlocal mean field games with applications. SIAM Journal on Numerical Analysis, 59(5):2639–2668, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lai:2022:MSB

[LL22a] Ming-Jun Lai and Jinsil Lee. A multivariate spline based collocation method for numerical solution of partial differential equations. SIAM Journal on Numerical Analysis, 60(5):2405–2434, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469602.

Lou:2022:IUB

[LL22b] Yimin Lou and Christoph Lehrenfeld. Isoparametric unfitted BDF–Finite element method for PDEs on evolving domains. SIAM Journal on Numerical Analysis, 60(4):2069–2098, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M142126X.

Li:2022:SVS

[LlL22] Zhaoyi Li and Hong lin Liao. Stability of variablestep BDF2 and BDF3 methods. SIAM Journal on Numerical Analysis, 60(4):2253–2272, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462398.

Liao:2023:DGS

[lLLL23] Hong lin Liao, Nan Liu, and Pin Lyu. Discrete gradient structure of a second-order variable-step method for nonlinear integro-differential models. SIAM Journal on Numerical Analysis, 61(5):2157–2181, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1520050.

Li:2023:ESM

[LLMR23] Buyang Li, Yanping Lin, Shu Ma, and Qiqi Rao. An exponential spectral method using VP means for semilinear subdiffusion equations with rough data. SIAM Journal on Numerical Analysis, 61(5):2305–2326, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1512041.

Lu:2021:OFD

[LLS21] Jianfang Lu, Yong Liu, and Chi-Wang Shu. An oscillationfree discontinuous Galerkin method for scalar hyperbolic conservation laws. SIAM Journal on Numerical Analysis, 59 (3):1299–1324, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liao:2020:ESM

[lLTZ20] Hong lin Liao, Tao Tang, and Tao Zhou. On energy stable, maximum-principle preserving, second-order BDF scheme with variable steps for the Allen–Cahn equation. SIAM Journal on Numerical Analysis, 58(4):2294–2314, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:2022:NPF

[LLZZ22] Shu Liu, Wuchen Li, Hongyuan Zha, and Haomin Zhou. Neural parametric Fokker–Planck equation. SIAM Journal on Numerical Analysis, 60 (3):1385–1449, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1344986.

Li:2022:ECQ

[LM22a] Buyang Li and Shu Ma. Exponential convolution quadrature for nonlinear subdiffusion equations with nonsmooth initial data. SIAM Journal on Numerical Analysis, 60(2):503–528, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421386.

Li:2022:NMI

[LM22b] Shukai Li and Sanjay Mehrotra. Numerical methods for integral equations of the second kind with NonSmooth solutions of bounded variation. SIAM Journal on Numerical Analysis, 60(5):2751–2780, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1480422.

Lundquist:2023:MLF

[LMN23] Tomas Lundquist, Arnaud G.

Malan, and Jan Nordström. A method-of-lines framework for energy stable arbitrary Lagrangian–Eulerian methods. SIAM Journal on Numerical Analysis, 61(5):2327–2351, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1514945.

Liu:2023:CDS

[LMR23] Chen Liu, Rami Masri, and Beatrice Riviere. Convergence of a decoupled splitting scheme for the Cahn–Hilliard– Navier–Stokes system. SIAM Journal on Numerical Analysis, 61(6):2651–2694, November 2023. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2022:SIE

[LMS22] Buyang Li, Shu Ma, and Katharina Schratz. A semiimplicit exponential lowregularity integrator for the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 60(4):2273–2292, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1437007.

Leclerc:2020:LDC

[LMSS20] Hugo Leclerc, Quentin Mérigot, Filippo Santambrogio, and Federico Stra. Lagrangian discretization of crowd motion and linear diffusion. SIAM Journal on Numerical Analysis, 58(4):2093–2118, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liu:2023:PAS

[LO23] Sifan Liu and Art B. Owen. Preintegration via active subspace. SIAM Journal on Numerical Analysis, 61(2): 495–514, ???? 2023. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1479129.

Leimkuhler:2024:CCR

[LPW24] Benedict J. Leimkuhler, Daniel Paulin, and Peter A. Whalley.

Contraction and convergence rates for discretized kinetic Langevin dynamics. SIAM Journal on Numerical Analysis, 62(3):1226–1258, May 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lederer:2020:PRE

[LR20] Philip L. Lederer and Sander Rhebergen. A pressurerobust embedded discontinuous Galerkin method for the Stokes problem by reconstruction operators. SIAM Journal on Numerical Analysis, 58(5): 2915–2933, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lu:2022:AIO

[LRK22] Peipei Lu, Andreas Rupp, and Guido Kanschat. Analysis of injection operators in geometric multigrid solvers for HDG methods. SIAM Journal on Numerical Analysis, 60(4):2293–2317, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1400110.

Li:2020:EAS

[LS20a] Xiaoli Li and Jie Shen. Error analysis of the SAV–MAC scheme for the Navier–Stokes equations. SIAM Journal on Numerical Analysis, 58(5): 2465–2491, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Locke:2020:NPO

[LS20b] Sarah Locke and John Singler. New proper orthogonal decomposition approximation theory for PDE solution data. SIAM Journal on Numerical Analysis, 58(6):3251– 3285, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2022:CNS

[LS22] Yang Li and Bangwei She. On convergence of numerical solutions for the compressible MHD system with exactly divergence-free magnetic field. SIAM Journal on Numerical Analysis, 60(4):2182–2202, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1431011.

Loscher:2024:STF

[LS24] Richard Löscher and Olaf Steinbach. Space-time finite element methods for distributed optimal control of the wave equation. SIAM Journal on Numerical Analysis, 62(1): 452–475, February 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Langer:2021:STF

[LSTY21] Ulrich Langer, Olaf Steinbach, Fredi Tröltzsch, and Huidong Yang. Space-time finite element discretization of parabolic optimal control problems with energy regularization. SIAM Journal on Numerical Analysis, 59(2):675– 695, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2024:OEA

[LSXY24] Buyang Li, Weiwei Sun, Yupei Xie, and Wenshan Yu. Optimal L^{2} error analysis of a loosely coupled finite element scheme for thin-structure interactions. SIAM Journal on Numerical Analysis, 62 (4):1782–1813, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1578401.

Leng:2021:ACR

[LTTF21] Yu Leng, Xiaochuan Tian, Nathaniel Trask, and John T. Foster. Asymptotically compatible reproducing kernel collocation and meshfree integration for nonlocal diffusion. SIAM Journal on Numerical Analysis, 59(1):88–118, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2020:SOS

[LUZ20] Buyang Li, Yuki Ueda, and Guanyu Zhou. A secondorder stabilization method for linearizing and decoupling nonlinear parabolic systems. SIAM Journal on Numerical Analysis, 58(5):2736–2763, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2023:SHO

[LWC23] Shishun Li, Jing-Yuan Wang, and Xiao-Chuan Cai. Astable high-order block implicit methods for parabolic equations. SIAM Journal on Numerical Analysis, 61 (4):1858–1884, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M152880X.

Li:2022:SEA

[LWS22] Xiaoli Li, Weilong Wang, and Jie Shen. Stability and error analysis of IMEX SAV schemes for the magneto-hydrodynamic equations. SIAM Journal on Numerical Analysis, 60(3): 1026–1054, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1430376.

Liu:2022:CAV

[LWWW22] Chun Liu, Cheng Wang, Yiwei Wang, and Steven M. Wise. Convergence analysis of the variational operator splitting scheme for a reaction–diffusion system with detailed balance. SIAM Journal on Numerical Analysis, 60(2):781–803, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421283.

Li:2020:CLL

[LWX20] Buyang Li, Jilu Wang, and Liwei Xu. A convergent linearized Lagrange finite element method for the magneto-hydrodynamic equations in two-dimensional nonsmooth and nonconvex domains. SIAM Journal on Numerical Analysis, 58(1):430– 459, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Liang:2024:TLB

[LWX24] Qigang Liang, Wei Wang, and Xuejun Xu. A two-level block preconditioned Jacobi– Davidson method for multiple and clustered eigenvalues of elliptic operators. SIAM Journal on Numerical Analysis, 62 (2):998–1019, April 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lyu:2020:CAS

[LWXZ20] Junlong Lyu, Zhongjian Wang, Jack Xin, and Zhiwen Zhang. Convergence analysis of stochastic structure-preserving schemes for computing effective diffusivity in random flows. SIAM Journal on Numerical Analysis, 58(5):3040–3067, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lyu:2022:CIP

[LWXZ22] Junlong Lyu, Zhongjian Wang, Jack Xin, and Zhiwen Zhang. A convergent interacting particle method and computation of KPP front speeds in chaotic flows. SIAM Journal on Numerical Analysis, 60(3):1136–1167, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1410786.

Li:2023:MCG

[LWY23] Qin Li, Li Wang, and Yunan Yang. Monte Carlo gradient in optimization constrained by radiative transport equation. SIAM Journal on Numerical Analysis, 61(6):2744–2774, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2020:LTA

[LWZ20] Buyang Li, Kai Wang, and Zhi Zhou. Long-time accurate symmetrized implicit-explicit BDF methods for a class of parabolic equations with nonself-adjoint operators. SIAM Journal on Numerical Analysis, 58(1):189–210, ???? 2020.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2023:GTE

[LWZ23] Bing Li, Yifei Wu, and Xiaofei Zhao. Gauge-transformed exponential integrator for generalized KdV equations with rough data. SIAM Journal on Numerical Analysis, 61(4):1689–1715, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1530501.

Liang:2023:TLP

[LX23] Qigang Liang and Xuejun Xu. A two-level preconditioned Helmholtz subspace iterative method for Maxwell eigenvalue problems. SIAM Journal on Numerical Analysis, 61(2):642–674, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1392012.

Li:2020:SLS

[LY20] Ruo Li and Fanyi Yang. A sequential least squares method for Poisson equation using a patch reconstructed space. SIAM Journal on Numerical Analysis, 58(1):353–374, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Li:2023:OCN

[LYC23] Yangrong Li, Shuang Yang, and Tomás Caraballo. Optimization and convergence of numerical attractors for discrete-time quasi-linear lattice system. SIAM Journal on Numerical Analysis, 61(2):905–928, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1461642.

Liu:2024:FEM

[LYL24] Fuhao Liu, Wei Yang, and Jichun Li. A finite element method for hyperbolic metamaterials with applications for hyperlens. SIAM Journal on Numerical Analysis, 62(3):1420–1442, June 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1591207.

Li:2023:AAT

[LZ23] Jichun Li and Li Zhu. Analysis and application of two novel finite element methods for solving Ziolkowski’s PML model in the integro-differential form. SIAM Journal on Numerical Analysis, 61(5):2209–2236, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1506936.

Luo:2022:CAD

[LZCZ22] Yue Luo, Xiangcheng Zheng, Xiangle Cheng, and Lei Zhang. Convergence analysis of discrete high-index saddle dynamics. SIAM Journal on Numerical Analysis, 60 (5):2731–2750, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1487965.

Li:2020:DGM

[LZM^{+}20] Jia Li, Dazhi Zhang, Xiong Meng, Boying Wu, and Qiang Zhang. Discontinuous Galerkin methods for nonlinear scalar conservation laws: Generalized local Lax– Friedrichs numerical fluxes. SIAM Journal on Numerical Analysis, 58(1):1–20, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Maier:2021:HOA

[Mai21] Roland Maier. A high-order approach to elliptic multiscale problems with general unstructured coefficients. SIAM Journal on Numerical Analysis, 59 (2):1067–1089, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ma:2020:ECG

[MCL20] Chupeng Ma, Liqun Cao, and Yanping Lin. Energy conserving Galerkin finite element methods for the Maxwell– Klein–Gordon system. SIAM Journal on Numerical Analysis, 58(2):1339–1366, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Metzger:2021:ECF

[Met21]  Stefan Metzger.  An effi-

cient and convergent finite element scheme for Cahn– Hilliard equations with dynamic boundary conditions. SIAM Journal on Numerical Analysis, 59(1):219–248, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mlinaric:2023:UFI

[MG23] Petar Mlinarić and Serkan Gugercin. A unifying framework for interpolatory L2optimal reduced-order modeling. SIAM Journal on Numerical Analysis, 61(5):2133–2156, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1516920.

Manns:2020:MSR

[MK20] Paul Manns and Christian Kirches. MultiDimensional sum-up rounding for elliptic control systems. SIAM Journal on Numerical Analysis, 58 (6):3427–3447, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Masri:2024:DGM

[MKR24] Rami Masri, Miroslav Kuchta, and Beatrice Riviere. Discontinuous Galerkin methods for 3D–1D systems. SIAM Journal on Numerical Analysis, 62(4):1814–1843, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1627390.

Munoz-Matute:2024:MDP

[MMD24] Judit Muñoz-Matute and Leszek Demkowicz. Multistage discontinuous Petrov–Galerkin time-marching scheme for nonlinear problems. SIAM Journal on Numerical Analysis, 62 (4):1956–1978, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1598088.

Medeiros:2021:SOF

[MNO21] Debora D. Medeiros, Hirofumi Notsu, and Cassio M. Oishi. Second-order finite difference approximations of the upper-convected time derivative. SIAM Journal on Numerical Analysis, 59(6):2955–2988, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mohamad:2021:NIF

[MO21] Haidar Mohamad and Marcel Oliver. Numerical integration of functions of a rapidly rotating phase. SIAM Journal on Numerical Analysis, 59(4): 2310–2319, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Miraci:2020:MAE

[MPV20] Ani Miraçi, Jan Papez, and Martin Vohralík. A multilevel algebraic error estimator and the corresponding iterative solver with p-robust behavior. SIAM Journal on Numerical Analysis, 58(5):2856– 2884, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ma:2023:ESM

[MQ23] Limin Ma and Zhonghua Qiao. An energy stable and maximum bound principle preserving scheme for the dynamic Ginzburg–Landau equations under the temporal gauge. SIAM Journal on Numerical Analysis, 61(6):2695–2717, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mishra:2021:EAD

[MR21] Siddhartha Mishra and T. Konstantin Rusch. Enhancing accuracy of deep learning algorithms by training with lowdiscrepancy sequences. SIAM Journal on Numerical Analysis, 59(3):1811–1834, ???? 2021. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic).

Muga:2023:ASM

[MRV23] Ignacio Muga, Sergio Rojas, and Patrick Vega. An adaptive superconvergent mixed finite element method based on local residual minimization. SIAM Journal on Numerical Analysis, 61(5):2084–2105, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1526307.

Meyer:2020:PEA

[MS20] Christian Meyer and Michael Sievers. A priori error analysis of local incremental minimization schemes for rateindependent evolutions. SIAM Journal on Numerical Analysis, 58(4):2376–2403, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Madureira:2021:HLS

[MS21] Alexandre L. Madureira and Marcus Sarkis. Hybrid localized spectral decomposition for multiscale problems. SIAM Journal on Numerical Analysis, 59(2):829–863, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Marcati:2023:ECD

[MS23] Carlo Marcati and Christoph Schwab. Exponential convergence of deep operator networks for elliptic partial differential equations. SIAM Journal on Numerical Analysis, 61(3):1513–1545, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1465718.

Ma:2022:NDA

[MSD22] Chupeng Ma, Robert Scheichl, and Tim Dodwell. Novel design and analysis of generalized finite element methods based on locally optimal spectral approximations. SIAM Journal on Numerical Analysis, 60(1):244–273, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1406179.

Muller:2020:OOF

[MSS20] Stefan Müller, Florian Schweiger, and Endre Süli. Optimalorder finite difference approximation of generalized solutions to the biharmonic equation in a cube. SIAM Journal on Numerical Analysis, 58 (1):298–329, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Melenk:2020:WEA

[MST20] Jens M. Melenk, Stefan A. Sauter, and Céline Torres. Wavenumber explicit analysis for Galerkin discretizations of lossy Helmholtz problems. SIAM Journal on Numerical Analysis, 58(4):2119–2143, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mardal:2022:RPE

[MST22] Kent-Andre Mardal, Jarle Sogn, and Stefan Takacs. Robust preconditioning and error estimates for optimal control of the convection–diffusion– reaction equation with limited observation in isogeometric analysis. SIAM Journal on Numerical Analysis, 60(1):195–221, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M139147X.

Maire:2020:WCS

[MT20] P.-H. Maire and N. Therme.

Weak consistency of a staggered finite volume scheme for Lagrangian hydrodynamics. SIAM Journal on Numerical Analysis, 58(3):1592–1612, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Mu:2020:URW

[Mu20] Lin Mu. A uniformly robust H(DIV) weak Galerkin finite element methods for Brinkman problems. SIAM Journal on Numerical Analysis, 58(3): 1422–1439, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Manns:2022:SNM

[MU22] Paul Manns and Stefan Ulbrich. A simplified Newton method to generate snapshots for POD models of semilinear optimal control problems. SIAM Journal on Numerical Analysis, 60(5):2807–2833, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1439821.

Mustapha:2020:AFR

[Mus20] Kassem Mustapha. An L_{1} approximation for a fractional reaction–diffusion equation, a second-order error analysis over time-graded meshes. SIAM Journal on Numerical Analysis, 58(2):1319–1338, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Muga:2020:DLP

[MvdZ20] Ignacio Muga and Kristoffer G. van der Zee. Discretization of linear problems in Banach spaces: Residual minimization, nonlinear Petrov–Galerkin, and monotone mixed methods. SIAM Journal on Numerical Analysis, 58(6):3406–3426, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Meng:2024:KML

[MZ24] Shixu Meng and Bo Zhang.

A kernel machine learning for inverse source and scattering problems. SIAM Journal on Numerical Analysis, 62(3):1443–1464, June 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1597381.

Ma:2022:FOU

[MZZ22] Chuwen Ma, Qinghai Zhang, and Weiying Zheng. A fourth-order unfitted characteristic finite element method for solving the advection– diffusion equation on timevarying domains. SIAM Journal on Numerical Analysis, 60(4):2203–2224, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1483475.

Nie:2022:UCA

[ND22] Daxin Nie and Weihua Deng.

A unified convergence analysis for the fractional diffusion equation driven by fractional Gaussian noise with Hurst index H ∈ (0, 1). SIAM Journal on Numerical Analysis, 60(3):1548–1573, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1422616.

Nordstrom:2020:NBC

[NH20] Jan Nordström and Thomas M. Hagstrom. The number of boundary conditions for initial boundary value problems. SIAM Journal on Numerical Analysis, 58(5):2818–2828, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ni:2021:ALR

[Ni21] Angxiu Ni. Approximating linear response by nonintrusive shadowing algorithms. SIAM Journal on Numerical Analysis, 59(6):2843–2865, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Neilan:2021:DFS

[NO21] Michael Neilan and Baris Otus. Divergence-free Scott– Vogelius elements on curved domains. SIAM Journal on Numerical Analysis, 59(2): 1090–1116, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Noelle:2022:CBC

[NPT22] Sebastian Noelle, Martin Parisot, and Tabea Tscherpel. A class of boundary conditions for time-discrete Green–Naghdi equations with bathymetry. SIAM Journal on Numerical Analysis, 60 (5):2681–2712, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1426031.

Novo:2021:EAP

[NR21] Julia Novo and Samuele Rubino. Error analysis of proper orthogonal decomposition stabilized methods for incompressible flows. SIAM Journal on Numerical Analysis, 59 (1):334–369, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nochetto:2022:GCP

[NRY22] Ricardo H. Nochetto, Michele Ruggeri, and Shuo Yang. Gamma-convergent projectionfree finite element methods for nematic liquid crystals: The Ericksen model. SIAM Journal on Numerical Analysis, 60(2):856–887, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1407495.

Nie:2022:SCO

[NSD22] Daxin Nie, Jing Sun, and Weihua Deng. Strong convergence order for the scheme of fractional diffusion equation driven by fractional Gaussian noise. SIAM Journal on Numerical Analysis, 60 (4):1879–1904, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1356270.

Nakatsukasa:2021:ELB

[NT21a] Yuji Nakatsukasa and Alex Townsend. Error localization of best L_{1} polynomial approximants. SIAM Journal on Numerical Analysis, 59(1):314– 333, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Nakatsukasa:2021:RLA

[NT21b] Yuji Nakatsukasa and Lloyd N. Trefethen. Reciprocal-log approximation and planar PDE solvers. SIAM Journal on Numerical Analysis, 59(6):2801– 2822, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ning:2022:EET

[NWZ22] Cui Ning, Yifei Wu, and Xiaofei Zhao. An embedded exponential-type lowregularity integrator for mKdV equation. SIAM Journal on Numerical Analysis, 60 (3):999–1025, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1408166.

Ohsawa:2023:PQI

[Ohs23]  Tomoki Ohsawa.  Preser-

vation of quadratic invariants by semiexplicit symplectic integrators for nonseparable Hamiltonian systems. SIAM Journal on Numerical Analysis, 61(3):1293–1315, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1517718.

Owen:2024:GCS

[OP24] Art B. Owen and Zexin Pan.

Gain coefficients for scrambled Halton points. SIAM Journal on Numerical Analysis, 62(3): 1021–1038, May 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osborne:2024:ANA

[OS24] Yohance A. P. Osborne and Iain Smears. Analysis and numerical approximation of stationary second-order mean field game partial differential inclusions. SIAM Journal on Numerical Analysis, 62(1): 138–166, January 2024. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ostermann:2023:EER

[OSV23] Alexander Ostermann, Fardin Saedpanah, and Nasrin Vaisi. Explicit exponential Runge– Kutta methods for semilinear integro-differential equations. SIAM Journal on Numerical Analysis, 61(3):1405–1425, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1504056.

Otarola:2022:FSO

[Otá22] Enrique Otárola. Fractional semilinear optimal control: Optimality conditions, convergence, and error analysis. SIAM Journal on Numerical Analysis, 60(1):1–27, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1356294.

Papadopoulos:2022:NAD

[Pap22] Ioannis P. A. Papadopoulos. Numerical analysis of a discontinuous Galerkin method for the Borrvall–Petersson topology optimization problem. SIAM Journal on Numerical Analysis, 60(5):2538–2564, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1438943.

Park:2020:ASM

[Par20]  Jongho Park.  Additive

Schwarz methods for convex optimization as gradient methods. SIAM Journal on Numerical Analysis, 58(3):1495–1530, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peng:2021:SEA

[PCQL21] Zhichao Peng, Yingda Cheng, Jing-Mei Qiu, and Fengyan Li. Stability-enhanced AP IMEX1-LDG method: Energybased stability and rigorous AP property. SIAM Journal on Numerical Analysis, 59(2): 925–954, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Pareschi:2022:MPF

[PR22] Lorenzo Pareschi and Thomas Rey. Moment preserving Fourier–Galerkin spectral methods and application to the Boltzmann equation. SIAM Journal on Numerical Analysis, 60(6):3216–3240, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1423452.

Pradovera:2020:IRM

[Pra20]  Davide Pradovera.  Inter-

polatory rational model order reduction of parametric problems lacking uniform inf– sup stability. SIAM Journal on Numerical Analysis, 58(4): 2265–2293, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Peszynska:2020:ASC

[PS20] Małgorzata Peszynska and Ralph E. Showalter. Approximation of scalar conservation law with hysteresis. SIAM Journal on Numerical Analysis, 58(2):962–987, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Potts:2021:AHD

[PS21] Daniel Potts and Michael Schmischke. Approximation of high-dimensional periodic functions with Fourier-based methods. SIAM Journal on Numerical Analysis, 59(5): 2393–2429, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Perrin:2022:NSS

[PS22] Charlotte Perrin and Khaled Saleh. Numerical staggered schemes for the freecongested Navier–Stokes equations. SIAM Journal on Numerical Analysis, 60(4): 1824–1852, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1436488.

Parker:2024:SLP

[PS24] Charles Parker and Endre Süli. Stable lifting of polynomial traces on triangles. SIAM Journal on Numerical Analysis, 62(2):692–717, March 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Przybylowicz:2022:EAS

[PSS22] Paweł Przybyłowicz, Michał Sobieraj, and Lukasz Stepień. Efficient approximation of SDEs driven by countably dimensional Wiener process and Poisson random measure. SIAM Journal on Numerical Analysis, 60(2):824–855, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1442747.

Pozharska:2022:NSR

[PU22] Kateryna Pozharska and Tino Ullrich. A note on sampling recovery of multivariate functions in the uniform norm. SIAM Journal on Numerical Analysis, 60(3):1363–1384, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1410580.

Qian:2021:FMG

[QS21] Jianliang Qian and Chao Song. Fast multiscale Gaussian beam method for threedimensional elastic wave equations in bounded domains. SIAM Journal on Numerical Analysis, 59(5):2536–2570, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Qi:2020:EES

[QW20] Ruisheng Qi and Xiaojie Wang. Error estimates of semidiscrete and fully discrete finite element methods for the Cahn–Hilliard–Cook equation. SIAM Journal on Numerical Analysis, 58(3):1613–1653, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Quan:2023:NSC

[QW23] Chaoyu Quan and Xu Wu. H^{1}-norm stability and convergence of an L2-type method on nonuniform meshes for subdiffusion equation. SIAM Journal on Numerical Analysis, 61(5):2106–2132, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1506468.

Qin:2022:VTS

[QWL22] Yi Qin, Yongshuai Wang, and Jian Li. A variable time step time filter algorithm for the geothermal system. SIAM Journal on Numerical Analysis, 60(5):2781–2806, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1464828.

Qiu:2020:AFO

[QZ20] Weifeng Qiu and Shun Zhang. Adaptive first-order system least-squares finite element methods for second-order elliptic equations in nondivergence form. SIAM Journal on Numerical Analysis, 58(6): 3286–3308, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ranocha:2020:ESE

[RK20] Hendrik Ranocha and David I.

Ketcheson. Energy stability of explicit Runge–Kutta methods for nonautonomous or nonlinear problems. SIAM Journal on Numerical Analysis, 58(6): 3382–3405, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ruggiu:2020:EAS

[RN20] Andrea Alessandro Ruggiu and Jan Nordström. Eigenvalue analysis for summationby-parts finite difference time discretizations. SIAM Journal on Numerical Analysis, 58 (2):907–928, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reusken:2021:ASD

[RS21a] Arnold Reusken and Benjamin Stamm. Analysis of the Schwarz domain decomposition method for the conductor-like screening continuum model. SIAM Journal on Numerical Analysis, 59 (2):769–796, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rousset:2021:GFL

[RS21b] Frédéric Rousset and Katharina Schratz. A general framework of low regularity integrators. SIAM Journal on Numerical Analysis, 59(3):1735– 1768, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rumpf:2022:FEA

[RSS22] Martin Rumpf, Stefan Simon, and Christoph Smoch. Finite element approximation of large-scale isometric deformations of parametrized surfaces. SIAM Journal on Numerical Analysis, 60(5):2945–2962, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1455292.

Rubino:2020:NAP

[Rub20] Samuele Rubino. Numerical analysis of a projection-based stabilized POD-ROM for incompressible flows. SIAM Journal on Numerical Analysis, 58(4):2019–2058, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rieger:2024:ACD

[RW24] Christian Rieger and Holger Wendland. On the approximability and curse of dimensionality of certain classes of highdimensional functions. SIAM Journal on Numerical Analysis, 62(2):842–871, March 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sharan:2022:TSS

[SBL22] Nek Sharan, Peter T. Brady, and Daniel Livescu. Time stability of strong boundary conditions in finite-difference schemes for hyperbolic systems. SIAM Journal on Numerical Analysis, 60(3): 1331–1362, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1419957.

Shao:2023:CAN

[SC23a] Nian Shao and Wenbin Chen.

Convergence analysis of Newton– Schur method for symmetric elliptic eigenvalue problem. SIAM Journal on Numerical Analysis, 61(1):315–342, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1448847.

Shi:2023:HOB

[SC23b] Jiankang Shi and Minghua Chen. High-order BDF convolution quadrature for subdiffusion models with a singular source term. SIAM Journal on Numerical Analysis, 61(6):2559–2579, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scheel:2023:NEM

[Sch23]  Arnd Scheel.  Nonlinear

eigenvalue methods for linear pointwise stability of nonlinear waves. SIAM Journal on Numerical Analysis, 61(2):592–616, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1492969.

Sarna:2020:CAG

[SGT20] Neeraj Sarna, Jan Giesselmann, and Manuel Torrilhon. Convergence analysis of Grad’s Hermite expansion for linear kinetic equations. SIAM Journal on Numerical Analysis, 58 (2):1164–1194, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scholz:2021:UHO

[SJ21] Felix Scholz and Bert Jüttler. Using high-order transport theorems for implicitly defined moving curves to perform quadrature on planar domains. SIAM Journal on Numerical Analysis, 59(4):2138– 2162, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sprekeler:2024:HNF

[Spr24] Timo Sprekeler. Homogenization of nondivergence-form elliptic equations with discontinuous coefficients and finite element approximation of the homogenized problem. SIAM Journal on Numerical Analysis, 62(2):646–666, March 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sonner:2020:SOP

[SR20] Florian Sonner and Thomas Richter. Second order pressure estimates for the Crank– Nicolson discretization of the incompressible Navier–Stokes equations. SIAM Journal on Numerical Analysis, 58(1): 375–409, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sheng:2020:FFL

[SST^{+}20] Changtao Sheng, Jie Shen, Tao Tang, Li-Lian Wang, and Huifang Yuan. Fast Fourier-like mapped Chebyshev spectral-Galerkin methods for PDEs with integral fractional Laplacian in unbounded domains. SIAM Journal on Numerical Analysis, 58(5):2435–2464, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sheng:2023:EMC

[SSX23] Changtao Sheng, Bihao Su, and Chenglong Xu. Efficient Monte Carlo method for integral fractional Laplacian in multiple dimensions. SIAM Journal on Numerical Analysis, 61(5):2035–2061, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1504706.

Shi:2023:EDD

[SSZ23] Feng Shi, Yizhong Sun, and Haibiao Zheng. Ensemble domain decomposition algorithm for the fully mixed random Stokes–Darcy model with the Beavers–Joseph interface conditions. SIAM Journal on Numerical Analysis, 61 (3):1482–1512, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1482846.

Stamm:2022:QOF

[ST22] Benjamin Stamm and Lambert Theisen. A quasi-optimal factorization preconditioner for periodic Schrödinger eigenstates in anisotropically expanding domains. SIAM Journal on Numerical Analysis, 60(5):2508–2537, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1456005.

Storn:2020:CLC

[Sto20] Johannes Storn. Computation of the LBB constant for the Stokes equation with a leastsquares finite element method. SIAM Journal on Numerical Analysis, 58(1):86–108, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sun:2021:ALO

[Sun21]  Weiwei Sun.  Analysis of

lowest-order characteristicsmixed FEMs for incompressible miscible flow in porous media. SIAM Journal on Numerical Analysis, 59(4):1875–1895, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Scarabel:2024:EID

[SV24] Francesca Scarabel and Rossana Vermiglio. Equations with infinite delay: Pseudospectral discretization for numerical stability and bifurcation in an abstract framework. SIAM Journal on Numerical Analysis, 62(4):1736–1758, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1581133.

Schluchtermann:2023:NSF

[SW23] Georg Schlüchtermann and Michael Wibmer. Numerical solution of free stochastic differential equations. SIAM Journal on Numerical Analysis, 61(6):2623–2650, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sun:2022:ELS

[SWW22] Zheng Sun, Yuanzhe Wei, and Kailiang Wu. On energy laws and stability of Runge–Kutta methods for linear seminegative problems. SIAM Journal on Numerical Analysis, 60(5):2448–2481, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1472218.

Sun:2020:FEA

[SWZ20] Ting Sun, Jilu Wang, and Chunxiong Zheng. Fast evaluation of artificial boundary conditions for advection diffusion equations. SIAM Journal on Numerical Analysis, 58(6): 3530–3557, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shao:2020:BRW

[SX20a] Sihong Shao and Yunfeng Xiong. Branching random walk solutions to the Wigner equation. SIAM Journal on Numerical Analysis, 58(5): 2589–2608, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Shen:2020:UBP

[SX20b] Jie Shen and Jie Xu. Unconditionally bound preserving and energy dissipative schemes for a class of Keller–Segel equations. SIAM Journal on Numerical Analysis, 58(3):1674– 1695, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Serna:2020:CRK

[SZ20] J. M. Sanz Serna and K. C.

Zygalakis. Contractivity of Runge–Kutta methods for convex gradient systems. SIAM Journal on Numerical Analysis, 58(4):2079–2092, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Serna:2021:CBL

[SZ21a] J. M. Sanz Serna and K. C.

Zygalakis. The connections between Lyapunov functions for some optimization algorithms and differential equations. SIAM Journal on Numerical Analysis, 59(3):1542– 1565, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sharma:2021:NSS

[SZ21b] Upanshu Sharma and Wei Zhang. NonReversible sampling schemes on submanifolds. SIAM Journal on Numerical Analysis, 59(6):2989– 3031, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tong:2024:PPM

[TC24] Fenghua Tong and Yongyong Cai. Positivity preserving and mass conservative projection method for the Poisson–Nernst–Planck equation. SIAM Journal on Numerical Analysis, 62(4):??, August 2024. CODEN SJNAAM.

ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1581649.

Tran:2024:DWD

[Tra24] Ngoc Tien Tran. Discrete weak duality of hybrid highorder methods for convex minimization problems. SIAM Journal on Numerical Analysis, 62(4):1492–1514, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1594534.

Takeda:2023:GEH

[TS23] Kota Takeda and Takashi Sakajo. Geometric ergodicity for Hamiltonian Monte Carlo on compact manifolds. SIAM Journal on Numerical Analysis, 61(6):2994–3013, December 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Tai:2023:UPN

[TWZZ23] Xue-Cheng Tai, Ragnar Winther, Xiaodi Zhang, and Weiying Zheng. A uniform preconditioner for a Newton algorithm for total variation minimization and minimumsurface problems. SIAM Journal on Numerical Analysis, 61(5):2062–2083, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1512776.

Vaccaro:2023:AOD

[VL23] Jennifer Vaccaro and Konstantin Lipnikov. Applying an oriented divergence theorem to swept face remap. SIAM Journal on Numerical Analysis, 61(5):2285–2304, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1518359.

vonWahl:2023:EAP

[vWR23] Henry von Wahl and Thomas Richter. Error analysis for a parabolic PDE model problem on a coupled moving domain in a fully Eulerian framework. SIAM Journal on Numerical Analysis, 61(1):286–314, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458417.

Wang:2023:AEL

[Wan23] Haiyong Wang. Analysis of error localization of Chebyshev spectral approximations. SIAM Journal on Numerical Analysis, 61(2):952–972, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1481452.

Wu:2021:EAE

[WB21] Yongke Wu and Yanhong Bai. Error analysis of energypreserving mixed finite element methods for the Hodge wave equation. SIAM Journal on Numerical Analysis, 59(3): 1433–1454, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Williams:2024:FSD

[WD24] Christopher Williams and Kenneth Duru. Full-spectrum dispersion relation preserving summation-by-parts operators. SIAM Journal on Numerical Analysis, 62 (4):1565–1588, August 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1586471.

Wei:2021:PDF

[WHL21] Huayi Wei, Xuehai Huang, and Ao Li. Piecewise divergencefree nonconforming virtual elements for Stokes problem in any dimensions. SIAM Journal on Numerical Analysis, 59 (3):1835–1856, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wu:2023:PPC

[WJS23] Kailiang Wu, Haili Jiang, and Chi-Wang Shu. Provably positive central discontinuous Galerkin schemes via geometric quasilinearization for ideal MHD equations. SIAM Journal on Numerical Analysis, 61(1):250–285, ???? 2023.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1486996.

Wang:2023:FDD

[WK23] Siyang Wang and Gunilla Kreiss. A finite differencediscontinuous Galerkin method for the wave equation in second order form. SIAM Journal on Numerical Analysis, 61(4):1962–1988, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1530690.

Wan:2020:PEE

[WL20] Andy T. S. Wan and Marc Laforest. A posteriori error estimation for the p-curl problem. SIAM Journal on Numerical Analysis, 58(1):460–491, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:2023:ESP

[WL23] Wansheng Wang and Shoufu Li. Efficient stability-preserving numerical methods for nonlinear coercive problems in vector space. SIAM Journal on Numerical Analysis, 61(2):872–904, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1443273.

Wu:2023:AFT

[WLF23] Jilian Wu, Ning Li, and Xinlong Feng. Analysis of a filtered time-stepping finite element method for natural convection problems. SIAM Journal on Numerical Analysis, 61(2):837–871, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1451476.

Wagner:2021:EAP

[WLPU21] Fabian Wagner, Jonas Latz, Iason Papaioannou, and Elisabeth Ullmann. Error analysis for probabilities of rare events with approximate models. SIAM Journal on Numerical Analysis, 59(4):1948–1975, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wu:2023:UTQ

[WM23] Bowei Wu and Per-Gunnar Martinsson. A unified trapezoidal quadrature method for singular and hypersingular boundary integral operators on curved surfaces. SIAM Journal on Numerical Analysis, 61(5):2182–2208, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1520372.

Wormell:2021:SCD

[WR21] Caroline L. Wormell and Sebastian Reich. Spectral convergence of diffusion maps: Improved error bounds and an alternative normalization. SIAM Journal on Numerical Analysis, 59(3):1687–1734, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:2023:OLT

[WS23] Dongling Wang and Martin Stynes. Optimal long-time decay rate of numerical solutions for nonlinear time-fractional evolutionary equations. SIAM Journal on Numerical Analysis, 61(5):2011–2034, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1494361.

Weidling:2020:OCR

[WSH20] Frederic Weidling, Benjamin Sprung, and Thorsten Hohage. Optimal convergence rates for Tikhonov regularization in Besov spaces. SIAM Journal on Numerical Analysis, 58(1):21–47, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:2020:PDW

[WW20] Chunmei Wang and Junping Wang. A primal-dual weak Galerkin finite element method for Fokker–Planck type equations. SIAM Journal on Numerical Analysis, 58(5):

2632–2661, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wu:2020:SAF

[WX20] Shuonan Wu and Jinchao Xu. Simplex-averaged finite element methods for H(grad), H(curl), and H(div) convection–diffusion problems. SIAM Journal on Numerical Analysis, 58(1):884–906, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wei:2020:CAP

[WYZ20a] Changkun Wei, Jiaqing Yang, and Bo Zhang. Convergence analysis of the PML method for time-domain electromagnetic scattering problems. SIAM Journal on Numerical Analysis, 58(3):1918– 1940, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Winckler:2020:AEE

[WYZ20b] Malte Winckler, Irwin Yousept, and Jun Zou. Adaptive edge element approximation for H(curl) elliptic variational inequalities of second kind. SIAM Journal on Numerical Analysis, 58(3):1941–1964, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:2024:EAF

[WYZ24] Li-Lian Wang, Jingye Yan, and Xiaolong Zhang. Error analysis of a first-order IMEX scheme for the logarithmic Schrödinger equation. SIAM Journal on Numerical Analysis, 62(1):119–137, January 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:2020:HOT

[WZ20] Kai Wang and Zhi Zhou. Highorder time stepping schemes for semilinear subdiffusion equations. SIAM Journal on Numerical Analysis, 58(6): 3226–3250, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:2021:EES

[WZ21] Bin Wang and Xiaofei Zhao.

Error estimates of some splitting schemes for chargedparticle dynamics under strong magnetic field. SIAM Journal on Numerical Analysis, 59(4): 2075–2105, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Wang:2023:GTS

[WZ23] Bin Wang and Xiaofei Zhao.

Geometric two-scale integrators for highly oscillatory system: Uniform accuracy and near conservations. SIAM Journal on Numerical Analysis, 61(3):1246–1277, ????

2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462908.

Wang:2022:SSB

[WZZ22] Xu Wang, Weidong Zhao, and Tao Zhou. Sinc-θ schemes for backward stochastic differential equations. SIAM Journal on Numerical Analysis, 60(4):1799–1823, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1444679.

Xiang:2021:CRS

[Xia21] Shuhuang Xiang. Convergence rates on spectral orthogonal projection approximation for functions of algebraic and logarithmatic regularities. SIAM Journal on Numerical Analysis, 59(3):1374– 1398, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Xiong:2024:CCD

[XLZZ24] Zhe Xiong, Lei Li, Ya-Nan Zhu, and Xiaoqun Zhang. On the convergence of continuous and discrete unbalanced optimal transport models for 1-Wasserstein distance. SIAM Journal on Numerical Analysis, 62(2):749–774, March 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Xu:2020:EEF

[XSZ20] Yuan Xu, Chi-Wang Shu, and Qiang Zhang. Error estimate of the fourth-order Runge– Kutta discontinuous Galerkin methods for linear hyperbolic equations. SIAM Journal on Numerical Analysis, 58(5): 2885–2914, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Xu:2022:CAI

[XZ22] Xuefeng Xu and Chen-Song Zhang. Convergence analysis of inexact two-grid methods: A theoretical framework. SIAM Journal on Numerical Analysis, 60(1):133–156, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1356075.

Ye:2021:CDF

[YC21] Changqing Ye and Junzhi Cui. Convergence of Dziuk’s fully discrete linearly implicit scheme for curve shortening flow. SIAM Journal on Numerical Analysis, 59(6):2823– 2842, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yu:2021:AAS

[YDS21] Yi Yu, Maksymilian Dryja, and Marcus Sarkis. From additive average Schwarz methods to nonoverlapping spectral additive Schwarz methods. SIAM Journal on Numerical Analysis, 59(5):2608–2638, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yu:2022:PHA

[YHLR22] Xiuchen Yu, Guanghui Hu, Wangtao Lu, and Andreas Rathsfeld. PML and highaccuracy boundary integral equation solver for wave scattering by a locally defected periodic surface. SIAM Journal on Numerical Analysis, 60(5):2592–2625, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1439705.

Yi:2022:LFE

[YLZ22] Son-Young Yi, Sanghyun Lee, and Ludmil Zikatanov. Locking-free enriched Galerkin method for linear elasticity. SIAM Journal on Numerical Analysis, 60(1):52–75, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1391353.

Ye:2020:SFW

[YZ20] Xiu Ye and Shangyou Zhang. A stabilizer free weak Galerkin method for the biharmonic equation on polytopal meshes. SIAM Journal on Numerical Analysis, 58(5):2572–2588, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yang:2023:PSS

[YZDCNC23] Minglei Yang, Guannan Zhang, Diego Del-Castillo-Negrete, and Yanzhao Cao. A probabilistic scheme for semilinear nonlocal diffusion equations with volume constraints. SIAM Journal on Numerical Analysis, 61(6):2718–2743, November 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Yang:2020:UPD

[YZZ20] Jie Yang, Weidong Zhao, and Tao Zhou. A unified probabilistic discretization scheme for FBSDEs: Stability, consistency, and convergence analysis. SIAM Journal on Numerical Analysis, 58(4):2351–2375, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhao:2021:SDM

[ZCPZ21] Lina Zhao, Eric T. Chung, Eun-Jae Park, and Guanyu Zhou. Staggered DG method for coupling of the Stokes and Darcy–Forchheimer problems. SIAM Journal on Numerical Analysis, 59(1):1–31, ???? 2021. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zheng:2020:SEA

[ZDMZ20] Chunxiong Zheng, Qiang Du, Xiang Ma, and Jiwei Zhang. Stability and error analysis for a second-order fast approximation of the local and nonlocal diffusion equations on the real line. SIAM Journal on Numerical Analysis, 58(3):1893–1917, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zaitzeff:2020:VEI

[ZEG20] Alexander Zaitzeff, Selim Esedoáğlu, and Krishna Garikipati. Variational extrapolation of implicit schemes for general gradient flows. SIAM Journal on Numerical Analysis, 58(5): 2799–2817, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhang:2022:ECP

[Zha22] Ruming Zhang. Exponential convergence of perfectly matched layers for scattering problems with periodic surfaces. SIAM Journal on Numerical Analysis, 60 (2):804–823, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1439043.

Zhang:2023:HOC

[Zha23a]  Ruming Zhang.  Higher-

order convergence of perfectly matched layers in threedimensional biperiodic surface scattering problems. SIAM Journal on Numerical Analysis, 61(6):2917–2939, December 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhang:2023:SDM

[Zha23b]  Ruming Zhang.  A spec-

tral decomposition method to approximate Dirichlet-to– Neumann maps in complicated waveguides. SIAM Journal on Numerical Analysis, 61(3):1195–1217, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1485425.

Zhou:2021:NAR

[ZL21] Jiaxin Zhou and Wangtao Lu.

Numerical analysis of resonances by a slab of subwavelength slits by Fouriermatching method. SIAM Journal on Numerical Analysis, 59 (4):2106–2137, ???? 2021. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhang:2023:CAL

[ZN23] Erchuan Zhang and Lyle Noakes. Convergence analysis of leapfrog for geodesics. SIAM Journal on Numerical Analysis, 61(5):2261–2284, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1515173.

Zouraris:2023:EER

[Zou23] Georgios E. Zouraris. Error estimation of the relaxation finite difference scheme for the nonlinear Schrödinger equation. SIAM Journal on Numerical Analysis, 61 (1):365–397, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1399130.

Zhong:2023:SAM

[ZQ23] Xiang Zhong and Weifeng Qiu. Spectral analysis of a mixed method for linear elasticity. SIAM Journal on Numerical Analysis, 61(4):1885–1917, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M148611X.

Zhang:2023:SNP

[ZRZ23] Yuhong Zhang, Yao Rong, and Haibiao Zheng. A study of numerical pollution of the decoupled algorithm for the convection model in superposed fluid and porous layers. SIAM Journal on Numerical Analysis, 61(2):1018–1056, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469699.

Zhang:2020:WPF

[ZSH20] Yuhong Zhang, Li Shan, and Yanren Hou. Well-posedness and finite element approximation for the convection model in superposed fluid and porous layers. SIAM Journal on Numerical Analysis, 58(1):541– 564, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zheng:2020:EEN

[ZW20a] Xiangcheng Zheng and Hong Wang. An error estimate of a numerical approximation to a hidden-memory variableorder space-time fractional diffusion equation. SIAM Journal on Numerical Analysis, 58(5): 2492–2514, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zheng:2020:OON

[ZW20b] Xiangcheng Zheng and Hong Wang. An optimal-order numerical approximation to variable-order space-fractional diffusion equations on uniform or graded meshes. SIAM Journal on Numerical Analysis, 58 (1):330–352, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhang:2022:HOF

[ZW22] Lu Zhang and Siyang Wang.

A high order finite difference method for the elastic wave equation in bounded domains with nonconforming interfaces. SIAM Journal on Numerical Analysis, 60(3):1516–1547, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1422586.

Zhou:2023:DAC

[ZW23] Yu Zhou and Haijun Wu. Dispersion analysis of CIP-FEM for the Helmholtz equation. SIAM Journal on Numerical Analysis, 61(3):1278–1292, ???? 2023. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M143827X.

Zhu:2024:BPB

[ZW24] Bingxin Zhu and Haijun Wu. The (p, p−1)-HDG method for the Helmholtz equation with large wave number. SIAM Journal on Numerical Analysis, 62(3):1394–1419, June 2024. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1562639.

Zhang:2020:ECM

[ZWC20] Wenzhong Zhang, Bo Wang, and Wei Cai. Exponential convergence for multipole and local expansions and their translations for sources in layered media: Two-dimensional acoustic wave. SIAM Journal on Numerical Analysis, 58(3):

1440–1468, ???? 2020. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Zhang:2022:EEE

[ZZZ22a] Lei Zhang, Pingwen Zhang, and Xiangcheng Zheng. Error estimates for Euler discretization of high-index saddle dynamics. SIAM Journal on Numerical Analysis, 60(5):2925–2944, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1458314.

Zhang:2022:IPD

[ZZZ22b] Zhengqi Zhang, Zhidong Zhang, and Zhi Zhou. Identification of potential in diffusion equations from terminal observation: Analysis and discrete approximation. SIAM Journal on Numerical Analysis, 60(5):2834–2865, ???? 2022. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1446708.